### v1

# Data-Driven Design Rules for TADF Emitters from a High-Throughput Screening of $747$ Molecules

Abstract
[…]
I Introduction
Organic Light-Emitting Diodes (OLEDs) based on Thermally Activated Delayed Fluorescence (TADF) have defined the third generation of electroluminescent materials, offering a viable path to achieving near-unity internal quantum efficiency (IQE) without resorting to the expensive and scarce noble metals found in phosphorescent emitters [1, 2, 3]. The TADF mechanism hinges on the efficient harvesting of non-emissive triplet excitons ($T_{1}$) by up-converting them to the emissive singlet state ($S_{1}$) through Reverse Intersystem Crossing (RISC). This process requires satisfying both a thermodynamic condition—a minimized singlet-triplet energy gap ($\Delta E_{\text{ST}}\qty$$\lesssim\qty{0.2}{}$)—and a kinetic condition: a sufficiently high rate of RISC ($k_{\text{RISC}}$) [4, 5].
The rational design of high-performance TADF emitters is a formidable challenge, dominated by a fundamental trade-off. In conventional Donor-Acceptor (D-A) architectures, achieving a small $\Delta E_{\text{ST}}$necessitates a significant spatial separation of the frontier molecular orbitals to minimize the electron exchange energy [6]. However, this separation inherently reduces the orbital overlap, which not only suppresses the oscillator strength ($f$) and thus the radiative decay rate, but can also hinder the spin-orbit coupling (SOC) necessary for efficient RISC [7, 8]. Successfully navigating this delicate balance requires exploring a vast and complex chemical space [9].
Computational chemistry is an indispensable tool for navigating this landscape, yet it presents its own dilemma. High-fidelity methods are computationally prohibitive for screening the thousands of potential candidates required for data-driven discovery [10, 11, 12]. This scalability bottleneck has historically limited computational studies to small, homologous series of molecules, hindering the extraction of universal design principles.
To bridge this gap, we employ a high-throughput virtual screening (HTVS) protocol based on accelerated semi-empirical methods, which we rigorously validated on a diverse dataset of $747$ experimentally known TADF emitters in our preceding companion paper (hereafter, Article 1) [13]. That study established our framework as a powerful and validated tool for large-scale molecular ranking and trend analysis.
In this article, we leverage our validated HTVS protocol to move beyond benchmarking and perform a large-scale, data-driven analysis aimed at elucidating the fundamental structure-property relationships that govern TADF performance. By analyzing the results from the entire 747-molecule dataset, we aim to distill the complex interplay of molecular structure, conformation, and electronic properties into a set of clear, quantitative, and actionable design guidelines. This study, therefore, seeks to: (i) establish a quantitative performance hierarchy among diverse molecular architectures (D-A, D-A-D, and MR systems); (ii) derive statistically robust design rules linking key structural parameters to both thermodynamic ($\Delta E_{\text{ST}}$) and kinetic (SOC, $k_{\text{RISC}}$) factors; (iii) identify and characterize distinct molecular families through data-driven clustering analysis; and (iv) propose a curated list of high-priority candidate molecules predicted to exhibit exceptional TADF properties. By translating a massive computational dataset into actionable design intelligence, this work provides a unified framework to guide and accelerate the discovery of next-generation TADF emitters.
This paper is structured as follows: Section II details our methodological framework. The main results are presented in the subsequent sections, including an analysis of the architectural landscape, a classification of molecular families, and a deep dive into the MR-TADF design paradigm. Finally, Section IV summarizes our conclusions and proposes future directions for TADF emitter design.
II Methodological framework for large-scale analysis
All structure-property relationships presented in this work were derived from a large-scale dataset of $747$ experimentally known TADF emitters, processed using a consistent and validated high-throughput virtual screening (HTVS) protocol. The full details, validation, and limitations of this computational framework are established in our preceding work (Article 1) [13]. Here, we provide a concise summary of the essential methodological components and their justifications.
II.1 Dataset and molecular structures
The dataset comprises a diverse set of $747$ TADF molecules curated from the literature [14]. To facilitate systematic analysis, molecules were classified into distinct architectural families (e.g., D-A, D-A-D, Multi-Resonance) using automated SMARTS pattern matching, followed by manual verification.
Initial 3D structures were generated from SMILES strings, and a rigorous conformational search was performed for each molecule using the CREST program coupled with the GFN2-xTB semi-empirical Hamiltonian [15, 16]. The lowest-energy conformer was then fully optimized at the GFN2-xTB level to obtain the final ground-state ($S_{0}$) geometry. This tight-binding method is specifically parameterized for accurate molecular structures and noncovalent interactions, making it well-suited for the complex and often flexible systems studied here [17].
II.2 Methodological approximations and justifications
II.2.1 Hybrid protocol and the vertical approximation compromise
The HTVS protocol relies on a hybrid computational strategy: ground-state ($S_{0}$) geometries are optimized using GFN2-xTB [16, 18], while excited-state properties ($S_{1}$, $T_{1}$) are calculated using the simplified methods, sTDA-xTB/sTD-DFT-xTB [19, 20]. This methodology implicitly employs the vertical approximation.
[…]
II.2.2 Limitations of the solvation model
The study of excited-state properties was performed using the ground-state-based ALPB implicit solvation model for toluene. We acknowledge that for the highly polarized charge-transfer (CT) states characteristic of TADF emitters, a state-specific solvation model (e.g., SS-PCM) is theoretically required for quantitative accuracy. However, such approaches are computationally prohibitive for HTS. The ALPB model is considered sufficient for screening as it has been shown that while the absolute solvation effect may be approximated, the relative ordering of molecules and the overall structure-property trends are conserved in the condensed phase.
II.2.3 Specific limitations for multi-resonance emitters
It is important to note that Multi-Resonance (MR) TADF emitters are inherently susceptible to multi-reference character [22], a feature that current single-reference based TD-DFT and its semi-empirical variants (sTDA-xTB/sTD-DFT-xTB) may not fully capture. Consequently, while our HTS trends provide excellent initial screening recommendations, definitive quantitative validation for MR-TADF candidates should rely on post-hoc applications of high-level methods such as SCS-CC2 or STEOM-DLPNO-CCSD [23].
II.3 Derivation of structure-property descriptors and theoretical framework
To translate the large computational dataset into actionable design rules, donor-acceptor (D-A) torsional angles were automatically extracted from all optimized geometries. To quantify the degree of charge-transfer and its relation to geometry, we calculated the HOMO-LUMO overlap integral ($S^{\prime}_{\text{HL}}$) and the distance between the centroids of the hole and electron densities ($\Delta r_{\text{CT}}$) using the Multiwfn package [24, 25]. Principal Component Analysis (PCA) was performed on the standardized matrix of all descriptors to identify the underlying variables governing TADF performance.
This analysis is grounded in the theoretical framework of TADF efficiency, which depends not only on a small $\Delta E_{\text{ST}}$(thermodynamics) but also on a high rate of reverse intersystem crossing ($k_{\text{RISC}}$), a kinetic parameter governed by spin-orbit coupling (SOC). According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation) [26, 27]. Our chosen descriptors—particularly the D-A torsional angle—are direct proxies for tuning this CT/LE character balance. While our protocol does not compute SOC explicitly due to its prohibitive cost, analyzing these structural and electronic descriptors provides a robust, qualitative assessment of the kinetic factors that, alongside $\Delta E_{\text{ST}}$, govern overall TADF performance.
III Results and discussion: from statistical trends to rational design principles
This section leverages the large-scale dataset of $747$ TADF emitters, processed with the validated HTVS protocol detailed in our companion paper (Article 1)[13], to extract quantitative, statistically robust structure-property relationships. Our analysis transitions from broad statistical trends to specific, actionable molecular design rules, with a focus on providing a deep mechanistic understanding that connects molecular structure to the key photophysical parameters governing TADF efficiency.
III.1 Low-dimensional nature of the TADF design space
[…]
III.2 Architectural hierarchy defines TADF potential
[…]
Figure 2: TADF performance stratified by molecular architecture. The box plots show the distribution of the singlet-triplet energy gap ($\Delta E_{\text{ST}}$) for each class. D-A-D architectures demonstrate statistically superior characteristics (lower median and narrower distribution of $\Delta E_{\text{ST}}$) compared to other systems, confirming the efficacy of this design principle for consistently achieving small energy gaps.
III.3 Data-driven clustering reveals distinct molecular families
To complement the structure-based classification, we performed an unsupervised K-means clustering based on the molecules’ photophysical properties. This data-driven approach identifies four distinct molecular families, as shown in Figure 3:
[…]
  • Cluster 3 (<<1%, n=1): Outliers. A single outlier with extreme properties requiring individual investigation.
This analysis confirms that while structural classification is useful, the photophysical property space naturally segregates molecules into performance-based families, with a clear cluster of high-potential candidates emerging directly from the data.
Figure 3: K-means clustering (k=4) of the $747$ TADF emitters based on their standardized photophysical properties in both the gas phase (a) and toluene solvent (b). The analysis partitions the dataset into distinct families, most notably identifying a cluster of high-efficiency candidates (Cluster 1) with desirable properties for TADF applications.
III.4 Multi-resonance TADF: a distinct paradigm for blue emitters
[…]
However, as noted in the methodology, it is important to interpret these results with caution. The inherent electronic complexity of MR-TADF systems makes them prone to multi-reference character [22], which is not fully captured by the single-reference nature of the semi-empirical methods used in this HTS. The trends presented should therefore be interpreted as a valuable guide for compound ranking rather than a source of definitive quantitative predictions, with final validation requiring higher levels of theory.
Table 1: Photophysical property comparison of MR-TADF with conventional D-A systems, calculated with sTDA-xTB in toluene. MR emitters achieve a small $\Delta E_{\text{ST}}$comparable to D-A-D systems but with significantly higher oscillator strength and blue-shifted emission, highlighting a different design strategy.
Property MR-TADF (n=67) D-A-D (n=182) Simple D-A (n=248)
[…]
Oscillator Strength ($f$) $0.61$ p m 0.58 $0.40$ p m 0.52 $0.51$ p m 0.58
[…]
III.5 Deconstructing performance: geometric and electronic design levers
III.5.1 Torsional angle as a primary control parameter
[…]
III.5.2 Quantifying the torsion-induced electronic trade-off
This geometric control is directly reflected in the electronic structure. As shown in Figure 5, twisting the molecule reduces the HOMO-LUMO overlap ($S^{\prime}_{\text{HL}}$), which correlates strongly with a smaller $\Delta E_{\text{ST}}$(Spearman $\rho=-0.45$). However, this comes at a cost: the oscillator strength, essential for a high radiative rate, is directly proportional to this overlap ($\rho=0.78$). This visualization makes the fundamental TADF design trade-off explicit: minimizing the energy gap via orbital separation inherently suppresses emission intensity. An optimal region for high-performance emitters lies where there is intermediate overlap (0.15 to 0.300.150.30) that balances these competing requirements.
To move beyond this qualitative picture and establish a more rigorous quantitative link, future work could employ methods such as Natural Transition Orbitals (NTOs) or the RespA procedure [29, 30]. These tools would allow for a compact representation of the virtual excitations and rigorously connect the molecular geometry to the precise charge-transfer character (e.g., via centroid distance $\Delta r$) and the ensuing coupling efficiency [31].
Figure 5: Correlation analysis between HOMO-LUMO overlap and key photophysical properties. The strong negative correlation with $\Delta E_{\text{ST}}$(left) validates the spatial separation model for TADF design, while the strong positive correlation with oscillator strength (right) highlights the fundamental design trade-off that molecular engineers must navigate.
III.6 Implications for reverse intersystem crossing ($k_{\text{RISC}}$)
While a small $\Delta E_{\text{ST}}$is the primary thermodynamic requirement for TADF, the overall efficiency is governed by the kinetics of RISC. While direct calculation of the $k_{\text{RISC}}$ rate is outside the scope of HTS, the derived structural principles must be explicitly connected to the kinetics of TADF. Within the Marcus theory framework, $k_{\text{RISC}}$ is proportional to the square of the spin-orbit coupling (SOC) matrix element and exponentially dependent on $\Delta E_{\text{ST}}$and the reorganization energy ($\lambda$) [32].
According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs. local-excitation, 1CT $\leftrightarrow$ 3LE) [26]. This creates a mechanistic challenge: the very act of creating a pure CT state in $S_{1}$ to minimize $\Delta E_{\text{ST}}$can lead to a $T_{1}$ state that also has pure CT character, making the transition spin-forbidden (1CT $\leftrightarrow$ 3CT) and suppressing $k_{\text{RISC}}$.
[…]
Table 2: Connection between molecular descriptors and the physical factors governing the RISC rate ($k_{\text{RISC}}$).
Molecular descriptor Influence on RISC factor Optimal range/strategy
HOMO-LUMO overlap Directly modulates $\Delta E_{\text{ST}}$ (thermodynamics) 0.15 to 0.300.150.30
[…]
Heteroatoms (S, Se) Directly increase SOC via heavy-atom effect Incorporate S-based donors
Conformational rigidity (e.g., in MR systems) Minimizes reorganization energy $\mathbf{\lambda}$ (kinetics) Use bridged or fused structures
III.7 Synthesis of a data-driven design strategy
The combination of these large-scale statistical analyses allows us to formulate a set of quantitative, data-driven design guidelines for accelerating the discovery of novel TADF emitters, summarized in Table 3. These rules provide a clear roadmap for molecular engineering, from selecting a high-performance architecture to fine-tuning its properties via geometric and electronic modifications.
Table 3: Quantitative TADF design guidelines derived from the high-throughput screening.
Parameter Target range / strategy Rationale
Architecture D-A-D or MR Statistically superior performance for general or blue emission, respectively.
[…]
HOMO-LUMO Overlap ($S^{\prime}_{\text{HL}}$) 0.15 to 0.300.150.30 Navigates the trade-off between $\Delta E_{\text{ST}}$  and oscillator strength.
Oscillator Strength ($f_{S_{1}}$) >0.1>0.1 Ensures a high radiative decay rate for bright emission.
III.8 Future outlook and data exploitation
III.8.1 The HTS dataset as a machine learning substrate
[…]
III.8.2 Actionable design targets
[…]
IV Conclusions
In this work, we performed a large-scale computational screening of $747$ TADF emitters to extract quantitative, statistically robust structure-property relationships for guiding molecular design. By systematically analyzing the impact of molecular architecture, conformational geometry, and electronic structure, we have moved beyond anecdotal evidence to establish a data-driven framework that integrates both the thermodynamic and kinetic principles of TADF engineering.
[…]
By providing a comprehensive set of design rules, a curated list of promising candidates, and a unified mechanistic view, this study significantly advances the goal of rational, predictive design for the next generation of TADF emitters. The experimental characterization of the identified candidates is the most important next step to validate our findings and realize the potential of this data-driven approach.
Conflicts of interest
There are no conflicts to declare.
V Associated content
V.1 Supporting information
The Supporting Information is available free of charge at https://pubs.acs.org/doi/...
Table 4: Molecular architecture and classification of the studied TADF emitters. This table provides the SMILES representation of each molecule, along with its structural classification.
Molecule Id SMILES Total donors Total acceptors Architecture Molecular weight logp
4CzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccccc32)c(C#N)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2ccccc21 4 4 Multi-D/A $788.914$ $13.818$
DMAC-DPS CC1(C)c2ccccc2N(c2ccc(S(=O)(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $632.829$ $10.738$
PSPCz C1=CC=C(C=C1)S(=O)(=O)C2=CC=CC(=C2)N3C4=CC=CC=C4C5=CC=CC=C53 1 0 1D-0A $383.472$ $5.617$
Px2BP C1=CC=C2C(=C1)N(C3=CC=CC=C3O2)C4=CC=C(C=C4)C(=O)C5=CC=C(C=C5)N6C7=CC=CC=C7OC8=CC=CC=C86 4 0 Multi-D/A $544.610$ $10.069$
CzS2 C1=CC=C2C(=C1)C3=CC=CC=C3N2C4=CC=C(C=C4)S(=O)(=O)C5=CC=C(C=C5)N6C7=CC=CC=C7C8=CC=CC=C86 2 0 2D-0A $548.667$ $8.714$
2TCz-DPS CC(C)(C)C1=CC2=C(C=C1)N(C3=C2C=C(C=C3)C(C)(C)C)C4=CC=C(C=C4)S(=O)(=O)C5=CC=C(C=C5)N6C7=C(C=C(C=C7)C(C)(C)C)C8=C6C=CC(=C8)C(C)(C)C 2 0 2D-0A $773.099$ $13.904$
TDBA-DI CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(-n4c5ccccc5c5c6c(c7ccccc7n6-c6ccccc6)c6c(c7ccccc7n6-c6ccccc6)c54)cc(c31)O2 3 0 Multi-D/A $877.897$ $14.300$
OPDPO O=CC1=CCC(c2ccccc2N2c3ccccc3Sc3ccccc32)(P(=O)(c2ccccc2)c2ccccc2)C=C1 2 0 2D-0A $581.677$ $8.916$
2CzTPEPCz C(=C(c1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 3 0 Multi-D/A $828.031$ $16.234$
TBRb CC(C)(C)c1ccc(-c2c3ccc(C(C)(C)C)cc3c(-c3ccccc3)c3c(-c4ccc(C(C)(C)C)cc4)c4ccc(C(C)(C)C)cc4c(-c4ccccc4)c23)cc1 0 0 Unknown $757.118$ $17.004$
TBPe CC(C)(C)c1cc2cc(C(C)(C)C)cc3c4cc(C(C)(C)C)cc5cc(C(C)(C)C)cc(c(c1)c23)c54 0 0 Unknown $476.748$ $10.927$
2CzPN N#Cc1cc(-c2cccc3c2[nH]c2ccccc23)c(-c2cccc3c2[nH]c2ccccc23)cc1C#N 2 4 Multi-D/A $458.524$ $8.033$
3CzFTFP Fc1nc(F)c(F)c(C2(c3ccc4c(c3)c3ccccc3n4-c3ccccc3)c3ccccc3-c3ccccc32)c1F 1 0 1D-0A $556.562$ $9.098$
DDPhCz-DCPP N#Cc1nc2c3ccc(-n4c5ccc(-c6ccccc6)cc5c5cc(-c6ccccc6)ccc54)cc3c3cc(-n4c5ccc(-c6ccccc6)cc5c5cc(-c6ccccc6)ccc54)ccc3c2nc1C#N 2 3 Multi-D/A $915.072$ $16.542$
ACRXTN CC1(C)c2ccccc2N(c2ccc3c(=O)c4ccccc4oc3c2)c2ccccc21 1 0 1D-0A $403.481$ $7.055$
TXO-TPA O=C1c2ccccc2S(=O)(=O)c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)cc21 1 0 1D-0A $487.580$ $7.201$
PCTXO O=C1c2ccccc2S(=O)(=O)c2ccc(-c3ccc(N(c4ccc(Br)cc4)c4ccc(Br)cc4)cc3)cc21 1 0 1D-0A $645.372$ $8.726$
IP-PPI c1ccc(-n2c(-c3ccc(-c4ccc(-c5cn6ccccc6n5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $562.676$ $9.981$
DMDHNP-DPS CC1(C)c2cnccc2Nc2ccncc21.O=S(=O)(c1ccccc1)c1ccccc1 0 0 Unknown $429.545$ $5.379$
4NTAZ-PPI c1ccc(-c2nnc(-c3ccc(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)cc3)n2-c2cccc3ccccc23)cc1 0 0 Unknown $715.860$ $12.734$
TAZ-PPI c1ccc(-c2nnc(-c3ccc(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)cc3)n2-c2ccccc2)cc1 0 0 Unknown $665.800$ $11.581$
PTZ-mPYR Cc1cc(N2c3ccccc3Sc3ccccc32)ccc1-c1cc(-c2ccc(N3c4ccccc4Sc4ccccc43)cc2C)ncn1 4 1 Multi-D/A $654.864$ $12.296$
pCNBCzoCF3 N#Cc1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccc(C#N)cc4C(F)(F)F)ccc32)c(C(F)(F)F)c1 2 4 Multi-D/A $670.616$ $11.329$
PTZTPE-2 CC1(C)OB(c2ccc(C(=C(c3ccccc3)c3ccccc3)c3ccccc3)cc2)OC1(C)C 0 0 Unknown $458.410$ $6.993$
4CzIPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(C#N)c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2ccccc21 4 4 Multi-D/A $788.914$ $13.818$
DMAC-TRZ CC1(C)c2ccccc2N(c2ccc(-c3nc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)c2ccccc21 1 1 D-A $516.648$ $8.982$
TCzTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ccccc54)c3)n2)cc1 3 1 Multi-D/A $804.957$ $14.164$
$\delta$-2CbPN N#Cc1cc(-n2c3ccccc3c3ncccc32)c(-n2c3ccccc3c3ncccc32)cc1C#N 0 4 Multi-D/A $460.500$ $6.414$
CzDCbTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ncccc54)c(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ncccc54)c3)n2)cc1 1 1 D-A $806.933$ $12.954$
$\alpha$-2CbPN N#Cc1cc(-n2c3ccccc3c3cccnc32)c(-n2c3ccccc3c3cccnc32)cc1C#N 0 4 Multi-D/A $460.500$ $6.414$
CN-QP N#Cc1ccc2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2c1 4 3 Multi-D/A $669.744$ $11.986$
TFM-QP FC(F)(F)c1ccc2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2c1 4 1 Multi-D/A $712.731$ $13.134$
NDN c1ccc2c(c1)c1cccnc1n2-c1ccc2sc3ccc(-n4c5ccccc5c5cccnc54)cc3c2c1 0 0 Unknown $516.629$ $9.039$
mCP c1cc(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $408.504$ $7.881$
ZDZ c1ccc2c(c1)c1ccccc1n2-c1ccc2sc3ccc(-n4c5ccccc5c5ccccc54)cc3c2c1 2 0 2D-0A $514.653$ $10.249$
ZDN c1ccc2c(c1)c1ccccc1n2-c1ccc2sc3ccc(-n4c5ccccc5c5cccnc54)cc3c2c1 1 0 1D-0A $515.641$ $9.644$
BTCz3CzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccccc32)c(C#N)c(-n2c3ccccc3c3c4sc5ccccc5c4ccc32)c1-n1c2ccccc2c2ccccc21 4 4 Multi-D/A $895.063$ $16.186$
2BTCz2CzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3c4sc5ccccc5c4ccc32)c(C#N)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2c3sc4ccccc4c3ccc21 4 4 Multi-D/A 1 001.2121\,001.212 $18.554$
APDC-DTPA N#Cc1nc2c(nc1C#N)-c1ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc-2c13 2 2 2D-2A $740.870$ $13.294$
BMIM N#Cc1nc2c(nc1C#N)-c1ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc-2c13 2 2 2D-2A $740.870$ $13.294$
ADO-TPA O=C1C(=O)c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3cccc1c23 1 0 1D-0A $425.487$ $7.356$
DCzTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)n2)cc1 2 1 D-A-D $639.762$ $11.067$
BDPCC c1ccc(-c2ccc3[nH]c4c(-c5ccc6c([nH]c7ccccc76)c5-c5cc(-c6ccccc6)cc6c5[nH]c5ccc(-c7ccccc7)cc56)cc(-c5ccccc5)cc4c3c2)cc1 3 0 Multi-D/A $801.993$ $16.592$
SpiroAC-TRZ c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(N4c5ccccc5C5(c6ccccc6-c6ccccc65)c5ccccc54)cc3)n2)cc1 1 1 D-A $638.774$ $11.019$
DBT-BZ-DMAC CC1(C)c2ccccc2Nc2cccc(C(=O)c3ccccc3)c21.c1ccc2c(c1)sc1ccccc12 0 0 Unknown $497.663$ $9.355$
DQP c1ccc2nc3c(nc2c1)c1nc2ccccc2nc1c1nc2ccccc2nc31 0 6 Multi-D/A $384.402$ $4.976$
TPA-DCPP N#Cc1nc2c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3c3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc3c2nc1C#N 2 3 Multi-D/A $766.908$ $13.953$
2PXZ-OXD c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)o2)cc1 4 0 Multi-D/A $584.635$ $10.555$
BmPyPhB CC1(C)c2ccccc2N(c2ccc(-c3nc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)nc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)n3)cc2)c2ccccc21 3 1 Multi-D/A $931.200$ $17.200$
MPPA COc1ccc(N(c2ccc(OC)cc2)c2ccc(-c3ccc4c(c3)C(=O)c3ccc(-c5ccc(N(c6ccc(OC)cc6)c6ccc(OC)cc6)cc5)cc3C4=O)cc2)cc1 2 0 2D-0A $814.938$ $12.770$
ACRFLCN N#Cc1ccc2c(c1)C1(c3cc(C#N)ccc3-2)c2ccccc2N(c2ccccc2)c2ccccc21 1 4 Multi-D/A $457.536$ $7.576$
PTZ-DCPP N#Cc1nc2c3ccc(N4c5ccccc5Sc5ccccc54)cc3c3cc(N4c5ccccc5Sc5ccccc54)ccc3c2nc1C#N 4 3 Multi-D/A $674.814$ $11.548$
BBA-Pyr c1ccc(-c2ccc(N(c3ccc(-c4ccccc4)cc3)c3ccc(-c4ccc(-c5cc(-c6ccccc6)nc(-c6ccccc6)n5)cc4)cc3)cc2)cc1 1 1 D-A $703.889$ $13.948$
PhICz-Pyr c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)cc4c5cc(-c6ccc(-c7cc(-c8ccccc8)nc(-c8ccccc8)n7)cc6)ccc5n3c24)cc1 2 1 D-A-D $699.857$ $13.629$
PhCz-Pyr c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)ccc2n3-c2ccc(-c3ccc(-c4cc(-c5ccccc5)nc(-c5ccccc5)n4)cc3)cc2)cc1 1 1 D-A $701.873$ $13.576$
PDA-DP CC1(C)c2cc(-c3ccccc3)ccc2N(c2ccc(S(=O)(=O)c3ccc(N4c5ccc(-c6ccccc6)cc5C(C)(C)c5cc(-c6ccccc6)ccc54)cc3)cc2)c2ccc(-c3ccccc3)cc21 2 0 2D-0A $937.221$ $17.406$
9CzFDBFDPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cc(-c3ccc(C4(c5ccc(-n6c7ccccc7c7ccccc76)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A 1 050.1471\,050.147 $15.992$
9PhCzFDBFDPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cc(-c3ccc(C4(c5ccc(-c6ccc(-n7c8ccccc8c8ccccc87)cc6)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A 1 126.2451\,126.245 $17.659$
9PhCzFDBFSPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1ccc(-c3ccc(C4(c5ccc(-c6ccc(-n7c8ccccc8c8ccccc87)cc6)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A $926.068$ $16.020$
9CzFDBFSPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1ccc(-c3ccc(C4(c5ccc(-n6c7ccccc7c7ccccc76)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A $849.970$ $14.353$
9ArFDBFxPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cc(-c3ccc(C4(c5ccc(-n6c7ccccc7c7ccccc76)cc5)c5ccccc5-c5ccccc54)cc3)cc12 1 0 1D-0A 1 050.1471\,050.147 $15.992$
12BTCzTPN N#Cc1c(-n2c3ccccc3c3ccccc32)c(-n2c3ccccc3c3ccc4c5ccccc5sc4c32)c(C#N)c(-n2c3ccccc3c3ccccc32)c1-n1c2ccccc2c2ccc3c4ccccc4sc3c21 4 4 Multi-D/A 1 001.2121\,001.212 $18.554$
oDTBPZ-DPXZ Cc1ccccc1-c1cc2nc3c4ccc(N5c6ccccc6Oc6ccccc65)cc4c4cc(N5c6ccccc6Oc6ccccc65)ccc4c3nc2cc1-c1ccccc1C 4 2 Multi-D/A $822.968$ $16.191$
3DMAC-BP-CN CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2c2nc4cc(N5c6ccccc6C(C)(C)c6ccccc65)c(C#N)cc4nc32)c2ccccc21 3 4 Multi-D/A $927.168$ $17.288$
3DMAC-BP-Br CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2c2nc4cc(N5c6ccccc6C(C)(C)c6ccccc65)c(Br)cc4nc32)c2ccccc21 3 2 Multi-D/A $981.054$ $18.179$
TPA-PZCN N#Cc1ccc2c(c1)c1cc(C#N)ccc1c1nc3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3nc21 2 6 Multi-D/A $816.968$ $15.106$
1PXZ-BP c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)c1ccccc1c1nc3ccccc3nc21 2 2 2D-2A $461.524$ $8.665$
PTZ-DBTO2 O=S1(=O)c2ccccc2-c2cc(N3c4ccccc4Sc4ccccc43)ccc21 2 0 2D-0A $413.523$ $6.434$
BACQ CC1(C)c2c(ccc3ccccc23)-n2c3ccccc3c(=O)c3cccc1c32 1 0 1D-0A $361.444$ $5.936$
2Cz-DMAC-TXO CC1(C)c2cc(-n3c4ccccc4c4ccccc43)ccc2N(c2ccc3sc4ccccc4c(=O)c3c2)c2ccc(-n3c4ccccc4c4ccccc43)cc21 3 0 Multi-D/A $749.939$ $13.718$
2Cz-DMAC-TTR CC1(C)c2cc(-n3c4ccccc4c4ccccc43)ccc2N(c2ccc3c(c2)S(=O)(=O)c2ccccc2S3(=O)=O)c2ccc(-n3c4ccccc4c4ccccc43)cc21 3 0 Multi-D/A $817.992$ $11.969$
2Cz-DMAC-BTB CC1(C)c2cc(-n3c4ccccc4c4ccccc43)ccc2N(c2ccc3c(c2)-c2ccccc2S3(=O)=O)c2ccc(-n3c4ccccc4c4ccccc43)cc21 3 0 Multi-D/A $753.927$ $12.803$
DMAC-BPP CC1(C)c2ccccc2N(c2cc(-c3cc(-c4ccc(-c5ccccc5)cc4)nc(-c4ccccc4)n3)cc(N3c4ccccc4C(C)(C)c4ccccc43)c2)c2ccccc21 2 1 D-A-D $799.034$ $15.363$
Me-MOC COc1ccc2c(c1)c1ccccc1n2-c1ccc(-c2c(C#N)c(C)nc(C)c2C#N)cc1 1 2 A-D-A $428.495$ $6.215$
Me-DOC COc1ccc2c(c1)c1cc(OC)ccc1n2-c1ccc(-c2c(C#N)c(C)nc(C)c2C#N)cc1 1 2 A-D-A $458.521$ $6.223$
3DCPO O=P(c1ccccc1)(c1ccc2c(c1)c1ccccc1n2-c1ccccc1)c1ccc2c(c1)c1ccccc1n2-c1ccccc1 2 0 2D-0A $608.681$ $9.520$
TDBA-TPDICz CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(-n4c5ccccc5c5c(-c6ccccc6)c6c(c(-c7ccccc7)c54)c4ccccc4n6-c4ccccc4)cc(c31)O2 2 0 2D-0A $864.898$ $14.537$
TRZ-pIC CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(-n4c5ccccc5c5c(-c6ccccc6)c6c(c(-c7ccccc7)c54)c4ccccc4n6-c4ccccc4)cc(c31)O2 2 0 2D-0A $864.898$ $14.537$
SBF-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)C1(c3ccccc3-c3ccccc31)c1cc(N3c4ccccc4Oc4ccccc43)ccc1-2 4 0 Multi-D/A $678.791$ $13.181$
PtOEP CCC1=C($C$)c2cc3[nH]c(cc4nc(cc5[nH]c(cc1n2)c($C$)c5CC)C(CC)=C4CC)c($C$)c3CC.[Pt+2] 0 0 Unknown $729.870$ $10.023$
PxCYP N#Cc1ccc(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)nc1-c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 1 Multi-D/A $618.696$ $11.438$
AcCYP CC1(C)c2ccccc2N(c2ccc(-c3ccc(C#N)c(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)n3)cc2)c2ccccc21 2 1 D-A-D $670.860$ $12.505$
BCN CC1(C)c2ccccc2N2c3ccc4cc(C#N)ccc4c3C(C)(C)c3cccc1c32 1 2 A-D-A $400.525$ $7.460$
DPPSPF O=S(=O)(c1ccccc1)c1ccc(-c2ccc3c(c2)C(c2ccccc2)(c2ccccc2)c2ccccc2-3)cc1 0 0 Unknown $534.680$ $8.550$
Me2ACBO CC(C)(c1ccc2c(c1)B1c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3Oc3cccc(c31)O2)c1ccc2c(c1)B1c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3Oc3cccc(c31)O2 2 0 2D-0A $994.809$ $13.686$
F2AcBO CC1(C)c2ccccc2N(c2ccc3c(c2)Oc2cccc4c2B3c2cc(C3(c5ccc6c(c5)B5c7ccc(N8c9ccccc9C(C)(C)c9ccccc98)cc7Oc7cccc(c75)O6)c5ccccc5-c5ccccc53)ccc2O4)c2ccccc21 2 0 2D-0A 1 116.9351\,116.935 $15.723$
PSPSF O=S(=O)(c1ccccc1)c1ccc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc1 0 0 Unknown $532.664$ $8.530$
TXO-P-Si O=C1c2ccccc2S(=O)(=O)c2ccc(-c3ccc([Si](c4ccccc4)(c4ccccc4)c4ccccc4)cc3)cc21 0 0 Unknown $578.765$ $5.108$
CzB-FMPIM FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc(-c3ccccc3)c2-c2ccccc2)cc1 1 0 1D-0A $681.761$ $12.656$
CzB-FMPPI FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 0 1D-0A $679.745$ $12.782$
BACF CC1(C)c2ccccc2N(c2ccccc2)c2ccc3cc(F)ccc3c21 1 0 1D-0A $353.440$ $7.088$
BACH CC1(C)c2ccccc2N(c2ccccc2)c2ccc3ccccc3c21 1 0 1D-0A $335.450$ $6.949$
PzTDBA CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(N4c5ccccc5N(c5cc6c7c(c5)Oc5ccc(C(C)(C)C)cc5B7c5cc(C(C)(C)C)ccc5O6)c5ccccc54)cc(c31)O2 2 0 2D-0A $942.818$ $13.585$
3TCPM O=C(c1ccccc1)c1ccc(-n2c3ccccc3c3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc32)cc1 2 0 2D-0A $590.726$ $11.151$
TPA-PPO c1ccc(-c2nc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cn3)oc2-c2ccccc2)cc1 1 0 1D-0A $541.654$ $10.207$
CzPh-PPO c1ccc(-c2nc(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cn3)oc2-c2ccccc2)cc1 1 0 1D-0A $539.638$ $9.835$
m-CF3PIBI FC(F)(F)c1cccc(-n2c(-c3ccc(-c4ccc(-c5ccc(-c6nc7ccccc7n6-c6ccccc6)cc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)c1 0 0 Unknown $782.869$ $14.358$
DPTPEA c1ccc(C(=C(c2ccccc2)c2ccc(N(c3ccccc3)c3ccccc3)cc2)c2ccccc2)cc1 1 0 1D-0A $499.657$ $10.164$
PICNAnCz N#Cc1ccc(-n2c(-c3ccc(-c4c5ccccc5c(Br)c5ccccc45)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 2 0D-2A $650.579$ $11.606$
DPA-DPQ c1ccc(-c2cc3ncc(N(c4ccccc4)c4ccccc4)nc3cc2-c2ccccc2)cc1 0 1 0D-1A $449.557$ $8.434$
PhQLPXZ c1ccc(-c2cc(-c3ccccc3)c3cc(N4c5ccccc5Oc5ccccc54)ccc3n2)cc1 2 0 2D-0A $462.552$ $9.144$
2DPA-DPQ c1ccc(-c2cc3nc(N(c4ccccc4)c4ccccc4)c(N(c4ccccc4)c4ccccc4)nc3cc2-c2ccccc2)cc1 0 1 0D-1A $616.768$ $11.903$
FDQPXZ Fc1cc2nc(-c3ccccc3)c(-c3ccccc3)nc2cc1N1c2ccccc2Oc2ccccc21 2 1 D-A-D $481.530$ $8.678$
DTRZSFX c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc4c(c3)C3(c5ccccc5Oc5ccccc53)c3cc(-c5nc(-c6ccccc6)nc(-c6ccccc6)n5)ccc3-4)n2)cc1 0 2 0D-2A $794.918$ $12.527$
E-o-DHMPA OCc1ccccc1-c1c2ccccc2c(-c2ccccc2CO)c2ccccc12 0 0 Unknown $390.482$ $6.312$
Z-o-DHMPA OCc1ccccc1-c1c2ccccc2c(-c2ccccc2CO)c2ccccc12 0 0 Unknown $390.482$ $6.312$
p-DHMPA OCc1ccc(-c2c3ccccc3c(-c3ccc(CO)cc3)c3ccccc23)cc1 0 0 Unknown $390.482$ $6.312$
STAC-BP-STAC O=C(c1ccc(N2c3ccccc3C3(c4ccccc4Sc4ccccc43)c3ccccc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4Sc4ccccc43)c3ccccc32)cc1 2 0 2D-0A $905.160$ $16.179$
SXAC-BP-SXAC O=C(c1ccc(N2c3ccccc3C3(c4ccccc4Oc4ccccc43)c3ccccc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4Oc4ccccc43)c3ccccc32)cc1 2 0 2D-0A $873.024$ $15.461$
m-DHMPA OCc1cccc(-c2c3ccccc3c(-c3cccc(CO)c3)c3ccccc23)c1 0 0 Unknown $390.482$ $6.312$
SFAC-BP-SFAC O=C(c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3ccccc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3ccccc32)cc1 2 0 2D-0A $841.026$ $15.210$
PBTPA CCN1c2ccccc2Sc2cc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c4nsnc34)ccc21 2 0 2D-0A $604.804$ $11.118$
HDP O=C1c2ccccc2C(=O)c2nc3c(cc(O)c4ccccc43)nc21 0 1 0D-1A $326.311$ $3.264$
PY-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4nc5c6ccc(C#N)cc6c6cc(C#N)ccc6c5nc4cc3-c3cccnc3)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 6 Multi-D/A $762.961$ $13.627$
CN-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4nc5c6ccc(C#N)cc6c6cc(C#N)ccc6c5nc4cc3C#N)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 8 Multi-D/A $710.885$ $12.436$
DMIC-TRZ CC1(C)CC(c2cccc(-c3nc(-c4ccccc4)nc(-c4ccccc4)n3)c2)C=c2[nH]c3cc4c(cc3c21)-c1ccccc1C=4 0 1 0D-1A $592.746$ $8.409$
NyDPAc CC1(C)c2ccccc2Nc2cccc(-c3ccc(-c4ccc5nc(-c6ccc(-c7cccc8c7C(C)(C)c7ccccc7N8)cc6)ccc5n4)cc3)c21 0 0 Unknown $696.898$ $13.064$
oCBP c1ccc(-n2c3ccccc3c3ccccc32)c(-c2ccccc2-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $484.602$ $9.548$
PABP Cc1cc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)ccc1-c1ccc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)cc1C 0 0 Unknown $686.898$ $15.251$
DHID-DPS CCn1c2ccccc2c2c1c1ccccc1n2-c1ccc(S(=O)(=O)c2ccc(-n3c4ccccc4c4c3c3ccccc3n4CC)cc2)cc1 0 0 Unknown $682.849$ $10.663$
4FBTPI CC(C)(C)c1ccc(-n2c(-c3ccc(-c4c($F)c($-c5ccc(-c6nc7c8ccccc8c8ccccc8c7n6-c6ccc(C(C)(C)C)cc6)cc5)c(F)c4F)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $999.168$ $18.797$
BCzBMes Cc1cc(C)c(B(c2c(C)cc(C)cc2C)n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)c(C)c1 2 0 2D-0A $656.682$ $11.063$
CCDMB Cc1cc(C)c(B(c2ccc(-n3c4ccccc4c4cc(-n5c6ccccc6c6ccccc65)ccc43)cc2)c2c(C)cc(C)cc2C)c(C)c1 2 0 2D-0A $656.682$ $10.248$
PCDMB Cc1cc(C)c(B(c2ccc(-n3c4ccccc4c4cc(N5c6ccccc6Oc6ccccc65)ccc43)cc2)c2c(C)cc(C)cc2C)c(C)c1 3 0 Multi-D/A $672.681$ $10.726$
BTrztCz CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccccc1-c1nc(-c2ccccc2)nc(-c2ccc(-c3nc(-c4ccccc4)nc(-c4ccccc4-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)n3)cc2)n1 2 2 2D-2A 1 095.4521\,095.452 $19.443$
DHID-DBS CCCCn1c2ccccc2c2c1c1ccccc1n2-c1ccc2c(c1)-c1cc(-n3c4ccccc4c4c3c3ccccc3n4CCCC)ccc1S2(=O)=O 0 0 Unknown $736.941$ $12.204$
CZ-DPS-DMAC CC1(C)c2ccccc2N(c2ccc(S(=O)(=O)c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)cc2)c2ccccc21 3 0 Multi-D/A $755.943$ $12.823$
XT-T Cc1ccc(N(c2ccc(C)cc2)c2ccccc2C2(c3ccccc3N(c3ccc(C)cc3)c3ccc(C)cc3)CCCCC2)cc1 2 0 2D-0A $626.888$ $13.110$
3DPTIA CC1(C)c2ccccc2N(c2nc(-c3ccccc3)nc(-c3ccccc3)n2)c2cc3c4ccccc4n(-c4ccccc4)c3cc21 1 1 D-A $605.745$ $10.412$
PyIAnTPh c1ccc(-c2cc(-c3ccccc3)cc(-c3c4ccccc4c(-c4ccc(-c5nc6c7cccc8ccc9cccc(c9c87)c6n5-c5ccccc5)cc4)c4ccccc34)c2)cc1 0 0 Unknown $798.989$ $16.564$
2CzP-INN N#Cc1c(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cncc1-c1ccc(-n2c3ccccc3c3ccccc32)cc1 2 1 D-A-D $586.698$ $10.481$
VBN0 C=Cc1ccc(COc2ccc3c(c2)c2ccccc2n3-c2cccc(-n3c4ccccc4c4cc(OCc5ccc(C=C)cc5)ccc43)c2C#N)cc1 2 2 2D-2A $697.838$ $12.196$
PIAnTPh c1ccc(-c2cc(-c3ccccc3)cc(-c3c4ccccc4c(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)c4ccccc34)c2)cc1 0 0 Unknown $774.967$ $15.973$
BCZ-DPS-AD CC1(C)c2ccccc2N(c2cc(N3c4ccccc4C(C)(C)c4ccccc43)cc(S(=O)(=O)c3ccc4c(c3)c3ccccc3n4-c3ccccc3)c2)c2ccccc21 3 0 Multi-D/A $798.024$ $13.835$
2BTPA-INN N#Cc1c(-c2ccc(N(c3ccc(-c4ccccc4)cc3)c3ccc(-c4ccccc4)cc3)cc2)cncc1-c1ccc(N(c2ccc(-c3ccccc3)cc2)c2ccc(-c3ccccc3)cc2)cc1 2 1 D-A-D $895.122$ $17.895$
SpiroAc c1ccc2c(c1)Nc1ccccc1C21c2ccccc2-c2ccccc21 0 0 Unknown $331.418$ $6.107$
BuPCzBCO Cc1ccc2c3cc(Br)c(C)cc3n(-c3ccc(C(C)(C)C)cc3)c2c1 1 0 1D-0A $406.367$ $7.461$
AQ-S-CZ O=C1c2ccccc2C(=O)c2cc(Sc3ccc(-n4c5ccccc5c5ccccc54)cc3)ccc21 1 0 1D-0A $481.576$ $7.710$
DBP-3TPA c1ccc(N(c2ccccc2)c2ccc(-c3ccc4nc5c6ccc(-c7ccc(N(c8ccccc8)c8ccccc8)cc7)cc6c6cc(-c7ccc(N(c8ccccc8)c8ccccc8)cc7)ccc6c5nc4c3)cc2)cc1 3 2 Multi-D/A 1 010.2571\,010.257 $20.500$
DCzIPN N#Cc1cc(C#N)c(-n2c3ccccc3c3ccccc32)cc1-n1c2ccccc2c2ccccc21 2 4 Multi-D/A $458.524$ $7.624$
DMeCzIPN Cc1cccc2c3ccccc3n(-c3cc(-n4c5ccccc5c5cccc(C)c54)c(C#N)cc3C#N)c12 2 4 Multi-D/A $486.578$ $8.241$
SAFDPA c1ccc(N(c2ccccc2)c2ccc3c(c2)-c2ccccc2C32c3ccccc3N(c3ccccc3)c3ccccc32)cc1 2 0 2D-0A $574.727$ $11.303$
t-DABNA CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cccc2c43)cc1 2 0 2D-0A $644.756$ $10.959$
3-AcBPtCN CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(C#N)ccc2c2nc4cc(C#N)c(C#N)cc4nc32)c2ccccc21 1 8 Multi-D/A $562.636$ $8.814$
3-AcBPdCN CC1(C)c2ccccc2N(c2ccc3c(c2)c2ccccc2c2nc4cc(C#N)c(C#N)cc4nc32)c2ccccc21 1 6 Multi-D/A $537.626$ $8.942$
11-AcBPdCN CC1(C)c2ccccc2N(c2ccc3nc4c5ccc(C#N)cc5c5cc(C#N)ccc5c4nc3c2)c2ccccc21 1 6 Multi-D/A $537.626$ $8.942$
PID-3 c1ccc(-n2c(-c3ccc(-c4cccs4)s3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $458.611$ $8.789$
3-ACBPmCN CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(C#N)ccc2c2nc4ccccc4nc32)c2ccccc21 1 4 Multi-D/A $512.616$ $9.070$
3-DTPA-BBI O=c1c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc(c32)c2nc3ccccc3n12 2 0 2D-0A $756.909$ $13.865$
IABTCN CC1(C)c2ccccc2-c2cc3c(cc21)N(c1ccccc1)c1ccc(-c2ccc(C#N)c4nsnc24)cc1C3(c1ccccc1)c1ccccc1 1 2 A-D-A $684.868$ $11.702$
BTZPP CC(C)(C)c1ccc(-n2c(-c3ccc(-c4cccc5nsnc45)cc3)cc3c2cc(-c2ccc(-c4cccc5nsnc45)cc2)n3-c2ccc(C(C)(C)C)cc2)cc1 0 0 Unknown $791.062$ $13.694$
[…]
BCzPh c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)cc1 2 0 2D-0A $484.602$ $9.548$
TBPM2CN N#Cc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(C#N)cc4)ccc2c2c3nc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)n2-c2ccccc2)cc1 1 4 Multi-D/A $815.980$ $15.213$
TPM2CN N#Cc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(C#N)cc4)ccc2c2c3nc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)n2-c2ccccc2)cc1 1 4 Multi-D/A $739.882$ $13.546$
PT-BZ-DMAC CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(Sc4ccccc4)cc3)cc2)c2ccccc21 1 0 1D-0A $497.663$ $9.178$
PS-BZ-DMAC CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(S(=O)(=O)c4ccccc4)cc3)cc2)c2ccccc21 1 0 1D-0A $529.661$ $7.860$
2PCzBN-tPh CC(C)(C)c1ccc(-c2cc(-n3c4ccc(-c5ccccc5)cc4c4cc(-c5ccccc5)ccc43)c(C#N)c(-n3c4ccc(-c5ccccc5)cc4c4cc(-c5ccccc5)ccc43)c2)cc1 2 2 2D-2A $870.112$ $17.385$
2PCzBN N#Cc1c(-n2c3ccc(-c4ccccc4)cc3c3cc(-c4ccccc4)ccc32)cccc1-n1c2ccc(-c3ccccc3)cc2c2cc(-c3ccccc3)ccc21 2 2 2D-2A $737.906$ $14.420$
TPA-2 O=C1c2ccccc2C(=O)c2c1ccc1nc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)nc21 2 1 D-A-D $746.870$ $12.679$
36DTPAFM N#CC(C#N)=C1c2ccc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)cc2-c2cc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)ccc21 2 2 2D-2A $714.872$ $13.790$
AQTC N#Cc1cc2nc3c(nc2c(C#N)c1C#N)-c1cccc2cccc-3c12 0 7 Multi-D/A $329.322$ $4.045$
3TPAFM N#CC(C#N)=C1c2ccccc2-c2cc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)ccc21 1 2 A-D-A $471.563$ $8.653$
BBPA CCCCOc1ccccc1-c1c2ccccc2c(-c2ccccc2OCCCC)c2ccccc12 0 0 Unknown $474.644$ $9.685$
DMPA COc1ccc(-c2c3ccccc3c(-c3ccc(OC)cc3OC)c3ccccc23)c(OC)c1 0 0 Unknown $450.534$ $7.361$
PXZ-NAI O=C1c2cccc3c(N4c5ccccc5Oc5ccccc54)ccc(c23)C(=O)N1c1ccccc1 2 0 2D-0A $454.485$ $7.216$
PTZ-NAI O=C1c2cccc3c(N4c5ccccc5Sc5ccccc54)ccc(c23)C(=O)N1c1ccccc1 2 0 2D-0A $470.553$ $7.575$
AI-4Cz O=C1c2c(c(-n3c4ccccc4c4ccccc43)c(-n3c4ccccc4c4ccccc43)c(-n3c4ccccc4c4ccccc43)c2-n2c3ccccc3c3ccccc32)C(=O)N1c1ccccc1 4 0 Multi-D/A $884.011$ $14.876$
DAcIPN CC1(C)c2ccccc2N(c2cc(N3c4ccccc4C(C)(C)c4ccccc43)c(C#N)cc2C#N)c2ccccc21 2 4 Multi-D/A $542.686$ $9.648$
TCz1 CCCCC($C$)Cn1c2ccc(-n3c4ccccc4c4ccccc43)cc2c2cc(-n3c4ccccc4c4ccccc43)ccc21 3 0 Multi-D/A $609.817$ $12.205$
MBF F[B-]1(F)OC(c2ccccc2)=Cc2cc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc[n+]21 2 1 D-A-D $503.317$ $7.363$
MPA O=C(Cc1cc(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)ncn1)c1ccccc1 2 1 D-A-D $455.517$ $7.145$
3,3-NHC O=c1oc2ccccc2c(O)c1C(c1ccc2ccccc2c1)c1c(O)c2ccccc2oc1=O 0 0 Unknown $462.457$ $5.644$
3CP-DPS-DMAC CC1(C)c2ccccc2N(c2ccc(S(=O)(=O)c3ccc4c(c3)c3ccccc3n4-c3ccccc3)cc2)c2ccccc21 2 0 2D-0A $590.748$ $9.726$
3CP-DPS-PXZ O=S(=O)(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc2c(c1)c1ccccc1n2-c1ccccc1 3 0 Multi-D/A $564.666$ $9.192$
3CzAnN c1ccc(-n2c3ccccc3c3cc(-c4c5ccccc5c(-c5ccc6ccccc6c5)c5ccccc45)ccc32)cc1 1 0 1D-0A $545.685$ $11.577$
DPQ-DMAC CC1(C)c2ccccc2N(c2ccc3nc(-c4ccccn4)c(-c4ccccn4)nc3c2)c2ccccc21 1 1 D-A $491.598$ $7.863$
DPAA c1ccc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)cc1 0 0 Unknown $330.430$ $7.327$
t3Cz-SO CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc2c(c1)c1cc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)ccc1n2-c1ccc(S(=O)(=O)c2ccccc2)cc1 3 0 Multi-D/A $938.294$ $17.001$
Cz-SO O=S(=O)(c1ccccc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 0 1D-0A $383.472$ $5.617$
DPQ-PXZ c1ccc(-c2nc3ccc(N4c5ccccc5Oc5ccccc54)cc3nc2-c2ccccn2)nc1 2 1 D-A-D $465.516$ $7.329$
pCzAnN c1ccc2cc(-c3c4ccccc4c(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)c4ccccc34)ccc2c1 1 0 1D-0A $545.685$ $11.577$
3-BHH O=c1oc2ccccc2c(O)c1C(c1ccc(Br)s1)c1c(O)c2ccccc2oc1=O 0 0 Unknown $497.322$ $5.315$
DPP-PXZ c1ccc(-c2nc3cc(N4c5ccccc5Oc5ccccc54)cnc3nc2-c2ccccn2)nc1 1 0 1D-0A $466.504$ $6.724$
APPT-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c3c(cccc13)-c1nc3c4cccnc4c4ncccc4c3nc1-2 2 1 D-A-D $537.582$ $9.102$
HITPE Oc1ccc(-c2ccc(C(=C(c3ccccc3)c3ccccc3)c3ccccc3)cc2)cc1-c1nc(-c2ccccc2)c(-c2ccccc2)n1-c1ccccc1 0 0 Unknown $718.900$ $13.253$
DV-3CzCN C=Cc1ccc(COc2cccc3c2c2ccccc2n3-c2cc(-n3c4ccccc4c4c(OCc5ccc(C=C)cc5)cccc43)c(C#N)c(-n3c4ccccc4c4c(OCc5ccc(C=C)cc5)cccc43)c2)cc1 3 2 Multi-D/A $995.195$ $17.516$
SF-DPSO O=S(=O)(c1ccc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc1)c1ccc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc1 0 0 Unknown $847.051$ $14.540$
BDQ-tBuCz CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c2nc(-c3ccccc3)c(-c3ccccc3)nc12 2 1 D-A-D $837.168$ $16.348$
BDQDMAC CC1(C)c2ccccc2N(c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)c3nc(-c4ccccc4)c(-c4ccccc4)nc23)c2ccccc21 2 1 D-A-D $696.898$ $13.182$
BDQPCz c1ccc(-c2nc3c(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)c3nc2-c2ccccc2)cc1 2 1 D-A-D $764.932$ $14.492$
4-SCZ c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2c1ccc1[nH]c3ccccc3c21 1 0 1D-0A $405.500$ $7.665$
Py-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4nc5c6ccc(-c7ccncc7)nc6c6nc(-c7ccncc7)ccc6c5nc4cc3C#N)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 4 Multi-D/A $817.013$ $13.607$
DPACPhTPI CC(C)(C)c1ccc(-n2c(-c3ccc(-c4ccc(N5c6ccccc6C(c6ccccc6)(c6ccccc6)c6ccccc65)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 0 1D-0A $834.079$ $16.129$
oPTBC N#Cc1ccc(-c2cc(C#N)cc(-c3ccc(C#N)cc3)c2N2c3ccccc3Oc3ccccc32)cc1 2 6 Multi-D/A $486.534$ $8.211$
mPTBC N#Cc1ccc(-c2cc(N3c4ccccc4Oc4ccccc43)cc(-c3ccc(C#N)cc3)c2C#N)cc1 2 6 Multi-D/A $486.534$ $8.211$
P3C CCCCC($C$)Cn1c2cc(C#N)ccc2c2ccc(C#Cc3ccc4ccc5cccc6ccc3c4c56)cc21 1 2 A-D-A $528.699$ $10.180$
DBCP c1ccc2c(c1)Oc1ccccc1N2c1cc2c3ccccc3c3ccc4ccc5c6ccccc6c1c1c5c4c3c21 2 0 2D-0A $531.614$ $11.648$
NAQ3 O=C1c2ccccc2C(=O)c2cc(N(c3ccc4c(c3)C(=O)c3ccccc3C4=O)c3ccc4c(c3)C(=O)c3ccccc3C4=O)ccc21 1 0 1D-0A $635.631$ $7.483$
mCP-BP-DMAC O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc2c(c1)c1ccccc1n2-c1cccc(-n2c3ccccc3c3ccccc32)c1 4 0 Multi-D/A $693.806$ $12.687$
BFCZPZ3 c1ccc2c(c1)oc1ccc3c4ccccc4n(-c4cnc(-n5c6ccccc6c6ccc7oc8ccccc8c7c65)cn4)c3c12 2 0 2D-0A $590.642$ $10.470$
BTCZPZ1 c1ccc2c(c1)sc1c2ccc2c3ccccc3n(-c3cnc(-n4c5ccccc5c5ccc6c7ccccc7sc6c54)cn3)c21 2 0 2D-0A $622.778$ $11.407$
BFCZPZ2 c1ccc2c(c1)oc1c2ccc2c1c1ccccc1n2-c1cnc(-n2c3ccccc3c3c4oc5ccccc5c4ccc32)cn1 2 0 2D-0A $590.642$ $10.470$
CZ9CZPZ c1ccc2c(c1)c1ccccc1n2-c1ccc2c(c1)c1ccccc1n2-c1cnc(-n2c3ccccc3c3cc(-n4c5ccccc5c5ccccc54)ccc32)cn1 4 0 Multi-D/A $740.870$ $12.865$
BFCZPZ1 c1ccc2c(c1)oc1c2ccc2c3ccccc3n(-c3cnc(-n4c5ccccc5c5ccc6c7ccccc7oc6c54)cn3)c21 2 0 2D-0A $590.642$ $10.470$
TPBPPI-PY c1ccc(-c2cc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(-c5cccnc5)cc4)cc3)c(-c3ccccc3)c(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $828.031$ $16.396$
TPBPPI-PBI c1ccc(-c2cc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(-c5ccc(-c6nc7ccccc7n6-c6ccccc6)cc5)cc4)cc3)c(-c3ccccc3)c(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown 1 019.2641\,019.264 $20.007$
PTZ-2PTO O=S1(=O)c2ccccc2N(c2ccccc2)c2cc(N3c4ccccc4Sc4ccccc43)ccc21 4 0 Multi-D/A $504.636$ $8.237$
4F-m-$\nu$-DABNA Cc1cccc(N(c2cccc(C)c2)c2cc3c4c(c2)N(c2c(F)cccc2F)c2cc5c(cc2B4c2ccc(C)cc2N3c2cccc(C)c2)B2c3ccc(C)cc3N(c3cccc(C)c3)c3cc(N(c4cccc(C)c4)c4cccc(C)c4)cc(c32)N5c2c(F)cccc2F)c1 6 0 Multi-D/A 1 281.1321\,281.132 $19.815$
m-$\nu$-DABNA Cc1cccc(N(c2cccc(C)c2)c2cc3c4c(c2)N(c2ccccc2)c2cc5c(cc2B4c2ccc(C)cc2N3c2cccc(C)c2)B2c3ccc(C)cc3N(c3cccc(C)c3)c3cc(N(c4cccc(C)c4)c4cccc(C)c4)cc(c32)N5c2ccccc2)c1 6 0 Multi-D/A 1 209.1721\,209.172 $19.259$
TPA-PyI Brc1cccc(-n2c(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)nc(-c3ccccc3)c2-c2ccccc2)c1 1 0 1D-0A $694.676$ $12.773$
4F-$\nu$-DABNA Fc1cccc(F)c1N1c2cc3c(cc2B2c4ccccc4N(c4ccccc4)c4cc(N(c5ccccc5)c5ccccc5)cc1c42)B1c2ccccc2N(c2ccccc2)c2cc(N(c4ccccc4)c4ccccc4)cc(c21)N3c1c(F)cccc1F 6 0 Multi-D/A 1 168.9161\,168.916 $17.348$
CBPTZ c1ccc(-c2nc(-c3ccccc3)nc(-c3cccc(-c4cc(-n5c6ccccc6c6ccccc65)ccc4-c4ccc(-n5c6ccccc6c6ccccc65)cc4)c3)n2)cc1 2 1 D-A-D $791.958$ $14.401$
CBPBM c1ccc(-n2c(-c3ccc(-c4cc(-n5c6ccccc6c6ccccc65)ccc4-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc3ccccc32)cc1 2 0 2D-0A $752.921$ $14.221$
CBPPY c1cncc(-c2cc(-n3c4ccccc4c4ccccc43)ccc2-c2ccc(-n3c4ccccc4c4ccccc43)cc2)c1 2 0 2D-0A $561.688$ $10.610$
N-B Cc1cc(C)c(B(c2ccccc2)c2c(C)cc(C)cc2C)c(C)c1 0 0 Unknown $326.292$ $4.053$
DPACpBr Cc1cc(N2c3ccccc3C(c3ccccc3)(c3ccccc3)c3ccccc32)cc(C)c1Br 1 0 1D-0A $516.482$ $9.232$
DPCN CC1(C)c2ccccc2N(c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(C#N)cc4)cc3)cc2)c2ccccc21 1 2 A-D-A $678.839$ $12.647$
BZC-PXZ O=C(CC(=O)c1ccccc1O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 2 0 2D-0A $421.452$ $6.423$
CzBN N#Cc1c(-n2c3ccccc3c3ccccc32)cccc1-n1c2ccccc2c2ccccc21 2 2 2D-2A $433.514$ $7.752$
2CzBN CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cccc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1C#N 2 2 2D-2A $657.946$ $12.942$
TPA-DPPI c1ccc(-c2nc3c4ccccc4c4ccccc4c3n2-c2ccc(-n3c(-c4ccc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4)nc4c5ccccc5c5ccccc5c43)cc2)cc1 1 0 1D-0A $906.105$ $17.448$
ANQDC-PSTA N#Cc1cc2nc3c(nc2cc1C#N)-c1ccc(N2c4ccc(-c5ccccc5)cc4C4(c5ccccc5-c5ccccc54)c4ccc5c(sc6ccccc65)c42)c2cccc-3c12 1 5 Multi-D/A $815.961$ $14.355$
pDPAPQ-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c3c(cccc13)-c1nc3cc(-c4ccncc4)c(-c4ccncc4)cc3nc1-2 2 1 D-A-D $589.658$ $10.130$
m-CZ-DPS-DMAC CC1(C)c2ccccc2N(c2cccc(S(=O)(=O)c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)c2)c2ccccc21 3 0 Multi-D/A $755.943$ $12.823$
pDPBPZ-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)c1ccccc1c1nc3cc(-c4ccncc4)c(-c4ccncc4)cc3nc21 2 2 2D-2A $615.696$ $10.789$
DDPhCz-DPPZ c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)ccc2n3-c2cc3nc4c5cccnc5c5ncccc5c4nc3cc2-n2c3ccc(-c4ccccc4)cc3c3cc(-c4ccccc4)ccc32)cc1 2 2 2D-2A $917.088$ $16.742$
DCz-DPPZ c1cnc2c(c1)c1nc3cc(-n4c5ccccc5c5ccccc54)c(-n4c5ccccc5c5ccccc54)cc3nc1c1cccnc12 2 2 2D-2A $612.696$ $10.074$
Oxd-bCZ c1ccc2c(c1)c1ccccc1n2-c1ccc(-c2ccc(-c3nnc(-c4ccc(-c5ccc(-n6c7ccccc7c7ccccc76)cc5)cc4)o3)cc2)cc1 2 0 2D-0A $704.833$ $12.932$
Fene c1ccc2c(c1)Oc1ccccc1N2c1ccc2cccc(N3c4ccccc4Oc4ccccc43)c2n1 3 0 Multi-D/A $491.550$ $9.386$
Fens c1ccc2c(c1)Sc1ccccc1N2c1ccc2cccc(N3c4ccccc4Sc4ccccc43)c2n1 3 0 Multi-D/A $523.686$ $10.104$
Yad CC1(C)c2ccccc2N(c2ccc3cccc(N4c5ccccc5C(C)(C)c5ccccc54)c3n2)c2ccccc21 1 0 1D-0A $543.714$ $10.453$
DTPCZPYTZ CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3ccc4c(-c5ccc(-c6nnc(-c7ccccc7)n6-c6ccccc6)cc5)ccc5ccc2c3c54)cc1 1 0 1D-0A $851.110$ $16.525$
TPADSO2 O=S1(=O)c2ccccc2S(=O)(=O)c2cc(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)ccc21 1 0 1D-0A $523.635$ $6.802$
B4PyPPM c1ccc(-c2nc(-c3cc(-c4ccncc4)cc(-c4ccncc4)c3)cc(-c3cc(-c4ccncc4)cc(-c4ccncc4)c3)n2)cc1 0 1 0D-1A $616.728$ $9.726$
mAP CC1(C)c2ccccc2N(c2cccc(N3c4ccccc4C(C)(C)c4ccccc43)c2)c2ccccc21 2 0 2D-0A $492.666$ $9.905$
DBBPZ-DPXZ c1ccc(-c2cc3nc4c5ccc(N6c7ccccc7Oc7ccccc76)cc5c5cc(N6c7ccccc7Oc7ccccc76)ccc5c4nc3cc2-c2ccccc2)cc1 4 2 Multi-D/A $794.914$ $15.574$
DPIBPZ-DPXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(c1)c1cc(N3c4ccccc4Oc4ccccc43)ccc1c1nc3cc(-c4cncnc4)c(-c4cncnc4)cc3nc21 4 4 Multi-D/A $798.866$ $13.154$
m-DTPACO O=C(c1ccc(N(c2ccccc2)c2ccccc2)cc1)c1cccc(C(=O)c2ccc(N(c3ccccc3)c3ccccc3)cc2)c1 2 0 2D-0A $620.752$ $11.088$
mDPBPZ-DPXZ c1cncc(-c2cc3nc4c5ccc(N6c7ccccc7Oc7ccccc76)cc5c5cc(N6c7ccccc7Oc7ccccc76)ccc5c4nc3cc2-c2cccnc2)c1 4 2 Multi-D/A $796.890$ $14.364$
p-DTPACO O=C(c1ccc(C(=O)c2ccc(N(c3ccccc3)c3ccccc3)cc2)cc1)c1ccc(N(c2ccccc2)c2ccccc2)cc1 2 0 2D-0A $620.752$ $11.088$
2TDBA-SBA CC(C)(C)c1ccc2c(c1)B1c3cc(C(C)(C)C)ccc3Oc3cc(N4c5ccccc5C5(c6ccccc64)c4ccccc4N(c4cc6c7c(c4)Oc4ccc(C(C)(C)C)cc4B7c4cc(C(C)(C)C)ccc4O6)c4ccccc45)cc(c31)O2 2 0 2D-0A 1 107.0251\,107.025 $16.281$
TPA-APm c1ccc(N(c2ccccc2)c2ccc(-c3cnc4nc5c(nc4n3)-c3cccc4cccc-5c34)cc2)cc1 1 0 1D-0A $499.577$ $8.357$
TPA-AP c1ccc(N(c2ccccc2)c2ccc(-c3ccc4nc5c(nc4c3)-c3cccc4cccc-5c34)cc2)cc1 1 1 D-A $497.601$ $9.567$
TDBA-SBA CC(C)(C)c1ccc(N2c3ccccc3C3(c4ccccc42)c2ccccc2N(c2cc4c5c(c2)Oc2ccc(C(C)(C)C)cc2B5c2cc(C(C)(C)C)ccc2O4)c2ccccc23)cc1 2 0 2D-0A $858.935$ $14.256$
TPA-APy c1ccc(N(c2ccccc2)c2ccc(-c3cnc4nc5c(nc4c3)-c3cccc4cccc-5c34)cc2)cc1 1 0 1D-0A $498.589$ $8.962$
TPD4PA c1ccc(N(c2ccccc2)c2cc3c4c(c2)N(c2ccccc2)c2cc5c6c(c2B4c2ccccc2N3c2ccccc2)Oc2ccccc2B6c2ccccc2O5)cc1 3 0 Multi-D/A $779.517$ $9.967$
o-Ac2BBP CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccccc3C(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $700.882$ $12.367$
TPA-PyF2 Cc1ccc(N(c2ccc(C3(c4cccnc4)c4cc(C(C)(C)C)ccc4-c4ccc(C(C)(C)C)cc43)cc2)c2ccc(C3(c4cccnc4)c4cc(C(C)(C)C)ccc4-c4ccc(C(C)(C)C)cc43)cc2)cc1 1 0 1D-0A $966.370$ $18.171$
o-Px2BBP O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccccc1C(=O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $648.718$ $11.300$
D-BP-PXZ O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccccc1-c1ccccc1C(=O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $724.816$ $12.967$
BTD-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2nsnc2c1 2 0 2D-0A $317.373$ $5.267$
tBuCzDBA Cc1cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)cc(C)c1B1c2ccccc2B(c2c(C)cc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cc2C)c2ccccc21 2 0 2D-0A $938.962$ $13.650$
5-Cz-Sac CC(C)(C)N1C(=O)c2cc(-n3c4ccccc4c4ccccc43)ccc2S1(=O)=O 1 0 1D-0A $404.491$ $4.727$
6-Cz-Sac CC(C)(C)N1C(=O)c2ccc(-n3c4ccccc4c4ccccc43)cc2S1(=O)=O 1 0 1D-0A $404.491$ $4.727$
DiCz-Sac CC(C)(C)N1C(=O)c2cc(-n3c4ccccc4c4ccccc43)c(-n3c4ccccc4c4ccccc43)cc2S1(=O)=O 2 0 2D-0A $569.686$ $7.824$
2CM CCCCCCCCn1c2ccccc2c2cc(Nc3nc(Nc4ccc5c(c4)c4ccccc4n5CCCCCCCC)nc(OC)n3)ccc21 2 1 D-A-D $695.956$ $12.304$
CzoB Cc1cc(C)c(B(c2ccccc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $491.487$ $7.150$
CzBroB Cc1cc(C)c(B(c2ccc(Br)cc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $570.383$ $7.913$
PxCN N#Cc1c(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)cncc1-c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 1 Multi-D/A $618.696$ $11.438$
AcCN CC1(C)c2ccccc2N(c2ccc(-c3cncc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)c3C#N)cc2)c2ccccc21 2 1 D-A-D $670.860$ $12.505$
CzCNoB Cc1cc(C)c(B(c2ccc(C#N)cc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 2 A-D-A $516.497$ $7.022$
CNPP-TPA N#Cc1nc2c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)nc3c3nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc3c2nc1C#N 2 3 Multi-D/A $768.884$ $12.743$
SPAC-FO O=C1c2ccc(N3c4ccccc4C4(c5ccccc5-c5ccccc54)c4ccccc43)cc2-c2cc(N3c4ccccc4C4(c5ccccc5-c5ccccc54)c4ccccc43)ccc21 2 0 2D-0A $839.010$ $15.191$
DMAC-FO CC1(C)c2ccccc2N(c2ccc3c(c2)-c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2C3=O)c2ccccc21 2 0 2D-0A $594.758$ $11.116$
6AcBIQ CC1(C)c2ccccc2N(c2ccc3c4c(cccc24)C(=O)N(c2ccccc2)C3=O)c2ccccc21 1 0 1D-0A $480.567$ $7.750$
PtB CC(C)(C)c1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $552.721$ $11.068$
BCPBT c1ccc2c(c1)c1ccccc1n2-c1cc(-c2ccc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)c3nsnc23)cc(-n2c3ccccc3c3ccccc32)c1 4 0 Multi-D/A $949.155$ $17.414$
InDM CN(C)c1ccc(-n2c(-c3ccccc3)nc(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $415.540$ $6.939$
PnB CCCCc1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $552.721$ $11.113$
BANE c1ccc(-c2ccc(-c3c4ccccc4c(-c4c5ccccc5c(-c5ccc(-c6ccccc6)cc5)c5ccccc45)c4ccccc34)cc2)cc1 0 0 Unknown $658.844$ $14.634$
CBM CCCCn1c2ccc(C=C(C#N)C#N)cc2c2cc(C=C(C#N)C#N)ccc21 1 4 Multi-D/A $375.435$ $5.456$
DCzBNPh-1 N#Cc1ccc(-c2cccc(-c3ccc(C#N)c(-n4c5ccccc5c5ccccc54)c3)c2)cc1-n1c2ccccc2c2ccccc21 2 4 Multi-D/A $610.720$ $10.958$
MCZ-P2-DTM COc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(OC)cc4)ccc2n3-c2ccc(C(=O)c3ccc4sc5ccccc5c4c3)cn2)cc1 1 0 1D-0A $666.802$ $11.129$
DBTO-PTZ O=S1(=O)c2ccc(N3c4ccccc4Oc4ccccc43)cc2-c2cc(N3c4ccccc4Oc4ccccc43)ccc21 4 0 Multi-D/A $578.649$ $9.651$
$\alpha$-DMAC-DBP CC1(C)c2ccccc2N(c2cccc3nc4c5ccccc5c5ccccc5c4nc23)c2ccccc21 1 2 A-D-A $487.606$ $9.198$
$\alpha$-2DMAC-DBP CC1(C)c2ccccc2N(c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)c3nc4c5ccccc5c5ccccc5c4nc23)c2ccccc21 2 2 2D-2A $694.882$ $13.308$
$\beta$-DI-DBP c1ccc(-n2c3ccccc3c3c2c2c4ccccc4n(-c4ccc5nc6c7ccccc7c7ccccc7c6nc5c4)c2c2c4ccccc4n(-c4ccccc4)c32)cc1 3 2 Multi-D/A $775.915$ $14.381$
DIC-TRZ c1ccc(-c2nc(-c3ccccc3)nc(-n3c4ccccc4c4ccc5c6ccccc6n(-c6ccccc6)c5c43)n2)cc1 2 1 D-A-D $563.664$ $9.400$
PO-TPA CC(C)(C)c1ccc(N(c2ccc(-c3cc4c(cc3C#N)=NC3c5ccc(P(=O)(c6ccccc6)c6ccccc6)nc5-c5nc(P(=O)(c6ccccc6)c6ccccc6)ccc5C3N=4)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 2 A-D-A 1 065.2111\,065.211 $13.075$
DMAC-Cz-TTR CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2n3-c2ccc3c(c2)S(=O)(=O)c2ccccc2S3(=O)=O)c2ccccc21 3 0 Multi-D/A $860.073$ $12.981$
DCzPO Cc1ccc2c(c1)c1cc(C)ccc1n2-c1cccc(-n2c3ccc(C)cc3c3cc(C)ccc32)c1P(=O)(c1ccccc1)c1ccccc1 2 0 2D-0A $664.789$ $10.754$
IQ-oTPA O=C1c2ccccc2-c2nc3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3nc21 2 1 D-A-D $718.860$ $13.115$
2C N#CC(C#N)=Cc1ccc(-c2ccc3ccc4cccc5ccc2c3c45)cc1F 0 2 0D-2A $372.402$ $6.821$
C-H N#CC(C#N)=Cc1c(F)cc(-c2ccc3ccc4cccc5ccc2c3c45)cc1F 0 2 0D-2A $390.392$ $6.960$
PSPP O=S(=O)(c1cccnc1)c1cccc(N2c3ccccc3Oc3ccccc32)c1 2 0 2D-0A $400.459$ $5.490$
2DN CC(C)c1cccc(C(C)C)c1N1C(=O)c2cc(N3c4ccccc4C(C)(C)c4ccccc43)cc3cc(N4c5ccccc5C(C)(C)c5ccccc54)cc(c23)C1=O 2 0 2D-0A $772.005$ $14.105$
PSPBP O=S(=O)(c1cccnc1)c1cc(N2c3ccccc3Oc3ccccc32)cc(N2c3ccccc3Oc3ccccc32)c1 4 0 Multi-D/A $581.653$ $9.065$
2PN CC(C)c1cccc(C(C)C)c1N1C(=O)c2cc(N3c4ccccc4Oc4ccccc43)cc3cc(N4c5ccccc5Oc5ccccc54)cc(c23)C1=O 4 0 Multi-D/A $719.841$ $13.038$
pCbz-PYR c1ccc(-c2ccc3c(c2)c2cc(-c4ccccc4)ccc2n3-c2cc(-n3c4ccc(-c5ccccc5)cc4c4cc(-c5ccccc5)ccc43)ncn2)cc1 2 1 D-A-D $714.872$ $13.339$
PPI-An-NPIM c1ccc(-c2nc(-c3ccccc3)n(-c3ccc(-c4c5ccccc5c(-c5ccc(-c6nc7c8ccccc8c8ccccc8c7n6-c6ccccc6)cc5)c5ccccc45)cc3)c2-c2ccccc2)cc1 0 0 Unknown $917.128$ $17.826$
2PS-2FPh Fc1cc(N2c3ccccc3Sc3ccccc32)c(F)cc1N1c2ccccc2Sc2ccccc21 4 0 Multi-D/A $508.618$ $9.834$
9CzTPE C(=Cc1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc(C(=C(c2ccccc2)c2ccccc2)c2ccccc2)cc1 1 0 1D-0A $599.777$ $11.961$
2CzTPE CCn1c2ccccc2c2ccc(-c3ccc(C=Cc4ccc(C(=C(c5ccccc5)c5ccccc5)c5ccccc5)cc4)cc3)cc21 1 0 1D-0A $627.831$ $12.659$
FTAT-FBO O=P(c1ccccc1)(c1ccccc1)c1ccc([Si](c2ccccc2)(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $536.687$ $4.703$
DmPy-PPI c1ccc(-n2c(-c3ccc(-c4cc(-c5cccc(-c6cccnc6)c5)cc(-c5cccc(-c6cccnc6)c5)c4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $752.921$ $14.124$
mCP-BmPy CC(C)(C)c1ccc2c(c1)C(c1cc(-c3cccnc3)cc(-c3cccnc3)c1)(c1ccc3c(c1)c1ccccc1n3-c1cccc(-n3c4ccccc4c4ccccc43)c1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $915.197$ $16.963$
mCP-Py CC(C)(C)c1ccc2c(c1)C(c1cccnc1)(c1ccc3c(c1)c1ccccc1n3-c1cccc(-n3c4ccccc4c4ccccc43)c1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $762.013$ $14.234$
mCP-Ph CC(C)(C)c1ccc2c(c1)C(c1ccccc1)(c1ccc3c(c1)c1ccccc1n3-c1cccc(-n3c4ccccc4c4ccccc43)c1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $761.025$ $14.839$
DMTAC Cc1ccc(-c2c(-c3ccccc3)c(-c3ccccc3)c3c(N)cc4c5ccccc5c(N)cc4c3c2-c2ccc(C)c(C)c2)cc1C 0 0 Unknown $618.824$ $12.212$
IPTAC Cc1ccc(-c2c(-c3ccc(C(C)C)cc3)c(-c3ccc(C(C)C)cc3)c3c(c(N)cc4c5ccccc5c(N)cc43)c2-c2ccc(C)cc2)cc1 0 0 Unknown $674.932$ $13.842$
TBTAC Cc1ccccc1-c1c(-c2ccc(C(C)(C)C)cc2)c(-c2ccc(C(C)(C)C)cc2)c2c(c(N)cc3c4ccccc4c(N)cc32)c1-c1ccccc1C 0 0 Unknown $702.986$ $14.190$
p2Cz2Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ccccc54)c(-c4nc(-c5ccccc5)nc(-c5ccccc5)n4)cc3-n3c4ccccc4c4ccccc43)n2)cc1 2 2 2D-2A $871.020$ $14.253$
m2Cz2Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-c4nc(-c5ccccc5)nc(-c5ccccc5)n4)c(-n4c5ccccc5c5ccccc54)cc3-n3c4ccccc4c4ccccc43)n2)cc1 2 2 2D-2A $871.020$ $14.253$
2,4-CzPyCl3 Clc1nc(-n2c3ccccc3c3ccccc32)c(Cl)c(-n2c3ccccc3c3ccccc32)c1Cl 2 0 2D-0A $512.827$ $9.236$
DMAC-PCN CC1(C)c2ccccc2N(c2ccc(-c3c(C#N)c(-c4ccccc4)nc(-c4ccccc4)c3C#N)cc2)c2ccccc21 1 2 A-D-A $564.692$ $9.935$
2PXZP Cn1cnc2c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc21 4 1 Multi-D/A $648.726$ $10.848$
2,4,6-CzPyCl2 Clc1c(-n2c3ccccc3c3ccccc32)nc(-n2c3ccccc3c3ccccc32)c(Cl)c1-n1c2ccccc2c2ccccc21 3 0 Multi-D/A $643.577$ $11.680$
1PXZP Cn1cnc2c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)ncnc21 2 1 D-A-D $391.434$ $5.606$
3,6-T2C N#Cc1ccc(-c2nc3c4ccc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4c4cc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)ccc4c3n2-c2ccc(C#N)cc2)cc1 2 4 Multi-D/A $907.093$ $17.016$
3,6-TC N#Cc1ccc(-c2nc3c4ccc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4c4cc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)ccc4c3n2-c2ccccc2)cc1 2 2 2D-2A $882.083$ $17.144$
Ph-CF3 c1ccc2c(c1)c1ccccc1n2-c1ccc(Nc2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 2 0 2D-0A $499.617$ $9.624$
arylene c1ccc2c(c1)c1ccccc1n2-c1ccc(Nc2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 2 0 2D-0A $499.617$ $9.624$
TS-1 O=S(=O)(c1ccccc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 0 1D-0A $383.472$ $5.617$
3,6-MeCzOXD Cc1ccc2c(c1)c1cc(C)ccc1n2-c1ccccc1-c1nnc(-c2ccccc2-n2c3ccc(C)cc3c3cc(C)ccc32)o1 2 0 2D-0A $608.745$ $10.831$
p-PO15NCzDPA O=P(c1ccccc1)(c1ccccc1)c1ccc(-c2c3ccccc3c(-c3ccc(-n4c5cccnc5c5cccnc54)cc3)c3ccccc23)cc1 0 0 Unknown $697.778$ $10.853$
2,7-MeCzOXD Cc1ccc2c3ccc(C)cc3n(-c3ccccc3-c3nnc(-c4ccccc4-n4c5cc(C)ccc5c5ccc(C)cc54)o3)c2c1 2 0 2D-0A $608.745$ $10.831$
BPCN-Cz2Ph N#Cc1ccccc1-c1ccc(-n2c3ccc(-c4ccccc4)cc3c3cc(-c4ccccc4)ccc32)cc1 1 2 A-D-A $496.613$ $9.656$
o-CzOXD c1ccc(-n2c3ccccc3c3ccccc32)c(-c2nnc(-c3ccccc3-n3c4ccccc4c4ccccc43)o2)c1 2 0 2D-0A $552.637$ $9.598$
BPCN-2Cz N#Cc1ccccc1-c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)cc1 2 2 2D-2A $585.710$ $11.086$
Cbz-MI CCCCCCN1C(=O)C(c2ccc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)=C(c2ccc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)C1=O 2 0 2D-0A $816.017$ $14.075$
DBFDPO O=P(c1ccccc1)(c1ccccc1)c1cccc2c1oc1c(P(=O)(c3ccccc3)c3ccccc3)cccc12 0 0 Unknown $568.549$ $6.865$
XCBP Cc1cc(-n2c3ccccc3c3ccccc32)c(C)cc1-c1ccc(C(=O)c2ccccc2)cc1 1 0 1D-0A $451.569$ $8.299$
Cbz-PYR c1ccc2c(c1)c1ccccc1n2-c1cc(-n2c3ccccc3c3ccccc32)ncn1 2 1 D-A-D $410.480$ $6.671$
tCbz-mPYRs CSc1nc(-n2c3ccc(C(C)(C)C)cc3c3c(C(C)(C)C)cccc32)c(C)c(-n2c3ccc(C(C)(C)C)cc3c3c(C(C)(C)C)cccc32)n1 2 1 D-A-D $695.033$ $12.891$
PCBP O=C(c1ccccc1)c1ccc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cc1 1 0 1D-0A $423.515$ $7.682$
MeCZ-TRZ Cc1ccc2c(c1)c1ccccc1n2-c1ccc(-c2nc(-c3ccc(-n4c5ccccc5c5cc(C)ccc54)c(C)c3)nc(-c3ccc(-n4c5ccccc5c5cc(C)ccc54)c(C)c3)n2)cc1C 3 1 Multi-D/A $889.119$ $16.014$
3ca COc1ccc(-c2cc(-c3ccc(C(F)(F)F)cc3)c(-c3nc4ccccc4[nH]3)s2)cc1 0 0 Unknown $450.485$ $7.653$
2MeCZ-TRZ Cc1ccc2c(c1)c1cc(C)ccc1n2-c1ccc(-c2nc(-c3ccc(-n4c5ccc(C)cc5c5cc(C)ccc54)c(C)c3)nc(-c3ccc(-n4c5ccc(C)cc5c5cc(C)ccc54)c(C)c3)n2)cc1C 3 1 Multi-D/A $931.200$ $16.940$
CZ-TRZ Cc1cc(-c2nc(-c3ccc(-n4c5ccccc5c5ccccc54)c(C)c3)nc(-c3ccc(-n4c5ccccc5c5ccccc54)c(C)c3)n2)ccc1-n1c2ccccc2c2ccccc21 3 1 Multi-D/A $847.038$ $15.089$
DMAC-AQ CC1(C)c2ccccc2N(c2ccc3c(c2)C(=O)c2ccccc2C3=O)c2ccccc21 1 0 1D-0A $415.492$ $6.571$
DTPTCzDP-CN CC(C)(C)c1nc(-c2ccc(-n3c4ccc(N(c5ccccc5)c5ccccc5)cc4c4cc(N(c5ccccc5)c5ccccc5)ccc43)c(C#N)c2)nc(C(C)(C)C)n1 3 3 Multi-D/A $794.019$ $14.042$
mP2MPC Cc1cc(-c2cccc(-n3c4ccccc4c4ccccc43)c2)c(C)cc1-c1cccc(-c2nc3c4ccccc4c4ccccc4c3n2-c2ccccc2)c1 1 0 1D-0A $715.900$ $14.047$
XPhos CC(C)C1=CC(C(C)C)=C(c2ccccc2)C(C(C)C)(P(C2CCCCC2)C2CCCCC2)C1 0 0 Unknown $478.745$ $10.624$
TPXZ-as-TAZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)n2)cc1 6 0 Multi-D/A $852.954$ $15.599$
2Trz6CNDBT N#Cc1cccc2c1sc1ccc(-c3nc(-c4ccccc4)nc(-c4ccccc4)n3)cc12 0 3 Multi-D/A $440.531$ $7.112$
TPAPyPI c1ccc(N(c2ccccc2)c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4cccc(-c5ccc6ccc7cccc8ccc5c6c78)c4)cc3)cc2)cc1 1 0 1D-0A $814.004$ $16.700$
TPACzPI c1ccc(N(c2ccccc2)c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4cccc(-c5ccc6c(c5)c5ccccc5n6-c5ccccc5)c4)cc3)cc2)cc1 2 0 2D-0A $855.057$ $16.900$
NA-TNA CC(C)(C)c1ccc(N2C(=O)c3cccc4c(/C=C/c5ccc6cc(N(c7ccccc7)c7ccccc7)ccc6c5)ccc(c34)C2=O)cc1 1 0 1D-0A $648.806$ $11.731$
CzPrSBI O=C1c2ccccc2S(=O)(=O)N1CCCn1c2ccccc2c2ccccc21 1 0 1D-0A $390.464$ $4.029$
m-CzCNDMT Cc1cc(-c2cccc(-n3c4ccccc4c4cc(C#N)ccc43)c2)c(C)cc1-c1cccc(-n2c3ccccc3c3cc(C#N)ccc32)c1 2 4 Multi-D/A $638.774$ $11.575$
DTXO-TPA2 Cc1cc(N(c2ccccc2)c2cccc3ccccc23)ccc1-c1ccc(N(c2ccccc2)c2cccc3ccccc23)cc1C 2 0 2D-0A $616.808$ $13.216$
24PQ-Cz c1ccc2c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)nc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)nc2c1 2 1 D-A-D $612.736$ $11.158$
SFPC O=S(=O)(c1ccc(N2c3ccccc3Sc3ccccc32)cc1)c1ccc(-n2c3ccccc3c3ccccc32)cc1 3 0 Multi-D/A $580.734$ $9.551$
2CzMC COc1cc(=O)oc2cc3c(cc12)c1ccccc1n3-c1ccccc1 1 0 1D-0A $341.366$ $4.899$
PPI-PIM-2 FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-n5c(-c6ccccc6)nc(-c6ccccc6)c5-c5ccccc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $808.907$ $14.871$
4CzMC COc1cc(=O)oc2c1ccc1c2c2ccccc2n1-c1ccccc1 1 0 1D-0A $341.366$ $4.899$
PPI-PIM-1 FC(F)(F)c1ccc(-n2c(-c3ccc(-c4ccc(-c5nc(-c6ccccc6)c(-c6ccccc6)n5-c5ccccc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 0 0 Unknown $808.907$ $14.871$
CzTPA c1ccc(N(c2ccccc2)c2ccccc2-n2c3ccccc3c3ccccc32)cc1 2 0 2D-0A $410.520$ $8.254$
SFX-bPy c1ccc2c(c1)Oc1ccc(-c3ccc4ccccc4n3)cc1C21c2ccccc2-c2ccccc21 0 0 Unknown $459.548$ $8.371$
t–BuCz2BP CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(C(=O)c2ccc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cc2)cc1 2 0 2D-0A $737.044$ $14.302$
SAF-2CzB CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3c(c2)C2(c4cc(-c5ccc(-n6c7ccc(C(C)(C)C)cc7c7cc(C(C)(C)C)ccc76)cc5)ccc4-3)c3ccccc3N(c3ccccc3)c3ccccc32)cc1 3 0 Multi-D/A 1 114.5341\,114.534 $22.551$
F-2CzB CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3c(c2)C(C)(C)c2cc(-c4ccc(-n5c6ccc(C(C)(C)C)cc6c6cc(C(C)(C)C)ccc65)cc4)ccc2-3)cc1 2 0 2D-0A $901.295$ $18.711$
SAF-CzB CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc3c(c2)C2(c4ccccc4-3)c3ccccc3N(c3ccccc3)c3ccccc32)cc1 2 0 2D-0A $761.025$ $15.192$
BPy-TP2 c1ccc(-c2ccc(-c3ccc4c5ccc(-c6ccc(-c7ccccn7)nc6)cc5c5ccccc5c4c3)cn2)nc1 0 0 Unknown $536.638$ $9.394$
Cz–AQ O=C1c2ccccc2C(=O)c2cc(-n3c4ccccc4c4ccccc43)ccc21 1 0 1D-0A $373.411$ $5.559$
-ph-pCzAnBzt Cc1cc(-c2c3ccccc3c(-c3ccc(C#N)cc3)c3ccccc23)c(C)cc1-c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 2 A-D-A $624.787$ $12.580$
PyI c1cc2ccc3cccc4c5[nH]cnc5c(c1)c2c34 0 0 Unknown $242.281$ $4.460$
DVCz-2CzCN C=Cc1ccc(COc2cccc3c2c2ccccc2n3CCCCCCOc2cccc3c2c2ccccc2n3-c2cccc(-n3c4ccccc4c4c(OCCCCCCn5c6ccccc6c6c(OCc7ccc(C=C)cc7)cccc65)cccc43)c2C#N)cc1 4 2 Multi-D/A 1 228.5501\,228.550 $21.691$
ATDBT c1ccc(-c2c3ccccc3c(-c3ccc(-n4c5ccccc5c5c6sc7ccccc7c6ccc54)cc3)c3ccccc23)cc1 1 0 1D-0A $601.774$ $12.792$
DPhBCz c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)ccc32)cc1 2 0 2D-0A $484.602$ $9.548$
AcHPM CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)ncn3)cc2)c2ccccc21 2 1 D-A-D $646.838$ $12.029$
PhCO-DMAC CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccccc3)cc2)c2ccccc21 1 0 1D-0A $389.498$ $7.027$
Diphenylsulfane c1ccc(-c2cccc(-c3nc(-c4cccc(-c5ccccc5)c4)nc(-c4cccc(-c5ccccc5)c4)n3)c2)cc1 0 1 0D-1A $537.666$ $9.874$
2BrCPT Brc1ccc2c(c1)c1cc(Br)ccc1n2-c1ccc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)cc1 1 1 D-A $632.359$ $9.495$
DCbz-Ph-Bz c1ccc(-n2c(-c3ccc(-c4cc(-n5c6ccccc6c6ccccc65)cc(-n5c6ccccc6c6ccccc65)c4)cc3)nc3ccccc32)cc1 2 0 2D-0A $676.823$ $12.554$
DDPA-Ph-Bz c1ccc(N(c2ccccc2)c2cc(-c3ccc(-c4nc5ccccc5n4-c4ccccc4)cc3)cc(N(c3ccccc3)c3ccccc3)c2)cc1 2 0 2D-0A $680.855$ $13.299$
IDAC-MCO Cc1cc2cc(-n3c4ccccc4c4cc5c(cc43)C(C)(C)c3ccccc3N5c3ccccc3)ccc2c(=O)o1 2 0 2D-0A $532.643$ $9.308$
diarylbiindole c1ccc(-c2c(-c3ccccc3)c3ccc4cccc5cc(-c6ccccc6)c(c2-c2ccccc2)c3c45)cc1 0 0 Unknown $506.648$ $11.252$
CNDSB N#CC(=Cc1ccccc1)c1ccc(C(C#N)=Cc2ccccc2)cc1 0 2 0D-2A $332.406$ $5.815$
TMAB Cc1cc(C)c(CC#Cc2ccc(N(C)C)cc2)c(B(c2c(C)c(C)cc(C)c2CC#Cc2ccc(N(C)C)cc2)c2c(C)c(C)cc(C)c2CC#Cc2ccc(N(C)C)cc2)c1C 0 0 Unknown $840.021$ $9.959$
MOF c1cncc(CC(c2nc3ccccc3[nH]2)N(C(Cc2cccnc2)c2nc3ccccc3[nH]2)C(Cc2cccnc2)c2nc3ccccc3[nH]2)c1 0 0 Unknown $680.820$ $8.054$
DICzTrz N#Cc1cc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)ccc1-n1c2ccccc2c2cc3c4ccccc4n4c5ccccc5c(c21)c34 3 3 Multi-D/A $662.756$ $10.991$
PXZ2PTO O=S1(=O)c2ccccc2N(c2ccccc2)c2cc(N3c4ccccc4Oc4ccccc43)ccc21 4 0 Multi-D/A $488.568$ $7.878$
SPBP-DPAC O=C1c2ccccc2C2(c3cc(N4c5ccccc5C(c5ccccc5)(c5ccccc5)c5ccccc54)ccc31)c1ccccc1-c1ccccc12 1 0 1D-0A $675.831$ $11.760$
PX-TRZ Cc1cc(C)c(B(c2ccc(-c3nc(Cl)nc(-c4ccccc4)n3)cc2)c2c(C)cc(C)cc2C)c(C)c1 0 1 0D-1A $515.897$ $6.226$
C2MPI Cc1ccc(N(c2ccc(C)cc2)c2ccccc2C2(c3ccccc3N(c3ccc(C)cc3)c3ccc(C)cc3)CCCCC2)cc1 2 0 2D-0A $626.888$ $13.110$
p-PCzTPD CCCCCCCCN1C(=O)c2c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)sc(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)c2C1=O 2 0 2D-0A $747.964$ $13.233$
m-PCzTPD CCCCCCCCN1C(=O)c2c(-c3cccc(-n4c5ccccc5c5ccccc54)c3)sc(-c3cccc(-n4c5ccccc5c5ccccc54)c3)c2C1=O 2 0 2D-0A $747.964$ $13.233$
o-PCzTPD CCCCCCCCN1C(=O)c2c(-c3ccccc3-n3c4ccccc4c4ccccc43)sc(-c3ccccc3-n3c4ccccc4c4ccccc43)c2C1=O 2 0 2D-0A $747.964$ $13.233$
CPBzFC c1ccc2c(c1)c1ccccc1n2-c1ccc(-c2ccc(-c3ccc4c(c3)Cc3cc(-n5c6ccccc6c6ccccc65)ccc3-4)c3nnsc23)cc1 2 0 2D-0A $706.874$ $12.791$
SiCz2Py2 c1cncc(-c2ccc([Si](c3ccc(-c4cccnc4)cc3)(c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)c1 2 0 2D-0A $821.072$ $11.382$
SiCz1Py3 c1cncc(-c2ccc([Si](c3ccc(-c4cccnc4)cc3)(c3ccc(-c4cccnc4)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)c1 1 0 1D-0A $732.963$ $9.347$
SiCz3Py1 c1cncc(-c2ccc([Si](c3ccc(-n4c5ccccc5c5ccccc54)cc3)(c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)c1 3 0 Multi-D/A $909.181$ $13.417$
POBBPE CC(C)(C)c1ccc(Oc2ccc(C(C)(C)C)cc2P(=O)(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $482.604$ $7.713$
CzppT c1ccc(-c2c(-c3ccccc3)c(-c3ccc(-c4cccc5c4[nH]c4ccccc45)cc3)c(-c3ccccn3)c(-c3ccccn3)c2-c2ccc(-c3cccc4c3[nH]c3ccccc34)cc2)cc1 2 0 2D-0A $867.068$ $17.082$
SBDBQ-PXZ c1ccc(-c2nc3ccc(N4c5ccccc5Oc5ccccc54)cc3nc2-c2ccccc2)cc1 2 1 D-A-D $463.540$ $8.539$
DBQ-3PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nc3ccc(N4c5ccccc5Oc5ccccc54)cc3nc2-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)cc1 6 1 Multi-D/A $825.928$ $15.690$
P4CzCN–BCz C=Cc1ccc(COc2ccc(C(C)(C)c3ccc(O)cc3)cc2)cc1 0 0 Unknown $344.454$ $5.940$
NAP-1,5-DPA COc1ccc(N(c2ccc(OC)cc2)c2cccc3c(N(c4ccc(OC)cc4)c4ccc(OC)cc4)cccc23)cc1 2 0 2D-0A $582.700$ $9.814$
3DPA3CN N#Cc1c(-c2ccc(N(c3ccccc3)c3ccccc3)cc2)c(C#N)c(-c2ccc(N(c3ccccc3)c3ccccc3)cc2)c(C#N)c1-c1ccc(N(c2ccccc2)c2ccccc2)cc1 3 6 Multi-D/A $883.071$ $16.712$
TPA-AN-NA c1ccc(N(c2ccccc2)c2ccc(-c3c4ccccc4c(-c4ccc5ccccc5c4)c4ccccc34)cc2)cc1 1 0 1D-0A $547.701$ $11.950$
OCzBN COc1ccc2c(c1)c1ccccc1n2-c1cc(-n2c3ccccc3c3cc(OC)ccc32)c(C#N)c(-n2c3ccccc3c3cc(OC)ccc32)c1 3 2 Multi-D/A $688.787$ $10.875$
BBCz c1ccc2c(c1)[nH]c1cc3[nH]c4ccccc4c3cc12 2 0 2D-0A $256.308$ $4.956$
IQ-Se COc1c(C#N)c(C)cc2c([Se][Se]c3c(C)nc(N4CCCCC4)c4c(OC)c(C#N)c(C)cc34)c(C)nc(N3CCCCC3)c12 0 4 Multi-D/A $746.676$ $5.032$
dTolmp Cc1ccccc1P(CP(CP(c1ccccc1C)c1ccccc1C)c1ccccc1)c1ccccc1C 0 0 Unknown $562.614$ $8.208$
SiTCNCz N#Cc1cccc(-c2nc(-c3cccc([Si](c4ccccc4)(c4ccccc4)c4ccccc4)c3)nc(-n3c4ccccc4c4ccccc43)n2)c1 1 3 Multi-D/A $681.875$ $7.552$
dpmb Cc1cc(P(c2ccccc2)c2ccccc2)c(P(c2ccccc2)c2ccccc2)cc1C 0 0 Unknown $474.524$ $5.820$
3CzBN N#Cc1c(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)cc1-n1c2ccccc2c2ccccc21 3 2 Multi-D/A $598.709$ $10.850$
TCB O=C(c1ccccc1)c1ccc(-n2c3ccccc3c3cc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)ccc32)cc1 2 0 2D-0A $590.726$ $11.151$
CPzPyC c1ccc2c(c1)c1ccccc1n2-c1cncc(-c2cncc(-n3c4ccccc4c4ccccc43)n2)c1 2 0 2D-0A $487.566$ $7.733$
CPzPzC c1ccc2c(c1)c1ccccc1n2-c1cncc(-c2cncc(-n3c4ccccc4c4ccccc43)n2)n1 2 0 2D-0A $488.554$ $7.128$
CPzPC c1cc(-c2cncc(-n3c4ccccc4c4ccccc43)n2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $486.578$ $8.338$
CBIQD O=C1c2cccc3cccc(c23)C(=O)N1c1nc2ccc(Cl)cc2s1 0 0 Unknown $364.813$ $4.904$
PXZ-QL c1ccc2c(c1)Oc1ccccc1N2c1ccc2ccccc2n1 1 0 1D-0A $310.356$ $5.810$
SBA c1ccc2c(c1)Nc1ccccc1C21c2ccccc2Nc2ccccc21 0 0 Unknown $346.433$ $6.183$
TATP-BP O=C(c1ccc(-n2c3ccccc3c3c4c(c5ccccc5n4-c4ccccc4)c4c(c5ccccc5n4-c4ccccc4)c32)cc1)c1ccc(-n2c3ccccc3c3c4c(c5ccccc5n4-c4ccccc4)c4c(c5ccccc5n4-c4ccccc4)c32)cc1 6 0 Multi-D/A 1 173.3921\,173.392 $21.500$
Pd-1 CCCCN(CCCC)c1ccc(C=Nc2cc(C#N)c(C#N)cc2N=Cc2ccc(N(CCCC)CCCC)cc2O)c(O)c1 0 4 Multi-D/A $620.842$ $9.156$
C3F7 Cc1cc(C)c(B(c2ccccc2-n2c3ccccc3c3ccccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $491.487$ $7.150$
BuCzCF3oB Cc1cc(C)c(B(c2ccc(C(F)(F)F)cc2-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c2c(C)cc(C)cc2C)c(C)c1 1 0 1D-0A $671.700$ $10.764$
phenazasiline Brc1ccc2nc3c4ccc(Br)cc4c4ccccc4c3nc2c1 0 2 0D-2A $438.122$ $6.614$
D–A–D Brc1ccc2nc3c4ccc(Br)cc4c4ccccc4c3nc2c1 0 2 0D-2A $438.122$ $6.614$
pzpy c1ccc(-c2nc3ccccc3n2-c2ccc(-c3c4ccccc4c(-c4ccc(-n5c(-c6ccccc6)nc6ccccc65)cc4)c4ccccc34)cc2)cc1 0 0 Unknown $714.872$ $13.339$
m-PO-ABN N#Cc1ccc(-c2c3ccccc3c(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5o4)c3)c3ccccc23)cc1 0 2 0D-2A $572.667$ $11.313$
PO-ABN N#Cc1ccc(-c2c3ccccc3c(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5o4)c3)c3ccccc23)cc1 0 2 0D-2A $572.667$ $11.313$
p-PO-ABN N#Cc1ccc(-c2c3ccccc3c(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5o4)cc3)c3ccccc23)cc1 0 2 0D-2A $572.667$ $11.313$
teracridan c1ccc2c(c1)Cc1c(cccc1-c1ccc3c(c1-c1cccc4c1Cc1ccccc1N4)Cc1ccccc1N3)N2 0 0 Unknown $539.682$ $9.964$
PTPC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc(-c3cc(-c4ccccc4)c(-c4ccc(-c5nc6c7ccccc7c7ccccc7c6n5-c5ccccc5)cc4)cc3-c3ccccc3)cc2)cc1 1 0 1D-0A 1 028.3561\,028.356 $21.026$
TPTI O=c1[nH]c2ccsc2c2cc3c(=O)[nH]c4ccsc4c3cc12 0 0 Unknown $324.386$ $3.799$
BDBTCz-Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(-n4c5ccc(-c6ccc7sc8ccccc8c7c6)cc5c5cc(-c6ccc7sc8ccccc8c7c6)ccc54)cc3)n2)cc1 1 1 D-A $839.061$ $16.039$
BDBFCz-Trz c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(-n4c5ccc(-c6ccc7oc8ccccc8c7c6)cc5c5cc(-c6ccc7oc8ccccc8c7c6)ccc54)cc3)n2)cc1 1 1 D-A $806.925$ $15.102$
HAP-3MF Cc1cc(C2=NC3=NC(c4ccc(F)c(C)c4)=NC4=NC(c5ccc(F)c(C)c5)=NC(=N2)N34)ccc1F 0 0 Unknown $497.484$ $5.080$
FlCz CC1(C)c2ccccc2-c2ccc(-c3ncc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cn3)cc21 1 1 D-A $513.644$ $9.214$
SFCz c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3ncc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cn3)cc21 1 1 D-A $635.770$ $11.251$
o-TPA-AP-TPA c1ccc(N(c2ccccc2)c2ccccc2-c2c3ccccc3c(-c3ccc4ccc5c(-c6ccccc6N(c6ccccc6)c6ccccc6)ccc6ccc3c4c65)c3ccccc23)cc1 2 0 2D-0A $865.092$ $18.831$
p-TPA-AP-TPA c1ccc(N(c2ccccc2)c2ccc(-c3c4ccccc4c(-c4ccc5ccc6c(-c7ccc(N(c8ccccc8)c8ccccc8)cc7)ccc7ccc4c5c76)c4ccccc34)cc2)cc1 2 0 2D-0A $865.092$ $18.831$
TNPZ N#Cc1ccc(-c2ccc3c(c2)c2cc(-c4ccc(C#N)cc4)ccc2c2nc4cc(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)c(-c5ccc(N(c6ccccc6)c6ccccc6)cc5)cc4nc32)cc1 2 6 Multi-D/A $969.164$ $18.440$
m-TPA-AP-TPA c1ccc(N(c2ccccc2)c2cccc(-c3c4ccccc4c(-c4ccc5ccc6c(-c7cccc(N(c8ccccc8)c8ccccc8)c7)ccc7ccc4c5c76)c4ccccc34)c2)cc1 2 0 2D-0A $865.092$ $18.831$
5,6PXZ-PIDO CC1(C)C(=O)c2cc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)cc2C1=O 4 0 Multi-D/A $688.783$ $12.577$
TRZ-3SO2 O=S(=O)(c1ccccc1)c1cccc(-c2nc(-c3cccc(S(=O)(=O)c4ccccc4)c3)nc(-c3cccc(S(=O)(=O)c4ccccc4)c3)n2)c1 0 1 0D-1A $729.861$ $7.371$
35tCzBPym CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2cncnc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 1 D-A-D $739.020$ $13.092$
35CzBPym O=C(c1cncnc1)c1cc(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c1 2 1 D-A-D $514.588$ $7.902$
25CzBPym O=C(c1cncnc1)c1cc(-n2c3ccccc3c3ccccc32)ccc1-n1c2ccccc2c2ccccc21 2 1 D-A-D $514.588$ $7.902$
25tCzBPym CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c(C(=O)c2cncnc2)c1 2 1 D-A-D $739.020$ $13.092$
BPBiPA c1ccc(-c2nc3ccccc3n2-c2ccc(-c3c4ccccc4c(-c4ccc(-n5c(-c6ccccc6)nc6ccccc65)cc4)c4ccccc34)cc2)cc1 0 0 Unknown $714.872$ $13.339$
BTZ–DMAC CC1(C)c2ccccc2N(c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)c3nsnc23)c2ccccc21 2 0 2D-0A $550.731$ $9.909$
BTZ–DPA c1ccc(N(c2ccccc2)c2ccc(N(c3ccccc3)c3ccccc3)c3nsnc23)cc1 2 0 2D-0A $470.601$ $8.631$
2F c1ccc(N(c2ccccc2)c2ccc3c(c2)c2cc(N(c4ccccc4)c4ccccc4)ccc2n3-c2ccc(-c3ccc4nccnc4c3)cc2)cc1 3 1 Multi-D/A $705.865$ $13.333$
PFBP-2a CC(C)(C)c1ccc2c(c1)C(C)(C)c1cc(C(C)(C)C)ccc1N2c1c($F)c($-c2c(F)c(F)c(F)c(F)c2F)c(F)c1F 1 0 1D-0A $635.614$ $11.310$
PFBP-2b CC(C)(C)c1ccc2c(c1)C(C)(C)c1cc(C(C)(C)C)ccc1N2c1c($F)c($-c2c(F)c(F)c(N3c4ccc(C(C)(C)C)cc4C(C)(C)c4cc(C(C)(C)C)ccc43)c(F)c2F)c(F)c1F 2 0 2D-0A $937.116$ $17.875$
PXZ-PPO Cc1cc2nc(N3c4ccccc4Oc4ccccc43)ccc2c(=O)o1 1 0 1D-0A $342.354$ $5.072$
PXZ-BOO Cc1nc2cc(N3c4ccccc4Oc4ccccc43)ccc2c(=O)o1 2 0 2D-0A $342.354$ $5.072$
PPZTPI CC(C)(C)c1ccc(-n2c(-c3ccc(N4c5ccccc5N(c5ccccc5)c5ccccc54)cc3)nc3c4ccccc4c4ccccc4c32)cc1 2 0 2D-0A $682.871$ $13.549$
PPZPPI c1ccc(N2c3ccccc3N(c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccccc4)cc3)c3ccccc32)cc1 2 0 2D-0A $626.763$ $12.252$
Pra-2DMAC CC1(C)c2ccccc2N(c2ccnc(N3c4ccccc4C(C)(C)c4ccccc43)c2)c2ccccc21 0 0 Unknown $493.654$ $9.300$
BAPCN CC1(C)c2ccccc2N(c2ccccc2)c2ccc3cc(-c4ccc(C#N)cc4)ccc3c21 1 2 A-D-A $436.558$ $8.488$
BACN CC1(C)c2ccccc2N(c2ccccc2)c2ccc3cc(C#N)ccc3c21 1 2 A-D-A $360.460$ $6.821$
BDQC-2 c1ccc(-c2ccc(N(c3ccc(-c4ccccc4)cc3)c3ccc4c5ccccc5n(-c5ccc6nc(-c7ccccc7)c(-c7ccccc7)nc6c5)c4c3)cc2)cc1 2 1 D-A-D $766.948$ $14.865$
NIDPA-1 c1ccc(N(c2ccccc2)c2ccc3c(c2)cc2c4cccc(N(c5ccccc5)c5ccccc5)c4n4c5ccccc5c3c24)cc1 4 0 Multi-D/A $625.775$ $12.929$
NIDPA-2 c1ccc(N(c2ccccc2)c2ccc3c(c2)c2cc(N(c4ccccc4)c4ccccc4)cc4c5c6ccccc6ccc5n3c24)cc1 4 0 Multi-D/A $625.775$ $12.929$
SABIQ c1ccc(N2c3ccccc3C3(c4ccccc4-c4nc5cc6ccccc6cc5nc43)c3ccccc32)cc1 1 1 D-A $509.612$ $8.929$
Perylene c1cc2cccc3c4cccc5cccc(c(c1)c23)c54 0 0 Unknown $252.316$ $5.737$
SAIP c1ccc(N2c3ccccc3C3(c4ccccc4-c4nccnc43)c3ccccc32)cc1 1 0 1D-0A $409.492$ $6.623$
FAC c1ccc2c(c1)Nc1ccccc1C21c2ccccc2-c2ccccc21 0 0 Unknown $331.418$ $6.107$
phenylperylene c1ccc(-c2ccc3cccc4c5cccc6cccc(c2c34)c65)cc1 0 0 Unknown $328.414$ $7.404$
ImPy-2 COc1ccc2cc(-c3nc(-c4ccccc4)c4ccccn34)ccc2c1 0 0 Unknown $350.421$ $5.830$
ImPy-1 c1ccc(-c2nc(-c3ccc4ccccc4c3)n3ccccc23)cc1 0 0 Unknown $320.395$ $5.822$
ImPy-3 c1ccc(-c2nc(-c3ccc4ccc5cccc6ccc3c4c56)n3ccccc23)cc1 0 0 Unknown $394.477$ $7.566$
BTaz-Ar-PXZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nc3ccccc3s2)cc1 2 0 2D-0A $392.483$ $7.539$
FSF4A CC1(C)c2ccccc2-c2ccc(N(c3ccccc3-c3ccccc3)c3cccc4c3-c3ccccc3C43c4ccccc4-c4ccccc43)cc21 1 0 1D-0A $675.875$ $13.473$
TPA-L-BN N#Cc1ccc(CCCc2ccc(N(c3ccccc3)c3ccccc3)cc2)cc1 1 2 A-D-A $388.514$ $7.203$
IpCm-PhBzAc CC(C)c1ccc2occ(-c3ccc(N4c5ccccc5C(C)(C)c5cc6c(cc54)oc4ccccc46)cc3)c(=O)c2c1 1 0 1D-0A $561.681$ $10.592$
DphAn-Br O=C1c2ccccc2C(c2ccccc2)(c2ccc(Br)cc2)c2ccccc21 0 0 Unknown $425.325$ $6.376$
5-DBFBI c1ccc2c(c1)nc1n(-c3ccc(-c4ccc5oc6ccccc6c5c4)cn3)c3ccccc3n21 0 0 Unknown $450.501$ $7.393$
4-DBFBI c1ccc2c(c1)nc1n(-c3cc(-c4ccc5oc6ccccc6c5c4)ccn3)c3ccccc3n21 0 0 Unknown $450.501$ $7.393$
DMACMNPTO Cc1cc(N2c3ccccc3S(=O)(=O)c3ccccc32)c(C)cc1N1c2ccccc2C(C)(C)c2ccccc21 3 0 Multi-D/A $542.704$ $9.029$
1CzSO CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc2c(c1)-c1cc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)ccc1S2(=O)=O 2 0 2D-0A $771.083$ $13.884$
DpAn-BzAc CC1(C)c2ccccc2N(c2ccc(C3(c4ccccc4)c4ccccc4C(=O)c4ccccc43)cc2)c2c1ccc1c2oc2ccccc21 1 0 1D-0A $643.786$ $11.622$
BzITz N#Cc1cc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)ccc1-n1c2ccccc2n2c3ccccc3nc12 0 3 Multi-D/A $539.602$ $7.489$
TPXZ–as-TAZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)n2)cc1 6 0 Multi-D/A $852.954$ $15.599$
TPA-CN-N4-2PY CC(C)(C)c1ccc(N(c2ccc(-c3cc(N)c(N)cc3C#N)cc2)c2ccc(C(C)(C)C)cc2)cc1 1 2 A-D-A $488.679$ $8.454$
TPA-CN-N4 c1ccc2nc3c4cccnc4c4ncccc4c3nc2c1 0 2 0D-2A $282.306$ $3.879$
6,7-DCNQx-DICz N#Cc1cc2nc(-c3ccccc3)c(-c3ccc(-n4c5ccccc5c5c6c(c7ccccc7n6-c6ccccc6)c6c(c7ccccc7n6-c6ccccc6)c54)cc3)nc2cc1C#N 3 5 Multi-D/A $827.951$ $13.998$
PBICT c1ccc(-c2nc(-n3c4ccccc4c4ccc5c6ccccc6n(-c6ccccc6)c5c43)nc(-n3c4ccccc4c4ccc5c6ccccc6n(-c6ccccc6)c5c43)n2)cc1 4 1 Multi-D/A $817.956$ $13.927$
DCBNPE O=[N+]([O-])c1ccc(/C(=C(/c2ccc([N+](=O)[O-])cc2)n2c3ccccc3c3ccccc32)n2c3ccccc3c3ccccc32)cc1 2 0 2D-0A $600.634$ $9.641$
TPA-BT-Q Cc1ccc(N(c2ccc(C)cc2)c2ccc(C3(c4ccc(N(c5ccc(C)cc5)c5ccc(C)cc5)cc4)CCCCC3)cc2)cc1 2 0 2D-0A $626.888$ $13.110$
D5 c1cc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)cc(-c2ccc3c4ccccc4c4ccccc4c3c2)c1 1 0 1D-0A $545.685$ $11.577$
D6 c1cc(-c2ccc3c4ccccc4c4ccccc4c3c2)cc(-n2c3ccccc3c3ccccc32)c1 1 0 1D-0A $469.587$ $9.910$
PxPYM c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2cnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2)cc1 4 1 Multi-D/A $594.674$ $10.962$
D1–D6 c1cc(-c2ccc3c4ccccc4c4ccccc4c3c2)cc(-n2c3ccccc3c3ccccc32)c1 1 0 1D-0A $469.587$ $9.910$
DTPCZTZ CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc(-c3nnc(-c4ccccc4)n3-c3ccccc3)cc2)cc1 1 0 1D-0A $650.870$ $11.960$
DHPZ-2BI c1ccc(-n2c(-c3ccc(N4c5ccccc5N(c5ccc(-c6nc7ccccc7n6-c6ccccc6)cc5)c5ccccc54)cc3)nc3ccccc32)cc1 2 0 2D-0A $718.864$ $12.951$
DHPZ-2BN N#Cc1ccc(N2c3ccccc3N(c3ccc(C#N)cc3)c3ccccc32)cc1 2 4 Multi-D/A $384.442$ $6.683$
SAF-2NP c1ccc2c(c1)-c1ccccc1C21c2ccccc2N(c2ccc3nc4c5cccnc5c5ncccc5c4nc3c2)c2ccccc21 1 2 A-D-A $611.708$ $10.026$
BTPO O=S1(=O)c2ccccc2N2c3ccccc3S(=O)(=O)c3cccc1c32 3 0 Multi-D/A $369.423$ $3.449$
PyDCN–DMAC CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccc(C#N)cc4)nc(-c4ccc(C#N)cc4)c3)cc2)c2ccccc21 1 4 Multi-D/A $564.692$ $9.935$
PyDCN N#Cc1ccc(-c2cc(-c3ccccc3)cc(-c3ccc(C#N)cc3)n2)cc1 0 4 Multi-D/A $357.416$ $5.826$
PXZ-OXD c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)o2)cc1 2 0 2D-0A $403.441$ $6.979$
TCz-3PA-TCz CC(C)(C)c1ccc(-c2ccc3c(c2)c2cc(-c4cnc(-c5ccc6c(c5)c5cc(-c7ccc(C(C)(C)C)cc7)ccc5n6-c5ccc(C(C)(C)C)cc5)cn4)ccc2n3-c2ccc(C(C)(C)C)cc2)cc1 2 0 2D-0A $939.304$ $18.529$
NAI–BiFA Cc1c(-c2ccccc2)cc(N2C(=O)c3cccc4c(N(c5ccc(-c6ccccc6)cc5)c5ccc6c(c5)C(C)(C)c5ccccc5-6)ccc(c34)C2=O)cc1-c1ccccc1 1 0 1D-0A $798.986$ $14.726$
NAI–PhBiFA Cc1c(-c2ccccc2)cc(N2C(=O)c3cccc4c(-c5ccc(N(c6ccc(-c7ccccc7)cc6)c6ccc7c(c6)C(C)(C)c6ccccc6-7)cc5)ccc(c34)C2=O)cc1-c1ccccc1 1 0 1D-0A $875.084$ $16.393$
pDTCz-2DPyS CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(S(=O)(=O)c2ccc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cn2)nc1 2 0 2D-0A $775.075$ $12.694$
2SPBAC-BP O=C(c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3cc4ccccc4cc32)cc1)c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3cc4ccccc4cc32)cc1 2 0 2D-0A $941.146$ $17.517$
SPBAC-PH c1ccc(N2c3ccccc3C3(c4ccccc4-c4ccccc43)c3cc4ccccc4cc32)cc1 1 0 1D-0A $457.576$ $8.986$
DFPBI CC1(C)c2cc(-c3nc(-c4ccccc4)c(-c4ccccc4)n3-c3ccccc3)ccc2-c2ccc(N(c3ccccc3)c3ccccc3)cc21 1 0 1D-0A $655.845$ $12.649$
TDPA–TRZ c1ccc(N(c2ccccc2)c2ccc(-c3nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)n3)cc2)cc1 3 1 Multi-D/A $811.005$ $15.282$
DMAC-11-DPPZ CC1(C)c2ccccc2N(c2ccc3nc4c5ccccc5c5ccccc5c4nc3c2)c2ccccc21 1 2 A-D-A $487.606$ $9.198$
PXZ-11-DPPZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2nc3c4ccccc4c4ccccc4c3nc2c1 2 2 2D-2A $461.524$ $8.665$
PXZ-10-DPPZ c1ccc2c(c1)Oc1ccccc1N2c1cccc2nc3c4ccccc4c4ccccc4c3nc12 2 2 2D-2A $461.524$ $8.665$
CCC c1ccc2c(c1)Cc1c-2ccc2c1-c1ccccc1C2 0 0 Unknown $254.332$ $4.829$
CzmPPC c1ccc(-c2cccc(-c3nc(-c4cccc(-c5ccccc5)c4)nc(-c4cccc(-c5ccccc5)c4)n3)c2)cc1 0 1 0D-1A $537.666$ $9.874$
C2FLA-1 CCCC1(CCC)c2ccccc2-c2ccc(C#Cc3ccc4c(c3)C(CCC)(CCC)c3cc(C#N)ccc3-4)cc21 0 2 0D-2A $547.786$ $10.691$
MDP3FL Cc1cccc(N(c2ccccc2)c2ccc3c(c2)C(C)(C)c2cc(-c4ccc5c(c4)C(C)(C)c4cc(-c6ccc7c(c6)C(C)(C)c6cc(N(c8ccccc8)c8cccc(C)c8)ccc6-7)ccc4-5)ccc2-3)c1 2 0 2D-0A $941.275$ $19.496$
C3FLA-2 CCCCC1(CCCC)c2cc(C#Cc3ccc4c(c3)C(CCC)(CCC)c3cc(C#N)ccc3-4)ccc2-c2ccc(C#Cc3ccc4c(c3)C(CCC)(CCC)c3cc(C#N)ccc3-4)cc21 0 4 Multi-D/A $873.241$ $16.610$
9CzFDPEPO O=P(c1ccccc1)(c1ccccc1)c1ccccc1Oc1ccc(-c2ccc(C3(c4ccc(-n5c6ccccc6c6ccccc65)cc4)c4ccccc4-c4ccccc43)cc2)cc1P(=O)(c1ccccc1)c1ccccc1 1 0 1D-0A 1 052.1631\,052.163 $15.885$
9CzFDPESPO O=P(c1ccccc1)(c1ccccc1)c1ccccc1Oc1ccc(-c2ccc(C3(c4ccc(-n5c6ccccc6c6ccccc65)cc4)c4ccccc4-c4ccccc43)cc2)cc1 1 0 1D-0A $851.986$ $14.245$
CTM O=C(c1ccc(-n2c3ccccc3c3ccccc32)cc1)c1ccc2sc3ccccc3c2c1 1 0 1D-0A $453.566$ $8.383$
tCTM CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(C(=O)c2ccc3sc4ccccc4c3c2)cc1 1 0 1D-0A $565.782$ $10.978$
PSBC CC(C)(C)c1ccc(-n2c(-c3ccc(-c4ccc(S(=O)(=O)c5ccc(-c6ccc(-n7c8ccccc8c8ccccc87)cc6)cc5)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 0 1D-0A $884.117$ $15.560$
26DAcBPy CC1(C)c2ccccc2N(c2ccc(C(=O)c3cccc(C(=O)c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)n3)cc2)c2ccccc21 2 0 2D-0A $701.870$ $11.762$
25DAcBPy CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(C(=O)c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)nc3)cc2)c2ccccc21 2 0 2D-0A $701.870$ $11.762$
26DPXZBPy O=C(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1cccc(C(=O)c2ccc(N3c4ccccc4Oc4ccccc43)cc2)n1 4 0 Multi-D/A $649.706$ $10.695$
BTITrz N#Cc1cc(-c2nc(-c3ccccc3)nc(-c3ccccc3)n2)ccc1-n1c2ccccc2c2c3ccccc3sc21 0 3 Multi-D/A $555.666$ $9.056$
mTPA–PPI c1ccc(N(c2ccccc2)c2ccc(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccccc4)c3)cc2)cc1 1 0 1D-0A $613.764$ $12.136$
PPI c1ccc(-c2nc3c4ccccc4c4ccccc4c3n2-c2ccccc2)cc1 0 0 Unknown $370.455$ $6.999$
CzTPA-m-Trz c1ccc(-c2ccc(N(c3ccc(-c4ccc5c(c4)c4ccccc4n5-c4ccccc4)cc3)c3cccc(-c4nc(-c5ccccc5)nc(-c5ccccc5)n4)c3)cc2)cc1 2 1 D-A-D $793.974$ $14.773$
TPA–2PPITPA c1ccc(N(c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)cc2)c2ccc(-c3ccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccc(N(c5ccccc5)c5ccccc5)cc4)cc3)cc2)cc1 3 0 Multi-D/A 1 316.6251\,316.625 $26.055$
[…]
2Cz2tCzBn CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1c(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1C#N 4 2 Multi-D/A $988.336$ $19.137$
2tCz2CzBn CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c(-n2c3ccccc3c3ccccc32)c(C#N)c1-n1c2ccccc2c2ccccc21 4 2 Multi-D/A $988.336$ $19.137$
PmPmP c1cncc(-c2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)cc(-c3cncnc3)n2)c1 2 2 2D-2A $492.542$ $7.238$
PLQYs CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc2ccc3cc(-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)cc4ccc(c1)c2c34 2 0 2D-0A $757.078$ $15.968$
PXZ-ND c1ccc2c(c1)Oc1ccccc1N2c1ccc2ccc(N3c4ccccc4Oc4ccccc43)nc2n1 2 0 2D-0A $492.538$ $8.781$
CPPyC c1cc(-c2cncc(-n3c4ccccc4c4ccccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $485.590$ $8.943$
DMAC-ND CC1(C)c2ccccc2N(c2ccc3ccc(N4c5ccccc5C(C)(C)c5ccccc54)nc3n2)c2ccccc21 0 0 Unknown $544.702$ $9.848$
PrDPhAc c1ccc(C2(c3ccccc3)c3ccccc3N(c3cncc(N4c5ccccc5C(c5ccccc5)(c5ccccc5)c5ccccc54)n3)c3ccccc32)cc1 0 0 Unknown $742.926$ $12.808$
2Mi2SB Cc1nc2ccccc2n1-c1cc(S(=O)(=O)c2ccc(C(C)(C)C)cc2)c(-n2c(C)nc3ccccc32)cc1S(=O)(=O)c1ccc(C(C)(C)C)cc1 0 0 Unknown $730.956$ $9.242$
BFDMAc-NAI CC(C)(C)c1ccc(N2C(=O)c3cccc4c(N5c6ccccc6C(C)(C)c6ccc7c(oc8ccccc87)c65)ccc(c34)C2=O)cc1 1 0 1D-0A $626.756$ $10.946$
HPIPXZ c1ccc(N(c2ccc(-c3ccc(N(c4ccccc4)c4cccc5ccccc45)cc3)cc2)c2cccc3ccccc23)cc1 2 0 2D-0A $588.754$ $12.600$
3BPy-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2cccnc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $738.032$ $13.697$
2BPy-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2ccccn2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $738.032$ $13.697$
BP-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2ccccc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $737.044$ $14.302$
Cz2ICz c1ccc2c(c1)c1ccccc1n2-c1ccc2c(c1)c1cccc3c4cc(-n5c6ccccc6c6ccccc65)ccc4n2c13 4 0 Multi-D/A $571.683$ $11.031$
Ac-BPCN CC1(C)c2ccccc2N(c2ccc3nc4c5cc(C#N)ccc5c5ccc(C#N)cc5c4nc3c2)c2ccccc21 1 6 Multi-D/A $537.626$ $8.942$
5CzICz c1ccc2c(c1)c1ccccc1n2-c1ccc2c(c1)c1cccc3c4ccccc4n2c31 3 0 Multi-D/A $406.488$ $7.934$
DCPDAPM CC1(C)c2ccccc2N(c2ccc(C(=O)c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)cc2)c2ccccc21 3 0 Multi-D/A $719.888$ $13.221$
SPFS-PXZ O=S1(=O)c2ccc(N3c4ccccc4Oc4ccccc43)cc2C2(c3ccccc3-c3ccccc32)c2cc(N3c4ccccc4Oc4ccccc43)ccc21 4 0 Multi-D/A $742.856$ $12.347$
DPS-PXZ O=S(=O)(c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $580.665$ $9.670$
TX-CzPh O=c1c2ccccc2sc2ccc(-c3ccc4c(c3)c3ccccc3n4-c3ccccc3)cc12 1 0 1D-0A $453.566$ $8.179$
bpipH Brc1cccc(-c2nc3c4cccnc4c4ncccc4c3[nH]2)n1 0 0 Unknown $376.217$ $4.484$
Tri-PXZ-PCN N#Cc1c(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)nc(-c2ccc(N3c4ccccc4Oc4ccccc43)cc2)c(C#N)c1-c1ccc(N2c3ccccc3Oc3ccccc32)cc1 6 2 Multi-D/A $900.998$ $16.552$
ACID–BPPZ CC1(C)c2ccccc2N(c2ccc3c(c2)c2ccccc2c2nc4c5cccnc5c5ncccc5c4nc32)c2cc3c4ccccc4n(-c4ccccc4)c3cc21 2 2 2D-2A $754.897$ $13.392$
C-BPyIA c1ccc(-n2c(-c3ccc(-c4c5ccccc5c(-c5ccc(-c6nc7c8cccc9ccc%10cccc(c%10c98)c7n6-c6ccccc6)cc5)c5ccccc45)cc3)nc3c4cccc5ccc6cccc(c6c54)c32)cc1 0 0 Unknown $963.156$ $19.134$
N-BPyIA c1ccc(-c2nc3c4cccc5ccc6cccc(c6c54)c3n2-c2ccc(-c3c4ccccc4c(-c4ccc(-n5c(-c6ccccc6)nc6c7cccc8ccc9cccc(c9c87)c65)cc4)c4ccccc34)cc2)cc1 0 0 Unknown $963.156$ $19.134$
2-TA-AN c1ccc2c(c1)Sc1ccc(-c3c4ccccc4cc4ccccc34)cc1S2 0 0 Unknown $392.548$ $8.276$
ANOP-MeFAC Cc1ccc2c(c1)C1(c3ccccc3-c3ccccc31)c1cc(C)ccc1N2c1ccc2c3c(cccc13)-c1nc3nonc3nc1-2 1 0 1D-0A $603.685$ $9.577$
ANOP-DMAC CC1(C)c2ccccc2N(c2ccc3c4c(cccc24)-c2nc4nonc4nc2-3)c2ccccc21 1 0 1D-0A $453.505$ $6.923$
4PyCNICz N#Cc1c(-c2ccccc2)nc(-c2ccccc2)nc1-c1ccc(-n2c3ccccc3c3ccc4c(c5ccccc5n4-c4ccccc4)c32)cc1 2 2 2D-2A $663.784$ $11.543$
2PyCNBCz N#Cc1c(-c2ccccc2)nc(-c2ccc(-n3c4ccccc4c4cc(-c5ccc6c(c5)c5ccccc5n6-c5ccccc5)ccc43)cc2)nc1-c1ccccc1 2 2 2D-2A $739.882$ $13.210$
CzA-PPO Cc1ccc(N(c2ccc(C)cc2)c2c3ccccc3c(-c3ccc(P(=O)(c4ccccc4)c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)c3ccccc23)cc1 2 0 2D-0A $814.969$ $14.483$
2PyCNICz N#Cc1c(-c2ccccc2)nc(-c2ccc(-n3c4ccccc4c4ccc5c(c6ccccc6n5-c5ccccc5)c43)cc2)nc1-c1ccccc1 2 2 2D-2A $663.784$ $11.543$
TAA-PPO Cc1ccc(N(c2ccc(C)cc2)c2c3ccccc3c(-c3ccc(P(=O)(c4ccccc4)c4ccccc4)cc3)c3ccccc23)cc1 1 0 1D-0A $649.774$ $11.386$
pSFIAc2 Cc1ccc2c(c1)c1c3c4c(c5c1n2-c1ccc(C(C)(C)C)cc1C51c2ccccc2-c2ccccc21)c1cc(C)ccc1n4-c1ccc(C(C)(C)C)cc1C31c2ccccc2-c2ccccc21 2 0 2D-0A $873.156$ $16.446$
TBN-TPA CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cc(-n5c6ccc(C(C)(C)C)cc6c6cc(C(C)(C)C)ccc65)cc2c43)cc1 3 0 Multi-D/A $922.167$ $16.651$
CzBPDCb c1cc(-c2cccc(-n3c4ccccc4c4ncccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 1 0 1D-0A $485.590$ $8.943$
[2,1-b]IF C1=CC2(C=C3C1=CC1=C3C3(C4=c5ccccc5=CC4=C1)c1ccccc1Sc1ccccc13)c1ccccc1Sc1ccccc12 0 0 Unknown $618.826$ $9.206$
cyclohexanone BrCCCCCCC1(CCCCCCBr)c2cc(Br)ccc2-c2ccc(Br)cc21 0 0 Unknown $650.131$ $10.169$
thio-ether C1=CC2(C=C3C1=CC1=C3C3(C4=c5ccccc5=CC4=C1)c1ccccc1Sc1ccccc13)c1ccccc1Sc1ccccc12 0 0 Unknown $618.826$ $9.206$
ICz-2 c1ccc(-c2cccnc2-c2cc(-c3c[nH]nn3)cc(-c3ccccn3)n2)cc1 0 0 Unknown $376.423$ $4.658$
DDCzTrz c1ccc(-c2nc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)nc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)n2)cc1 4 1 Multi-D/A $970.152$ $17.261$
B-O-dpa CC(C)(C)c1cc2c3c(c1)N(c1ccccc1)c1ccccc1B3c1ccccc1O2 1 0 1D-0A $401.318$ $5.389$
B-dpa-SpiroAC CC(C)(C)c1cc2c3c(c1)N1c4ccccc4C4(c5ccccc5-c5ccccc54)c4cccc(c41)B3c1ccccc1N2c1ccccc1 2 0 2D-0A $638.623$ $9.743$
DCz-OTP c1cc(-c2ccccc2-c2cccc(-n3c4ccccc4c4ccccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $560.700$ $11.215$
B-dpa-Cz CC(C)(C)c1cc2c3c(c1)-n1c4ccccc4c4cccc(c41)B3c1ccccc1N2c1ccccc1 2 0 2D-0A $474.416$ $6.694$
c-D2T CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(N2c3ccccc3C3(c4ccccc4-c4cc5c(cc43)C3(c4ccccc4-5)c4ccccc4N(c4ccc(-c5nc(-c6ccccc6)nc(C6CCCCC6)n5)cc4)c4ccccc43)c3ccccc32)cc1 3 1 Multi-D/A 1 251.6351\,251.635 $23.238$
DPXZ-DPPM O=C(c1ccccc1)c1ccc2nc3c4ccc(N5c6ccccc6Oc6ccccc65)cc4c4cc(N5c6ccccc6Oc6ccccc65)ccc4c3nc2c1 4 2 Multi-D/A $746.826$ $13.471$
CBPMCN N#Cc1ccc(-n2c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 2 A-D-A $560.660$ $9.968$
CPBPMCN N#Cc1ccc(-n2c(-c3ccc(-c4ccc(-n5c6ccccc6c6ccccc65)cc4)cc3)nc3c4ccccc4c4ccccc4c32)cc1 1 2 A-D-A $636.758$ $11.635$
TPB-AC N#Cc1ccc(-c2cc(-c3ccccc3)c(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)cc2-c2ccccc2)cc1 1 2 A-D-A $574.727$ $11.696$
Me-DMAC Cc1nc(C)c(C#N)c(-c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)cc2)c1C#N 1 2 A-D-A $440.550$ $7.218$
iTPBI-CN N#Cc1c(-n2c(-c3ccccc3)nc3ccccc32)cc(-n2c(-c3ccccc3)nc3ccccc32)cc1-n1c(-c2ccccc2)nc2ccccc21 0 2 0D-2A $679.787$ $10.576$
TCP c1ccc2c(c1)c1ccccc1n2-c1cc(-n2c3ccccc3c3ccccc32)cc(-n2c3ccccc3c3ccccc32)c1 3 0 Multi-D/A $573.699$ $10.978$
CDBP-BP-PXZ Cc1cc(-n2c3ccccc3c3ccccc32)ccc1-c1ccc(-n2c3ccccc3c3cc(C(=O)c4ccc(N5c6ccccc6Oc6ccccc65)cc4)ccc32)cc1C 4 0 Multi-D/A $797.958$ $14.971$
CDBP-BP-DMAC Cc1cc(-n2c3ccccc3c3ccccc32)ccc1-c1ccc(-n2c3ccccc3c3cc(C(=O)c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4)ccc32)cc1C 3 0 Multi-D/A $824.040$ $15.505$
3CzPhpPM c1ccc(-c2ccc3c(c2)c2ccccc2n3-c2ccc(-c3cc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)cc1 1 1 D-A $549.677$ $10.242$
SFI23pPM c1ccc(-c2cc(-c3ccc(-n4c5ccccc5c5cc6c(cc54)C4(c5ccccc5-c5ccccc54)c4ccccc4-6)cc3)nc(-c3ccccc3)n2)cc1 1 1 D-A $711.868$ $12.918$
3PCzPFP c1ccc(-n2c3ccccc3c3cc(-c4cnc5oc6ncccc6c5c4)ccc32)cc1 1 0 1D-0A $411.464$ $7.140$
SFI34pPM c1ccc(-c2cc(-c3ccc(-n4c5ccccc5c5c6c(ccc54)-c4ccccc4C64c5ccccc5-c5ccccc54)cc3)nc(-c3ccccc3)n2)cc1 1 1 D-A $711.868$ $12.918$
SBA-2DPS O=S(=O)(c1ccccc1)c1ccc(N2c3ccccc3C3(c4ccccc42)c2ccccc2N(c2ccc(S(=O)(=O)c4ccccc4)cc2)c2ccccc23)cc1 2 0 2D-0A $778.955$ $11.301$
2DMAC-BP-F CC1(C)c2ccccc2N(c2ccc3c(c2)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2c2nc4c(F)cccc4nc32)c2ccccc21 2 2 2D-2A $712.872$ $13.447$
DCA N#Cc1cc(C#N)cc(-c2c3ccccc3c(-c3cc(C#N)cc(C#N)c3)c3ccccc23)c1 0 8 Multi-D/A $430.470$ $6.814$
DPFPhe CCCCCCN1C2=CCC(=CC3c4cc(-c5ccccc5)ccc4-c4ccc(-c5ccccc5)cc43)C=C2Oc2ccccc21 0 0 Unknown $597.802$ $11.710$
DFPFPhe CCCCCCN1C2=CCC(=CC3c4cc(-c5ccc(F)cc5)ccc4-c4ccc(-c5ccc(F)cc5)cc43)C=C2Oc2ccccc21 0 0 Unknown $633.782$ $11.988$
3ai CCCCc1cc2c(s1)-c1sc(CCCC)cc1P2(=O)c1ccc(-n2c3ccccc3c3ccccc32)cc1 1 0 1D-0A $565.744$ $9.212$
DNFPh CCCCCCN1C2=CCC(=CC3c4cc(-c5cccc6ccccc56)ccc4-c4ccc(-c5cccc6ccccc56)cc43)C=C2Sc2ccccc21 0 0 Unknown $713.990$ $14.730$
pCzAnBzt N#Cc1ccc(-c2c3ccccc3c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccccc23)cc1 1 2 A-D-A $520.635$ $10.296$
DPE-DPXZ C(#Cc1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $540.622$ $10.237$
3CzAnBzt N#Cc1ccc(-c2c3ccccc3c(-c3ccc4c(c3)c3ccccc3n4-c3ccccc3)c3ccccc23)cc1 1 2 A-D-A $520.635$ $10.296$
[…]
ADN c1ccc2cc(-c3c4ccccc4c(-c4ccc5ccccc5c4)c4ccccc34)ccc2c1 0 0 Unknown $430.550$ $9.633$
Et3SiH FC1=C(F)C(c2c(F)c(F)c(F)c(F)c2F)(c2c(F)c(F)c(F)c(F)c2F)C(c2c(F)c(F)c(F)c(F)c2F)(c2c(F)c(F)c(F)c(F)c2F)B1c1c(F)c(F)c(F)c(F)c1F 0 0 Unknown $932.132$ $11.081$
DPE-DDMAc CC1(C)c2ccccc2N(c2ccc(C#Cc3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $592.786$ $11.305$
mCz-TAn-CN CC(C)(C)c1ccc2c(-c3cccc(-n4c5ccccc5c5ccccc54)c3)c3cc(C(C)(C)C)ccc3c(-c3ccc(C#N)cc3)c2c1 1 2 A-D-A $632.851$ $12.891$
indenoacridine c1ccc2[nH]c3ccc4c5ccccc5cc4c3cc2c1 1 0 1D-0A $267.331$ $5.628$
MPAc Cc1ccc2c(c1)C(c1ccccc1)(c1ccccc1)c1cc(C)ccc1N2 0 0 Unknown $361.488$ $6.743$
TBAc CC(C)(C)C1=CC=CC(C(C)(C)C)(C2(C3(C(C)(C)C)C=CC=C(C(C)(C)C)C3)c3ccccc3N(c3ccccc3)c3ccccc32)C1 1 0 1D-0A $637.996$ $13.686$
poCz-SO O=P(c1ccccc1)(c1ccccc1)c1ccc2[nH]c3ccc(P(=O)(c4ccccc4)c4ccccc4)cc3c2c1 1 0 1D-0A $567.565$ $6.600$
H4TCPE O=C(O)c1ccc(CC(c2ccc(C(=O)O)cc2)(c2ccc(C(=O)O)cc2)c2ccc(C(=O)O)cc2)cc1 0 0 Unknown $510.498$ $5.057$
CbBPCb c1cc(-c2cccc(-n3c4ccccc4c4cccnc43)c2)cc(-n2c3ccccc3c3cccnc32)c1 0 0 Unknown $486.578$ $8.338$
TMOE COc1ccc(C(=C(c2ccc(OC)cc2)c2ccc(OC)cc2)c2ccc(OC)cc2)cc1 0 0 Unknown $452.550$ $6.728$
FAPbCl3 c1ccc(-c2ccc(-c3ccccc3)c(-c3ccccc3)c2-c2ccccc2)cc1 0 0 Unknown $382.506$ $8.355$
DBT-SADF O=S1(=O)c2ccccc2-c2cc(N3c4ccccc4C4(c5ccccc5-c5ccccc54)c4ccccc43)ccc21 1 0 1D-0A $545.663$ $8.646$
mPXTMP Cc1c(-c2ccc3c(c2)Oc2ccccc2S3)cccc1-c1ccc2c(c1)Oc1ccccc1S2 0 0 Unknown $488.633$ $9.843$
APDC-tPh N#Cc1nc2c(nc1C#N)-c1ccc(-c3cc(-c4ccccc4)cc(-c4ccccc4)c3)c3cccc-2c13 0 2 0D-2A $482.546$ $8.022$
POZDBPHZ c1ccc2c(c1)Oc1ccccc1N2c1ccc2c(ccc3nc4ccc5cc(N6c7ccccc7Oc7ccccc76)ccc5c4nc32)c1 4 2 Multi-D/A $642.718$ $12.240$
CPTBF CC(C)(C)c1ccc2c(c1)C(c1ccc(-n3c4ccccc4c4ccccc43)cc1)(c1ccc(-n3c4ccccc4c4ccccc43)cc1)c1cc(C(C)(C)C)ccc1-2 2 0 2D-0A $761.025$ $14.839$
trimesitylboron Cc1cc(C)c(B(c2c(C)cc(C)cc2C)c2c(C)cc(C)cc2C)c(C)c1 0 0 Unknown $368.373$ $4.979$
CMP c1ccc(C(c2ccccc2)(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $320.435$ $6.069$
CH3OH–CHCl3 O=P(c1ccccc1)(c1ccccc1)c1cccc(-c2ccc3ccc4cccc5ccc2c3c45)c1 0 0 Unknown $478.531$ $7.890$
s-triazine CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc([Si](c2ccccc2)(c2ccccc2)c2ccc(B(O)O)cc2)cc1 1 0 1D-0A $657.739$ $6.436$
[BMIM][PF6] CCCCCCC1(CCCCCC)c2ccccc2-c2ccc(Br)cc21 0 0 Unknown $413.443$ $8.656$
1SBFN c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3cccc4ccccc34)cc21 0 0 Unknown $442.561$ $8.850$
2SBFN c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3ccc4ccccc4c3)cc21 0 0 Unknown $442.561$ $8.850$
F2PA CC1(C)c2ccccc2-c2ccc(N(c3ccccc3)c3ccc4c(c3)C(C)(C)c3ccccc3-4)cc21 1 0 1D-0A $477.651$ $9.769$
TSBFB c1ccc2c(c1)-c1ccccc1C21c2ccccc2-c2ccc(-c3cc(-c4ccc5c(c4)C4(c6ccccc6-c6ccccc64)c4ccccc4-5)cc(-c4ccc5c(c4)C4(c6ccccc6-c6ccccc64)c4ccccc4-5)c3)cc21 0 0 Unknown 1 021.2751\,021.275 $19.718$
DPPP c1ccc(-c2cc(-c3ccccc3)cc(-c3ccccc3-c3ccc4ccc5c(-c6ccccc6-c6cc(-c7ccccc7)cc(-c7ccccc7)c6)ccc6ccc3c4c65)c2)cc1 0 0 Unknown $811.040$ $17.920$
TPA-NZP c1ccc(-c2c3ccccc3c(-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3nsnc23)cc1 1 0 1D-0A $505.646$ $9.648$
mTPA-PPI c1ccc(N(c2ccccc2)c2ccc(-c3cccc(-c4nc5c6ccccc6c6ccccc6c5n4-c4ccccc4)c3)cc2)cc1 1 0 1D-0A $613.764$ $12.136$
BSBFB c1cc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)cc(-c2ccc3c(c2)C2(c4ccccc4-c4ccccc42)c2ccccc2-3)c1 0 0 Unknown $706.888$ $13.708$
BN-DBC-Ph2 C1=CB2C(=C3C(=C4C=C(c5ccccc5)C=CN4c4ccc(-c5ccccc5)cc43)c3ccccc32)C=C1 0 0 Unknown $481.407$ $7.869$
TCbTrz c1ccc(-c2nc(-c3ccccc3)nc(-c3cc(-n4c5ccccc5c5ncccc54)c(-n4c5ccccc5c5ncccc54)c(-n4c5ccccc5c5ncccc54)c3)n2)cc1 0 1 0D-1A $807.921$ $12.349$
Xm-mCP O=c1c2ccccc2oc2ccc(-c3cc(-n4c5ccccc5c5ccccc54)cc(-n4c5ccccc5c5ccccc54)c3)cc12 2 0 2D-0A $602.693$ $10.807$
Xp-mCP CC1(C)OB(c2cc(-n3c4ccccc4c4ccccc43)cc(-n3c4ccccc4c4ccccc43)c2)OC1(C)C 2 0 2D-0A $534.468$ $8.180$
ethylacetate O=S(=O)(c1ccc(N2c3ccccc3C=Cc3ccccc32)cc1)c1ccc(N2c3ccccc3C=Cc3ccccc32)cc1 2 0 2D-0A $600.743$ $10.427$
3DMACIPN CC1(C)c2ccccc2N(c2c(F)c(N3c4ccccc4C(C)(C)c4ccccc43)c(C#N)c(N3c4ccccc4C(C)(C)c4ccccc43)c2C#N)c2ccccc21 3 4 Multi-D/A $767.952$ $13.896$
ICZ CCCCCc1c2[nH]c3ccccc3c2cc2[nH]c3ccccc3c12 2 0 2D-0A $326.443$ $6.688$
BDAVBi C(=Cc1ccc(N(c2ccccc2)c2ccccc2)cc1)c1ccc(-c2ccc(C=Cc3ccc(N(c4ccccc4)c4ccccc4)cc3)cc2)cc1 2 0 2D-0A $692.906$ $14.634$
DPVBi C(=C(c1ccccc1)c1ccccc1)c1ccc(-c2ccc(C=C(c3ccccc3)c3ccccc3)cc2)cc1 0 0 Unknown $510.680$ $10.531$
CBTD2 CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(-c2ccc(-c3ccc(-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)cc3)c3nsnc23)cc1 2 0 2D-0A $843.197$ $16.409$
MeOCzPCN COc1ccc2c(c1)c1cc(OC)ccc1n2-c1ncnc(-n2c3ccc(OC)cc3c3cc(OC)ccc32)c1C#N 2 2 2D-2A $555.594$ $6.577$
Na-DMAC CC1(C)c2ccccc2N(c2ccc(-c3c(C#N)c(-c4cccc5ccccc45)nc(-c4cccc5ccccc45)c3C#N)cc2)c2ccccc21 1 2 A-D-A $664.812$ $12.241$
3Py-DMAC CC1(C)c2ccccc2N(c2ccc(-c3c(C#N)c(-c4cccnc4)nc(-c4cccnc4)c3C#N)cc2)c2ccccc21 1 2 A-D-A $566.668$ $8.725$
6tBPA CC(C)(C)c1ccc(N(c2ccc(C(C)(C)C)cc2)c2c3ccc(C(C)(C)C)cc3c(N(c3ccc(C(C)(C)C)cc3)c3ccc(C(C)(C)C)cc3)c3ccc(C(C)(C)C)cc23)cc1 2 0 2D-0A $849.304$ $18.718$
C-O COc1ccc2c(c1)c1ccccc1n2-c1cc2c(cc1-n1c3ccccc3c3ccccc31)Cc1ccccc1O2 2 0 2D-0A $542.638$ $9.586$
IDFL-2DPA CC1(C)c2ccccc2-c2c1cc1c3cc(N(c4ccccc4)c4ccccc4)ccc3n3c4ccc(N(c5ccccc5)c5ccccc5)cc4c2c13 4 0 Multi-D/A $691.878$ $14.083$
ACRF N#Cc1ccc2c(c1)C1(c3cc(C#N)ccc3-2)c2ccccc2N(c2ccccc2)c2ccccc21 1 4 Multi-D/A $457.536$ $7.576$
TPSi-F c1ccc(C2(c3ccc([Si](c4ccccc4)(c4ccccc4)c4ccccc4)cc3)c3ccccc3-c3ccccc32)cc1 0 0 Unknown $576.815$ $7.427$
PAB c1ccc(N(c2ccccc2)c2cc3c4c(c2)N(c2ccccc2)c2ccccc2B4c2ccccc2N3c2ccccc2)cc1 3 0 Multi-D/A $587.535$ $9.239$
2tPAB CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cc(N(c5ccccc5)c5ccccc5)cc2c43)cc1 3 0 Multi-D/A $811.967$ $14.429$
pDTPA-DTPZ c1ccc(N(c2ccccc2)c2ccc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c4nc5c(nc34)c3sccc3c3ccsc35)cc2)cc1 2 2 2D-2A $779.006$ $15.486$
oDTPA-DTPZ c1ccc(N(c2ccccc2)c2ccc(-c3cc4nc5c(nc4cc3-c3ccc(N(c4ccccc4)c4ccccc4)cc3)c3sccc3c3ccsc35)cc2)cc1 2 2 2D-2A $779.006$ $15.486$
oTPAPO-DTPZ O=P(c1ccccc1)(c1ccccc1)c1cc2nc3c(nc2cc1-c1ccc(N(c2ccccc2)c2ccccc2)cc1)c1sccc1c1ccsc13 1 2 A-D-A $735.874$ $11.989$
BPZPP c1ccc(-c2cc3nc4c5ccc(N6c7ccccc7Oc7ccccc76)nc5c5nc(N6c7ccccc7Oc7ccccc76)ccc5c4nc3cc2-c2ccccc2)cc1 2 2 2D-2A $796.890$ $14.364$
BIPOCz c1ccc(-c2nc(-c3ccc(OCCCCc4cccc5c4[nH]c4ccccc45)cc3)n(-c3ccccc3)c2-c2ccccc2)cc1 1 0 1D-0A $609.773$ $10.909$
tDPAC-BN CC(C)(C)c1ccc(N2c3ccc(C(C)(C)C)cc3B3c4cc(C(C)(C)C)ccc4N(c4ccc(C(C)(C)C)cc4)c4cc(N5c6ccccc6C(c6ccccc6)(c6ccccc6)c6ccccc65)cc2c43)cc1 3 0 Multi-D/A $976.174$ $17.125$
tDMAC-BN CC(C)(C)c1ccc(N(c2ccc(C(C)(C)C)cc2)c2cc3c4c(c2)N2c5ccccc5C(C)(C)c5cccc(c52)B4c2cc(C(C)(C)C)ccc2N3c2ccc(C(C)(C)C)cc2)cc1 3 0 Multi-D/A $852.032$ $15.068$
2ICz-OXD c1ccc(-n2c3ccccc3c3c2ccc2c4ccccc4n(-c4ccc(-c5nnc(-c6ccc(-n7c8ccccc8c8ccc9c(c%10ccccc%10n9-c9ccccc9)c87)cc6)o5)cc4)c23)cc1 4 0 Multi-D/A $883.027$ $15.792$
SIR-B Cc1ccc2c(c1)C1(c3ccc(-c4ccccc4)cc3C(=O)c3ccc(-c4ccccc4)cc31)c1cc(C)ccc1N2c1ccc(C(C)(C)C)cc1 1 0 1D-0A $671.884$ $12.645$
SIR-C Cc1ccc2c(c1)C1(c3ccc(-c4ccccc4-c4ccccc4)cc3C(=O)c3ccc(-c4ccccc4-c4ccccc4)cc31)c1cc(C)ccc1N2c1ccc(C(C)(C)C)cc1 1 0 1D-0A $824.080$ $15.979$
SIR-A Cc1ccc2c(c1)C1(c3ccccc3C(=O)c3ccccc31)c1cc(C)ccc1N2c1ccc(C(C)(C)C)cc1 1 0 1D-0A $519.688$ $9.311$
TRZ3DPA c1ccc(N(c2ccccc2)c2ccc(-c3nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)nc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)n3)cc2)cc1 3 1 Multi-D/A $811.005$ $15.282$
BO3N Nc1ccc2c(c1)Oc1cc(N)cc3c1B2c1ccc(Nc2cccc4cc5ccc6cc7cc8ccccc8cc7cc6c5cc24)cc1O3 0 0 Unknown $641.539$ $9.088$
BPPyA Cc1ccc(C)c(N(c2ccc3ccc4c(N(c5cc(C)ccc5C)c5cccc6c5oc5ccccc56)ccc5ccc2c3c54)c2cccc3c2oc2ccccc23)c1 2 0 2D-0A $772.948$ $16.556$
SiCN2Cz N#Cc1cc(-c2nc(-n3c4ccccc4c4ccccc43)nc(-n3c4ccccc4c4ccccc43)n2)cc([Si](c2ccccc2)(c2ccccc2)c2ccccc2)c1 2 3 Multi-D/A $770.972$ $8.982$
PhPC c1ccc(-c2c3ccccc3c(-c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccccc23)cc1 1 0 1D-0A $495.625$ $10.424$
1MOBI COc1cc(/C=C/c2ccc(-n3cnc4ccccc43)cc2)c(OC)cc1/C=C/c1ccc(-n2cnc3ccccc32)cc1 0 0 Unknown $574.684$ $8.722$
3BrBI Brc1cc(/C=C/c2ccc(-n3cnc4ccccc43)cc2)c(Br)cc1/C=C/c1ccc(-n2cnc3ccccc32)cc1 0 0 Unknown $672.424$ $10.230$
DMBFTX CC1(C)c2ccccc2-c2cccc(-c3ccc4sc5ccc(-c6cccc7c6C(C)(C)c6ccccc6-7)cc5c(=O)c4c3)c21 0 0 Unknown $596.795$ $11.361$
SBF-AnCN N#Cc1ccc(-c2c3ccccc3c(-c3ccc4c(c3)C3(c5ccccc5-c5ccccc53)c3ccccc3-4)c3ccccc23)cc1 0 2 0D-2A $593.729$ $11.542$
DMF-AnCN CC1(C)c2ccccc2-c2ccc(-c3c4ccccc4c(-c4ccc(C#N)cc4)c4ccccc34)cc21 0 2 0D-2A $471.603$ $9.505$
DPF-AnCN N#Cc1ccc(-c2c3ccccc3c(-c3ccc4c(c3)C(c3ccccc3)(c3ccccc3)c3ccccc3-4)c3ccccc23)cc1 0 2 0D-2A $595.745$ $11.562$
DPA c1ccc(-c2c3ccccc3c(-c3ccccc3)c3ccccc23)cc1 0 0 Unknown $330.430$ $7.327$
TPA–DAP–TPA c1ccc(N(c2ccccc2)c2ccc(-c3c4ccccc4c(-c4ccc5ccc6c(-c7c8ccccc8c(-c8ccc(N(c9ccccc9)c9ccccc9)cc8)c8ccccc78)ccc7ccc4c5c76)c4ccccc34)cc2)cc1 2 0 2D-0A 1 041.3101\,041.310 $22.804$
AcDbp CC1(C)c2ccccc2N(c2ccc3c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4c4nc5ccccc5nc4c3c2)c2ccccc21 2 2 2D-2A $694.882$ $13.308$
4-Ac CC(C)(C)c1ccc2c(c1)C(C)(C)c1cc(C(C)(C)C)ccc1N2c1ccc(C(F)(F)F)cc1 1 0 1D-0A $465.603$ $9.409$
czpzpy c1cc(-c2cccc3c2[nH]c2ccccc23)nc(-n2cccn2)c1 1 0 1D-0A $310.360$ $4.569$
NyDPO c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2ccc3nc(-c4ccc(N5c6ccccc6Oc6ccccc65)cc4)ccc3n2)cc1 4 0 Multi-D/A $644.734$ $12.115$
1MPA Cc1nc(N2c3ccccc3C(c3ccccc3)(c3ccccc3)c3ccccc32)cc(N2c3ccccc3C(c3ccccc3)(c3ccccc3)c3ccccc32)n1 0 1 0D-1A $756.953$ $13.117$
LEt CCc1ccccc1P(c1ccccc1CC)c1ccccc1P(c1ccccc1CC)c1ccccc1CC 0 0 Unknown $558.686$ $7.453$
DC-TC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc(C(=O)C(=O)c2ccc(-n3c4ccc(C(C)(C)C)cc4c4cc(C(C)(C)C)ccc43)cc2)cc1 2 0 2D-0A $765.054$ $14.136$
DC-ACR CC1(C)c2ccccc2N(c2ccc(C(=O)C(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $624.784$ $10.970$
1G-F1 c1ccc(P(c2ccccc2)c2cscc2P(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $452.499$ $5.265$
P–Cu–C c1ccc(P(c2ccccc2)c2ccccc2-c2ccccc2P(c2ccccc2)c2ccccc2)cc1 0 0 Unknown $522.568$ $6.870$
BPCP-2CPC N#Cc1cc(-n2c3ccccc3c3cc(-c4ccc(C5(c6ccc(-n7c8ccccc8c8ccccc87)cc6)CCCCC5)cc4)ccc32)c(-n2c3ccccc3c3cc(-c4ccc(C5(c6ccc(-n7c8ccccc8c8ccccc87)cc6)CCCCC5)cc4)ccc32)cc1C#N 4 4 Multi-D/A 1 257.5981\,257.598 $23.653$
2H-Qz c1ccc(-c2cc(-c3ccccn3)cc(-c3nc(-c4ccc(N5c6ccccc6Oc6ccccc65)cc4)c4ccccc4n3)c2)nc1 2 1 D-A-D $617.712$ $10.663$
PPINCN-Cz N#Cc1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc3c4ccc(-c5ccc(-n6c7ccccc7c7ccccc76)cc5)cc4c4cc(-c5ccc(-n6c7ccccc7c7ccccc76)cc5)ccc4c32)c2ccccc12 2 2 2D-2A 1 052.2531\,052.253 $20.449$
PPINCN N#Cc1ccc(-n2c(-c3ccc4ccc5cccc6ccc3c4c56)nc3c4ccccc4c4ccccc4c32)c2ccccc12 0 2 0D-2A $569.667$ $10.921$
BDAPM O=C(c1ccc(N2c3ccccc3[Si](c3ccccc3)(c3ccccc3)c3ccccc32)cc1)c1ccc(N2c3ccccc3[Si](c3ccccc3)(c3ccccc3)c3ccccc32)cc1 2 0 2D-0A $877.208$ $9.239$
12AcCz-PM CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)c2c1ccc1c2c2ccccc2n1-c1ccccc1 2 1 D-A-D $680.855$ $12.684$
23AcCz-PM CC1(C)c2ccccc2N(c2ccc(-c3cc(-c4ccccc4)nc(-c4ccccc4)n3)cc2)c2cc3c(cc21)c1ccccc1n3-c1ccccc1 2 1 D-A-D $680.855$ $12.684$
CDDPI c1ccc2c(-n3c4ccccc4c4ccccc43)ccc(-c3nc4c5ccccc5c5ccccc5c4n3-c3ccc4c(c3)OCCO4)c2c1 1 0 1D-0A $643.746$ $11.020$
DDBP c1ccc(-c2nc(-c3ccc(-c4ccc(N(c5ccccc5)c5ccccc5)cc4)c4ccccc34)n(-c3ccc4c(c3)OCCO4)c2-c2ccccc2)cc1 1 0 1D-0A $723.876$ $12.934$
CCDD O=C(c1ccccc1)c1ccc(-n2c3ccc(N(c4ccccc4)c4ccccc4)cc3c3cc(N(c4ccccc4)c4ccccc4)ccc32)cc1 3 0 Multi-D/A $681.839$ $12.954$
DPNC COc1ccc(N(c2ccc(OC)cc2)c2ccc3c(c2)c2cc(N(c4ccc(OC)cc4)c4ccc(OC)cc4)ccc2n3-c2cccc3ccccc23)cc1 3 0 Multi-D/A $747.895$ $12.911$
LiPr CC(C)c1ccccc1P(c1ccccc1C(C)C)c1ccccc1P(c1ccccc1C(C)C)c1ccccc1C(C)C 0 0 Unknown $614.794$ $9.697$
Spiro-CN Cc1ccc(N(c2ccc(C)cc2)c2ccc3c(c2)C2(c4cc(C#N)ccc4-c4ccc(C#N)cc42)c2cc(N(c4ccc(C)cc4)c4ccc(C)cc4)ccc2-3)cc1 2 4 Multi-D/A $756.953$ $13.947$
TMDBP CC1(C)c2ccccc2Nc2c1ccc1c2C(C)(C)c2ccccc2N1 0 0 Unknown $340.470$ $6.452$
ACRSA O=C1c2ccccc2C2(c3ccccc31)c1ccccc1N(c1ccccc1)c1ccccc12 1 0 1D-0A $435.526$ $7.397$
CAPI c1ccc2c(-n3c(-c4ccc(-c5c6ccccc6c(-n6c7ccccc7c7ccccc76)c6ccccc56)cc4)nc4c5ccccc5c5ccccc5c43)cccc2c1 1 0 1D-0A $761.928$ $15.223$
PhCzSPOTz O=P(c1ccccc1)(c1ccccc1)c1nc(-c2ccccc2)nc(-n2c3ccccc3c3ccccc32)n1 1 1 D-A $522.548$ $6.275$
CCAPI N#Cc1ccc(-n2c(-c3ccc(-c4c5ccccc5c(-n5c6ccccc6c6ccccc65)c5ccccc45)cc3)nc3c4ccccc4c4ccccc4c32)c2ccccc12 1 2 A-D-A $786.938$ $15.094$
2CzCbPy c1cc(-n2c3ccccc3c3ccccc32)nc(-n2c3ccc(-n4c5ccccc5c5ccccc54)cc3c3ncccc32)c1 2 0 2D-0A $575.675$ $9.768$
CzCbPy c1ccc(-n2c3ccc(-n4c5ccccc5c5ccccc54)cc3c3ncccc32)nc1 1 0 1D-0A $410.480$ $6.671$
DCE CC(C)CCC[C@@H](C)[C@H]1CC[C@H]2[C@@H]3CC=C4C[C@@H](O)CC[C@]4(C)[C@H]3CC[C@]12C 0 0 Unknown $386.664$ $7.389$
PICFOCz FC(F)(F)c1cccc(-n2c(-c3ccc(OCCCCn4c5ccccc5c5ccccc54)cc3)nc3c4ccccc4c4ccccc4c32)c1 1 0 1D-0A $675.754$ $11.985$
dpmt Cc1sc(C)c(P(c2ccccc2)c2ccccc2)c1P(c1ccccc1)c1ccccc1 0 0 Unknown $480.553$ $5.881$
BICFOCz FC(F)(F)c1cccc(-n2c(-c3ccc(OCCCCn4c5ccccc5c5ccccc54)cc3)nc(-c3ccccc3)c2-c2ccccc2)c1 1 0 1D-0A $677.770$ $11.859$
19a c1ccc2cc(-c3ccc4cc5cc(-c6ccc7ccccc7c6)ccc5cc4c3)ccc2c1 0 0 Unknown $430.550$ $9.633$
BTPBFCz c1ccc(-c2nc(-c3ccccc3-n3c4ccccc4c4c5oc6ccccc6c5ccc43)c3sc4ccccc4c3n2)cc1 1 1 D-A $593.711$ $11.175$
Ph-OBNA c1ccc(-n2c3ccccc3c3c4c5c(cc32)Oc2ccccc2B5c2ccccc2O4)cc1 1 0 1D-0A $435.291$ $5.511$
HPVAnt CCCCCCc1ccc(C=Cc2ccc3cc4ccccc4cc3c2)cc1 0 0 Unknown $364.532$ $8.286$
BTPBFCz-D3 c1ccc(-c2nc(-c3ccc(-n4c5ccccc5c5ccc6c(c7ccccc7n6-c6ccccc6)c54)cc3-n3c4ccccc4c4c5oc6ccccc6c5ccc43)c3sc4ccccc4c3n2)cc1 3 1 Multi-D/A $924.101$ $17.369$
ICBNTrz4 CC1(C)c2ccccc2-c2cc3c4ccccc4n(-c4ccc(-c5nc(-c6ccccc6)nc(-c6ccccc6)n5)cc4C#N)c3cc21 1 3 Multi-D/A $615.740$ $10.148$
1PCTRz c1ccc(-c2nc(-c3ccccc3)nc(-c3ccc(-n4c5ccccc5c5cc(-n6c7ccccc7c7ccccc76)ccc54)c(-c4ccccc4)c3)n2)cc1 2 1 D-A-D $715.860$ $12.734$
DMACN-B Cc1cc(C)c(N(c2ccc3c(c2)C(c2ccccc2)(c2ccccc2)c2cccc4c2N3c2ccccc2C4(C)C)c2c(C)cc(C)cc2C)c(C)c1 2 0 2D-0A $700.970$ $13.812$
CzSO-1CO CC1(OCOC(=O)c2ccccc2)CC=Cc2c1c1ccccc1n2-c1ccc(S(=O)(=O)c2ccccc2)cc1 0 0 Unknown $549.648$ $6.927$
PPO21 O=P(c1ccccc1)(c1ccccc1)c1ccc(-n2c3ccccc3c3cc(P(=O)(c4ccccc4)c4ccccc4)ccc32)cc1 1 0 1D-0A $643.663$ $8.063$
OXD-7 CC(C)(C)c1ccc(-c2nnc(-c3cccc(-c4nnc(-c5ccc(C(C)(C)C)cc5)o4)c3)o2)cc1 0 0 Unknown $478.596$ $7.716$
TCZ1 CCCCC($C$)Cn1c2ccc(-n3c4ccccc4c4ccccc43)cc2c2cc(-n3c4ccccc4c4ccccc43)ccc21 3 0 Multi-D/A $609.817$ $12.205$
APPT c1cc2c3c(cccc3c1)-c1nc3c4cccnc4c4ncccc4c3nc1-2 0 1 0D-1A $356.388$ $5.527$
ANQDC N#Cc1cc2nc3c(nc2cc1C#N)-c1cccc2cccc-3c12 0 5 Multi-D/A $304.312$ $4.174$
ECDPTT CCn1c2ccccc2c2cc(C(=C(c3ccccc3)c3ccccc3)c3ccc(-c4ccc5c(c4)S(=O)(=O)c4ccccc4S5(=O)=O)cc3)ccc21 1 0 1D-0A $727.907$ $10.468$
1CM CCCCCCCCn1c2ccccc2c2cc(Nc3nc(OC)nc(OC)n3)ccc21 1 1 D-A $433.556$ $6.101$
NPE-AcDPS CC(c1ccccc1)N1c2cc(N3c4ccccc4C(C)(C)c4ccccc43)ccc2S(=O)(=O)c2ccc(N3c4ccccc4C(C)(C)c4ccccc43)cc21 3 0 Multi-D/A $749.980$ $12.950$
DMAC-Cz CC1(C)c2ccccc2N(c2ccc3[nH]c4ccc(N5c6ccccc6C(C)(C)c6ccccc65)cc4c3c2)c2ccccc21 3 0 Multi-D/A $581.763$ $11.539$
VBC-PTZ CCN($C$)c1ccc2c(c1)oc(=O)c1c(=O)oc3cc(N4c5ccccc5Sc5ccccc54)ccc3c12 2 0 2D-0A $532.621$ $7.833$
t-BuPCz CC(C)(C)c1ccc(-c2ccc3[nH]c4ccc(-c5ccc(C(C)(C)C)cc5)cc4c3c2)cc1 1 0 1D-0A $431.623$ $9.250$
4BPy-mDTC CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1cc(C(=O)c2ccncc2)cc(-n2c3ccc(C(C)(C)C)cc3c3cc(C(C)(C)C)ccc32)c1 2 0 2D-0A $738.032$ $13.697$
DC–TC Cc1ccc(N(c2ccc(C)cc2)c2ccc(C3(c4ccc(N(c5ccc(C)cc5)c5ccc(C)cc5)cc4)CCCCC3)cc2)cc1 2 0 2D-0A $626.888$ $13.110$
DFDBQPXZ Fc1cc2nc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)c(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)nc2cc1F 4 1 Multi-D/A $680.714$ $12.393$
DPE C(#Cc1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $540.622$ $10.237$
CBZ CCCCCCOc1ccc2c3ccc(OCCCCCC)cc3n(-c3ccnc(-c4ccc(F)c(C#N)c4F)c3)c2c1 1 2 A-D-A $581.707$ $9.914$
tBuCz2 CC(C)(C)c1ccc2c(c1)c1cc(C(C)(C)C)ccc1n2-c1ccc2[nH]c3ccc(-n4c5ccc(C(C)(C)C)cc5c5cc(C(C)(C)C)ccc54)cc3c2c1 3 0 Multi-D/A $722.033$ $14.705$
26DCzPPy c1cc(-c2cccc(-c3cccc(-n4c5ccccc5c5ccccc54)c3)n2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $561.688$ $10.610$
mCBP c1cc(-c2cccc(-n3c4ccccc4c4ccccc43)c2)cc(-n2c3ccccc3c3ccccc32)c1 2 0 2D-0A $484.602$ $9.548$
benzoxazole Cc1sc2ccccc2c1C1=C(c2c(C)sc3ccccc23)C(F)(F)C(F)(F)C1(F)F 0 0 Unknown $468.487$ $8.563$
dbG2TAZ c1ccc2c(c1)ccc1[nH]c3ccc4ccccc4c3c12 1 0 1D-0A $267.331$ $5.628$
V–TPA C=Cc1ccc(COc2cccc3c2c2ccccc2n3-c2ccc(N(c3ccc(-n4c5ccccc5c5ccccc54)cc3)c3ccc(-n4c5ccccc5c5ccccc54)cc3)cc2)cc1 4 0 Multi-D/A $873.072$ $16.670$
CCDC O=C(c1ccccc1)c1ccc(-n2c3ccc(-n4c5ccccc5c5ccccc54)cc3c3cc(-n4c5ccccc5c5ccccc54)ccc32)cc1 3 0 Multi-D/A $677.807$ $12.209$
PXZPDO O=C(C=C(O)c1ccc(N2c3ccccc3Oc3ccccc32)cc1)c1ccc(N2c3ccccc3Oc3ccccc32)cc1 4 0 Multi-D/A $586.647$ $10.619$
DMAC-BP CC1(C)c2ccccc2N(c2ccc(C(=O)c3ccc(N4c5ccccc5C(C)(C)c5ccccc54)cc3)cc2)c2ccccc21 2 0 2D-0A $596.774$ $11.136$
SBTF c1ccc2c(c1)-c1ccccc1C21c2c(ccc3cc(-c4ccc5ccccc5c4)ccc23)-c2ccc(-c3cc4ccccc4c4ccccc34)c3cccc1c23 0 0 Unknown $718.899$ $15.130$
3-DPM N#CC(C#N)=C(c1cccc(N(c2ccccc2)c2ccccc2)c1)c1cccc(N(c2ccccc2)c2ccccc2)c1 2 2 2D-2A $564.692$ $10.475$
4-DPM N#CC(C#N)=C(c1ccc(N(c2ccccc2)c2ccccc2)cc1)c1ccc(N(c2ccccc2)c2ccccc2)cc1 2 2 2D-2A $564.692$ $10.475$
2CzPy c1ccc(-c2ccc(-n3c4ccccc4c4ccccc43)cc2-n2c3ccccc3c3ccccc32)nc1 2 0 2D-0A $485.590$ $8.943$
2PBCzPy c1ccc(-n2c3ccccc3c3cc(-c4ccc5c(c4)c4ccccc4n5-c4ccc(-c5ccccn5)c(-n5c6ccccc6c6cc(-c7ccc8c(c7)c7ccccc7n8-c7ccccc7)ccc65)c4)ccc32)cc1 4 0 Multi-D/A $968.176$ $18.471$
2BFCzPy c1ccc(-c2ccc(-n3c4ccccc4c4c5oc6ccccc6c5ccc43)cc2-n2c3ccccc3c3c4oc5ccccc5c4ccc32)nc1 2 0 2D-0A $665.752$ $12.742$
CNTPA-CZ N#Cc1ccc(N(c2ccccc2)c2cc(-n3c4ccccc4c4ccccc43)cc(-n3c4ccccc4c4ccccc43)c2)cc1 3 2 Multi-D/A $600.725$ $11.222$
CNTPA-PXZ N#Cc1ccc(N(c2ccccc2)c2cc(N3c4ccccc4Oc4ccccc43)cc(N3c4ccccc4Oc4ccccc43)c2)cc1 5 2 Multi-D/A $632.723$ $12.179$
BMZ-TZ c1cc(-c2nc3ccccc3[nH]2)cc(-c2nnnc(-c3cccc(-c4nc5ccccc5[nH]4)c3)c2-c2cccc(-c3nc4ccccc4[nH]3)c2)c1 0 0 Unknown $657.741$ $9.502$
TPAPm N#Cc1c(-c2ccc(N(c3ccccc3)c3ccccc3)cc2)ncnc1-c1ccc(N(c2ccccc2)c2ccccc2)cc1 2 2 2D-2A $591.718$ $10.622$
CzPm N#Cc1c(-c2ccc(-n3c4ccccc4c4ccccc43)cc2)ncnc1-c1ccc(-n2c3ccccc3c3ccccc32)cc1 2 2 2D-2A $587.686$ $9.876$
TCPZ C1=C(c2cccc(-n3c4ccccc4c4ccccc43)c2)NN(c2cccc(-n3c4ccccc4c4ccccc43)c2)N=C1c1cccc(-n2c3ccccc3c3ccccc32)c1 3 0 Multi-D/A $806.973$ $13.748$
TrisCz-Trz c1ccc(-n2c3ccccc3c3ccccc32)c(-c2nc(-c3ccccc3-n3c4ccccc4c4ccccc43)nc(-c3ccccc3-n3c4ccccc4c4ccccc43)n2)c1 3 1 Multi-D/A $804.957$ $14.164$
PTZ–TAZ c1ccc2c(c1)Oc1ccccc1N2c1ccc(-c2nnc(-c3ccc(N4c5ccccc5Oc5ccccc54)cc3)o2)cc1 4 0 Multi-D/A $584.635$ $10.555$
DDMA-TXO2 CC1(C)c2ccccc2N(c2ccc3c(c2)C(C)(C)c2cc(N4c5ccccc5C(C)(C)c5ccccc54)ccc2S3(=O)=O)c2ccccc21 2 0 2D-0A $672.894$ $11.377$
UGH3 c1ccc([Si](c2ccccc2)(c2ccccc2)c2cccc([Si](c3ccccc3)(c3ccccc3)c3ccccc3)c2)cc1 0 0 Unknown $594.906$ $4.441$
V.2 Data availability
All computational data, analysis scripts, and molecular structures are available at https://github.com/TchapetNjafa.
Acknowledgments
We gratefully acknowledge Dr. Benjamin Panebei Samafou for his generous support and provision of computational resources, which were instrumental in enabling the numerical simulations and data analyses presented in this work.

### v2

@@
-Data-Driven Design Rules for TADF Emitters from a High-Throughput Screening of $747$ Molecules
+Data-Driven Design Guidelines for TADF Emitters from a High-Throughput Screening of $747$ Molecules
 Abstract
 […]
-I Introduction
-Organic Light-Emitting Diodes (OLEDs) based on Thermally Activated Delayed Fluorescence (TADF) have defined the third generation of electroluminescent materials, offering a viable path to achieving near-unity internal quantum efficiency (IQE) without resorting to the expensive and scarce noble metals found in phosphorescent emitters [1, 2, 3].
+keywords:
+TADF, OLED, high-throughput screening, structure-property relationships, singlet-triplet gap, xTB, molecular design
+INTRODUCTION
+Organic Light-Emitting Diodes (OLEDs) based on Thermally Activated Delayed Fluorescence (TADF) have defined the third generation of electroluminescent materials, offering a viable path to achieving near-unity internal quantum efficiency (IQE) without resorting to the expensive and scarce noble metals found in phosphorescent emitters 1, 2, 3.
 The TADF mechanism hinges on the efficient harvesting of non-emissive triplet excitons ($T_{1}$) by up-converting them to the emissive singlet state ($S_{1}$) through Reverse Intersystem Crossing (RISC).
-This process requires satisfying both a thermodynamic condition—a minimized singlet-triplet energy gap ($\Delta E_{\text{ST}}\qty$$\lesssim\qty{0.2}{}$)—and a kinetic condition: a sufficiently high rate of RISC ($k_{\text{RISC}}$) [4, 5].
-The rational design of high-performance TADF emitters is a formidable challenge, dominated by a fundamental trade-off.
-In conventional Donor-Acceptor (D-A) architectures, achieving a small $\Delta E_{\text{ST}}$necessitates a significant spatial separation of the frontier molecular orbitals to minimize the electron exchange energy [6].
-However, this separation inherently reduces the orbital overlap, which not only suppresses the oscillator strength ($f$) and thus the radiative decay rate, but can also hinder the spin-orbit coupling (SOC) necessary for efficient RISC [7, 8].
-Successfully navigating this delicate balance requires exploring a vast and complex chemical space [9].
-Computational chemistry is an indispensable tool for navigating this landscape, yet it presents its own dilemma.
-High-fidelity methods are computationally prohibitive for screening the thousands of potential candidates required for data-driven discovery [10, 11, 12].
-This scalability bottleneck has historically limited computational studies to small, homologous series of molecules, hindering the extraction of universal design principles.
-To bridge this gap, we employ a high-throughput virtual screening (HTVS) protocol based on accelerated semi-empirical methods, which we rigorously validated on a diverse dataset of $747$ experimentally known TADF emitters in our preceding companion paper (hereafter, Article 1) [13].
-That study established our framework as a powerful and validated tool for large-scale molecular ranking and trend analysis.
-In this article, we leverage our validated HTVS protocol to move beyond benchmarking and perform a large-scale, data-driven analysis aimed at elucidating the fundamental structure-property relationships that govern TADF performance.
-By analyzing the results from the entire 747-molecule dataset, we aim to distill the complex interplay of molecular structure, conformation, and electronic properties into a set of clear, quantitative, and actionable design guidelines.
-This study, therefore, seeks to: (i) establish a quantitative performance hierarchy among diverse molecular architectures (D-A, D-A-D, and MR systems); (ii) derive statistically robust design rules linking key structural parameters to both thermodynamic ($\Delta E_{\text{ST}}$) and kinetic (SOC, $k_{\text{RISC}}$) factors; (iii) identify and characterize distinct molecular families through data-driven clustering analysis; and (iv) propose a curated list of high-priority candidate molecules predicted to exhibit exceptional TADF properties.
-By translating a massive computational dataset into actionable design intelligence, this work provides a unified framework to guide and accelerate the discovery of next-generation TADF emitters.
-This paper is structured as follows: Section II details our methodological framework.
+This process requires satisfying both a thermodynamic condition—a minimized singlet-triplet energy gap ($\Delta E_{\text{ST}}\qty$$\lesssim\qty{0.2}{}$)—and a kinetic condition: a sufficiently high rate of RISC ($k_{\text{RISC}}$) 4, 5.
+The rational design of efficient TADF emitters faces an inherent trade-off that complicates optimization.
+In conventional Donor-Acceptor (D-A) architectures, achieving a small $\Delta E_{\text{ST}}$necessitates a significant spatial separation of the frontier molecular orbitals to minimize the electron exchange energy 6.
+However, this separation inherently reduces the orbital overlap, which not only suppresses the oscillator strength ($f$) and thus the radiative decay rate, but can also hinder the spin-orbit coupling (SOC) necessary for efficient RISC 7, 8.
+Optimizing both properties simultaneously demands exploration across diverse molecular scaffolds 9.
+Computational methods can accelerate this search, but they introduce a cost-accuracy trade-off.
+High-fidelity methods are computationally prohibitive for screening the thousands of potential candidates required for data-driven discovery 10, 11, 12.
+This cost barrier has confined most studies to small molecular series.
+Universal design guidelines are therefore still lacking.
+Here, we employ a high-throughput virtual screening (HTVS) protocol based on accelerated semiempirical methods.
+We validated this approach on a data set of $747$ experimentally known TADF emitters in our companion paper (hereafter, Article 1 13), which established its reliability for large-scale molecular ranking and trend analysis.
+In this article, we leverage our validated HTVS protocol to move beyond benchmarking and perform a large-scale analysis of structure-property relationships governing TADF performance.
+By analyzing the entire 747-molecule data set, we aim to establish quantitative design guidelines from the interplay of molecular structure, conformation, and electronic properties.
+We aim to establish an architecture ranking, derive structure-property rules for both $\Delta E_{\text{ST}}$ and $k_{\text{RISC}}$, classify molecules by clustering, and propose synthetic targets with favorable predicted properties.
+From this data set, we extract practical design guidelines.
+These can inform the synthesis of improved TADF materials.
+This paper is structured as follows: Section METHODOLOGICAL FRAMEWORK FOR LARGE-SCALE ANALYSIS details our methodological framework.
 The main results are presented in the subsequent sections, including an analysis of the architectural landscape, a classification of molecular families, and a deep dive into the MR-TADF design paradigm.
-Finally, Section IV summarizes our conclusions and proposes future directions for TADF emitter design.
-II Methodological framework for large-scale analysis
+Finally, Section CONCLUSIONS summarizes our conclusions and proposes future directions for TADF emitter design.
+METHODOLOGICAL FRAMEWORK FOR LARGE-SCALE ANALYSIS
 All structure-property relationships presented in this work were derived from a large-scale data set of $747$ experimentally known TADF emitters, processed using a consistent and validated high-throughput virtual screening (HTVS) protocol.
-The full details, validation, and limitations of this computational framework are established in our preceding work (Article 1) [13].
+The full details, validation, and limitations of this computational framework are established in our preceding work (Article 1 13).
 Here, we provide a concise summary of the essential methodological components and their justifications.
-II.1 Dataset and molecular structures
-The dataset comprises a diverse set of $747$ TADF molecules curated from the literature [14].
-To facilitate systematic analysis, molecules were classified into distinct architectural families (e.g., D-A, D-A-D, Multi-Resonance) using automated SMARTS pattern matching, followed by manual verification.
-Initial 3D structures were generated from SMILES strings, and a rigorous conformational search was performed for each molecule using the CREST program coupled with the GFN2-xTB semi-empirical Hamiltonian [15, 16].
+Data Set and Molecular Structures
+The data set comprises a diverse set of $747$ TADF molecules curated from the literature 14.
+To facilitate systematic analysis, molecules were classified into distinct architectural families (e.g., D-A, D-A-D, Multiresonance) using automated SMARTS pattern matching, followed by manual verification.
+Initial 3D structures were generated from SMILES strings, and a rigorous conformational search was performed for each molecule using the CREST program coupled with the GFN2-xTB semiempirical Hamiltonian 15, 16.
 The lowest-energy conformer was then fully optimized at the GFN2-xTB level to obtain the final ground-state ($S_{0}$) geometry.
-This tight-binding method is specifically parameterized for accurate molecular structures and noncovalent interactions, making it well-suited for the complex and often flexible systems studied here [17].
-II.2 Methodological approximations and justifications
-II.2.1 Hybrid protocol and the vertical approximation compromise
-The HTVS protocol relies on a hybrid computational strategy: ground-state ($S_{0}$) geometries are optimized using GFN2-xTB [16, 18], while excited-state properties ($S_{1}$, $T_{1}$) are calculated using the simplified methods, sTDA-xTB/sTD-DFT-xTB [19, 20].
+This tight-binding method is parametrized for molecular structures and noncovalent interactions, which is appropriate for the flexible systems in our data set 17.
+Methodological Approximations and Justifications
+Hybrid Protocol and the Vertical Approximation Compromise
+The HTVS protocol relies on a hybrid computational strategy: ground-state ($S_{0}$) geometries are optimized using GFN2-xTB 16, 18, while excited-state properties ($S_{1}$, $T_{1}$) are calculated using the simplified methods, sTDA-xTB/sTD-DFT-xTB 19, 20.
 This methodology implicitly employs the vertical approximation.
 […]
-II.2.2 Limitations of the solvation model
-The study of excited-state properties was performed using the ground-state-based ALPB implicit solvation model for toluene.
-We acknowledge that for the highly polarized charge-transfer (CT) states characteristic of TADF emitters, a state-specific solvation model (e.g., SS-PCM) is theoretically required for quantitative accuracy.
-However, such approaches are computationally prohibitive for HTS.
-The ALPB model is considered sufficient for screening as it has been shown that while the absolute solvation effect may be approximated, the relative ordering of molecules and the overall structure-property trends are conserved in the condensed phase.
-II.2.3 Specific limitations for multi-resonance emitters
-It is important to note that Multi-Resonance (MR) TADF emitters are inherently susceptible to multi-reference character [22], a feature that current single-reference based TD-DFT and its semi-empirical variants (sTDA-xTB/sTD-DFT-xTB) may not fully capture.
-Consequently, while our HTS trends provide excellent initial screening recommendations, definitive quantitative validation for MR-TADF candidates should rely on post-hoc applications of high-level methods such as SCS-CC2 or STEOM-DLPNO-CCSD [23].
-II.3 Derivation of structure-property descriptors and theoretical framework
+Solvation Model Selection and Trade-offs
+The treatment of solvation effects in excited-state calculations presents a fundamental trade-off between computational efficiency and physical accuracy, particularly for charge-transfer (CT) states characteristic of TADF emitters.
+This study employed the ground-state-based GBSA implicit solvation model for toluene, which warrants detailed discussion of its limitations and the rationale for this choice in high-throughput screening applications.
+Ground-state Vs. State-specific Solvation Models
+Traditional implicit solvation models like GBSA utilize the ground-state electron density to compute solvation energies for all electronic states.22 While computationally efficient, this approximation becomes problematic for CT states where significant charge redistribution occurs upon excitation.23 The solvent response to electronic excitation can differ substantially from ground-state solvation, as the reorganized electron density in CT states creates different electrostatic interactions with the surrounding medium.24
+State-specific solvation models, such as the state-specific polarizable continuum model (SS-PCM), address this limitation by computing separate solvation energies for each electronic state using the corresponding electron density.24, 25 For TADF emitters, where the $S_{1}$ and $T_{1}$ states often exhibit pronounced CT character with substantial charge separation between donor and acceptor moieties, SS-PCM can provide significantly more accurate solvation energies.6
+Implications for Tadf Properties
+The choice of solvation model has direct implications for predicted TADF properties.
+Solvent polarity modulates the singlet-triplet energy gap ($\Delta E_{\text{ST}}$) through differential solvation of the $S_{1}$ and $T_{1}$ states.26 CT states, with their spatially separated charges, typically exhibit stronger solvation effects than locally excited states.
+Consequently, ground-state-based models may systematically under- or overestimate $\Delta E_{\text{ST}}$  values, particularly in polar solvents.
+Recent computational studies of TADF emitters using state-specific approaches have demonstrated that solvent effects can alter $\Delta E_{\text{ST}}$  by 0.1-0.3 eV compared to gas-phase calculations, with the magnitude depending on the degree of charge separation and solvent polarity.6 Ground-state-based models may capture the overall trend but with reduced quantitative accuracy.
+Computational Cost Considerations
+The primary limitation of state-specific solvation models is computational cost.
+SS-PCM requires separate self-consistent field calculations for each electronic state, typically increasing computational time by a factor of 2-3 compared to ground-state models.25 For high-throughput screening of 747 molecules with multiple conformers, this translates to thousands of additional calculations, making SS-PCM computationally prohibitive within reasonable time and resource constraints.
+Validation of Gbsa for Screening Applications
+Despite its limitations, GBSA has been extensively validated for screening applications in drug discovery and materials science.27 While absolute solvation energies may contain systematic errors, the relative ordering of molecules and structure-property trends are often preserved.
+This makes GBSA suitable for the primary objective of this study: identifying structural patterns and relative rankings rather than predicting absolute TADF efficiencies.
+The key assumption underlying our approach is that the systematic errors introduced by the ground-state approximation affect all molecules similarly, preserving the relative trends essential for deriving design guidelines.
+This assumption is supported by the observation that GBSA models provide reliable relative free energies for structurally similar compounds.27
+Limitations and Future Directions
+We acknowledge several limitations of the GBSA approach for TADF emitters:
+  1. 1.
+    Quantitative accuracy: Absolute $\Delta E_{\text{ST}}$  values may deviate from experimental measurements due to inadequate treatment of CT state solvation.
+  2. 2.
+    Solvent dependence: The model cannot capture how different solvents might alter the relative rankings of molecules.
+  3. 3.
+    Strong CT states: Molecules with highly polarized CT states may be most affected by the ground-state approximation.
+Future studies should validate these findings using state-specific solvation models for representative subsets of molecules, particularly those identified as promising TADF candidates.
+Additionally, experimental validation in different solvents would help assess the transferability of the derived design guidelines.
+Justification for Current Approach
+Given the screening nature of this study and the computational constraints, GBSA represents an appropriate balance between accuracy and efficiency.
+The model enables the systematic analysis of 747 molecules while maintaining sufficient accuracy to identify meaningful structure-property relationships.
+The derived guidelines should be viewed as screening criteria for identifying promising candidates, with the understanding that quantitative predictions require more sophisticated solvation treatments.
+Specific Limitations for Multi-resonance Emitters
+MR-TADF emitters, characterized by their extended $\pi$-conjugation and multiple resonance structures involving boron-nitrogen doping, are known to exhibit multi-reference character in their electronic structure.28, 29 While single-reference methods like sTD-DFT-xTB may not fully capture the quantitative energetics of these systems, previous validation studies have demonstrated that qualitative trends in $\Delta E_{\text{ST}}$  and orbital properties remain consistent between semiempirical, TD-DFT, and post-Hartree-Fock methods.30, 31
+[…]
+Spin-orbit Coupling Calculations
+[…]
+Derivation of Structure-property Descriptors and Theoretical Framework
 To translate the large computational data set into actionable design rules, donor-acceptor (D-A) torsional angles were automatically extracted from all optimized geometries.
-To quantify the degree of charge-transfer and its relation to geometry, we calculated the HOMO-LUMO overlap integral ($S^{\prime}_{\text{HL}}$) and the distance between the centroids of the hole and electron densities ($\Delta r_{\text{CT}}$) using the Multiwfn package [24, 25].
+To quantify the degree of charge-transfer and its relation to geometry, we calculated the HOMO-LUMO overlap integral ($S^{\prime}_{\text{HL}}$) and the distance between the centroids of the hole and electron densities ($\Delta r_{\text{CT}}$) using the Multiwfn package 38, 39.
 Principal Component Analysis (PCA) was performed on the standardized matrix of all descriptors to identify the underlying variables governing TADF performance.
@@
 According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs.
-local-excitation) [26, 27].
+local-excitation) 40, 41.
 Our chosen descriptors—particularly the D-A torsional angle—are direct proxies for tuning this CT/LE character balance.
 While our protocol does not compute SOC explicitly due to its prohibitive cost, analyzing these structural and electronic descriptors provides a robust, qualitative assessment of the kinetic factors that, alongside $\Delta E_{\text{ST}}$, govern overall TADF performance.
-III Results and discussion: from statistical trends to rational design principles
-This section leverages the large-scale dataset of $747$ TADF emitters, processed with the validated HTVS protocol detailed in our companion paper (Article 1)[13], to extract quantitative, statistically robust structure-property relationships.
+RESULTS AND DISCUSSION: FROM STATISTICAL TRENDS TO RATIONAL DESIGN PRINCIPLES
+This section leverages the large-scale data set of $747$ TADF emitters, processed with the validated HTVS protocol detailed in our companion paper (Article 1 13), to extract quantitative, statistically robust structure-property relationships.
 Our analysis transitions from broad statistical trends to specific, actionable molecular design rules, with a focus on providing a deep mechanistic understanding that connects molecular structure to the key photophysical parameters governing TADF efficiency.
-III.1 Low-dimensional nature of the TADF design space
+Low-dimensional Nature of the Tadf Design Space
 […]
 […]
-III.2 Architectural hierarchy defines TADF potential
+Architectural Hierarchy Defines Tadf Potential
 […]
@@
 D-A-D architectures demonstrate statistically superior characteristics (lower median and narrower distribution of $\Delta E_{\text{ST}}$) compared to other systems, confirming the efficacy of this design principle for consistently achieving small energy gaps.
-III.3 Data-driven clustering reveals distinct molecular families
+Data-driven Clustering Reveals Distinct Molecular Families
 To complement the structure-based classification, we performed an unsupervised K-means clustering based on the molecules’ photophysical properties.
-This data-driven approach identifies four distinct molecular families, as shown in Figure 3:
+This data-driven approach identifies four distinct molecular families, as shown in Figure˜3:
 […]
@@
 The analysis partitions the data set into distinct families, most notably identifying a cluster of high-efficiency candidates (Cluster 1) with desirable properties for TADF applications.
-III.4 Multi-resonance TADF: a distinct paradigm for blue emitters
+Multi-resonance Tadf: a Distinct Paradigm for Blue Emitters
 […]
 However, as noted in the methodology, it is important to interpret these results with caution.
-The inherent electronic complexity of MR-TADF systems makes them prone to multi-reference character [22], which is not fully captured by the single-reference nature of the semi-empirical methods used in this HTS.
-The trends presented should therefore be interpreted as a valuable guide for compound ranking rather than a source of definitive quantitative predictions, with final validation requiring higher levels of theory.
+The inherent electronic complexity of MR-TADF systems makes them prone to multi-reference character 42, which is not fully captured by the single-reference nature of the semiempirical methods used in this HTS.
+These results guide compound ranking but require higher-level validation for quantitative accuracy.
 Table 1: Photophysical property comparison of MR-TADF with conventional D-A systems, calculated with sTDA-xTB in toluene.
@@
 […]
-III.5 Deconstructing performance: geometric and electronic design levers
-III.5.1 Torsional angle as a primary control parameter
+Deconstructing Performance: Geometric and Electronic Design Levers
+Torsional Angle as a Primary Control Parameter
 […]
 […]
-III.5.2 Quantifying the torsion-induced electronic trade-off
+Quantifying the Torsion-induced Electronic Trade-off
 This geometric control is directly reflected in the electronic structure.
-As shown in Figure 5, twisting the molecule reduces the HOMO-LUMO overlap ($S^{\prime}_{\text{HL}}$), which correlates strongly with a smaller $\Delta E_{\text{ST}}$(Spearman $\rho=-0.45$).
+As shown in Figure˜5, twisting the molecule reduces the HOMO-LUMO overlap ($S^{\prime}_{\text{HL}}$), which correlates strongly with a smaller $\Delta E_{\text{ST}}$  (Spearman $\rho=-0.45$).
 However, this comes at a cost: the oscillator strength, essential for a high radiative rate, is directly proportional to this overlap ($\rho=0.78$).
 This visualization makes the fundamental TADF design trade-off explicit: minimizing the energy gap via orbital separation inherently suppresses emission intensity.
-An optimal region for high-performance emitters lies where there is intermediate overlap (0.15 to 0.300.150.30) that balances these competing requirements.
-To move beyond this qualitative picture and establish a more rigorous quantitative link, future work could employ methods such as Natural Transition Orbitals (NTOs) or the RespA procedure [29, 30].
-These tools would allow for a compact representation of the virtual excitations and rigorously connect the molecular geometry to the precise charge-transfer character (e.g., via centroid distance $\Delta r$) and the ensuing coupling efficiency [31].
+An optimal region for high-performance emitters lies where there is intermediate overlap ($0.150.30$) that balances these competing requirements.
+To move beyond this qualitative picture and establish a more rigorous quantitative link, future work could employ methods such as Natural Transition Orbitals (NTOs) or the RespA procedure 43, 44.
+These tools would allow for a compact representation of the virtual excitations and rigorously connect the molecular geometry to the precise charge-transfer character (e.g., via centroid distance $\Delta r$) and the ensuing coupling efficiency 45.
 Figure 5: Correlation analysis between HOMO-LUMO overlap and key photophysical properties.
 The strong negative correlation with $\Delta E_{\text{ST}}$  (left) validates the spatial separation model for TADF design, while the strong positive correlation with oscillator strength (right) highlights the fundamental design trade-off that molecular engineers must navigate.
-III.6 Implications for reverse intersystem crossing ($k_{\text{RISC}}$)
+Implications for Reverse Intersystem Crossing ($k_{\text{risc}}$)
 While a small $\Delta E_{\text{ST}}$  is the primary thermodynamic requirement for TADF, the overall efficiency is governed by the kinetics of RISC.
 While direct calculation of the $k_{\text{RISC}}$ rate is outside the scope of HTS, the derived structural principles must be explicitly connected to the kinetics of TADF.
-Within the Marcus theory framework, $k_{\text{RISC}}$ is proportional to the square of the spin-orbit coupling (SOC) matrix element and exponentially dependent on $\Delta E_{\text{ST}}$and the reorganization energy ($\lambda$) [32].
+Within the Marcus theory framework, $k_{\text{RISC}}$ is proportional to the square of the spin-orbit coupling (SOC) matrix element and exponentially dependent on $\Delta E_{\text{ST}}$  and the reorganization energy ($\lambda$) 46.
 According to El-Sayed’s rules, SOC is maximized when the interconverting singlet ($S_{1}$) and triplet ($T_{1}$) states have different orbital characters (e.g., charge-transfer vs.
-local-excitation, 1CT $\leftrightarrow$ 3LE) [26].
+local-excitation, 1CT $\leftrightarrow$ 3LE) 40.
 This creates a mechanistic challenge: the very act of creating a pure CT state in $S_{1}$ to minimize $\Delta E_{\text{ST}}$  can lead to a $T_{1}$ state that also has pure CT character, making the transition spin-forbidden (1CT $\leftrightarrow$ 3CT) and suppressing $k_{\text{RISC}}$.
@@
 Molecular descriptor Influence on RISC factor Optimal range/strategy
-HOMO-LUMO overlap Directly modulates $\Delta E_{\text{ST}}$ (thermodynamics) 0.15 to 0.300.150.30
+HOMO-LUMO overlap Directly modulates $\Delta E_{\text{ST}}$ (thermodynamics) $0.150.30$
 […]
@@
 Conformational rigidity (e.g., in MR systems) Minimizes reorganization energy $\mathbf{\lambda}$ (kinetics) Use bridged or fused structures
-III.7 Synthesis of a data-driven design strategy
-The combination of these large-scale statistical analyses allows us to formulate a set of quantitative, data-driven design guidelines for accelerating the discovery of novel TADF emitters, summarized in Table 3.
+Synthesis of a Data-driven Design Strategy
+The combination of these large-scale statistical analyses allows us to formulate a set of quantitative, data-driven design guidelines for accelerating the discovery of novel TADF emitters, summarized in Table˜3.
 These rules provide a clear roadmap for molecular engineering, from selecting a high-performance architecture to fine-tuning its properties via geometric and electronic modifications.
@@
 […]
-HOMO-LUMO Overlap ($S^{\prime}_{\text{HL}}$) 0.15 to 0.300.150.30 Navigates the trade-off between $\Delta E_{\text{ST}}$  and oscillator strength.
+HOMO-LUMO Overlap ($S^{\prime}_{\text{HL}}$) $0.150.30$ Navigates the trade-off between $\Delta E_{\text{ST}}$  and oscillator strength.
 Oscillator Strength ($f_{S_{1}}$) >0.1>0.1 Ensures a high radiative decay rate for bright emission.
-III.8 Future outlook and data exploitation
-III.8.1 The HTS dataset as a machine learning substrate
+Future Outlook and Data Exploitation
+The Hts Data Set as a Machine Learning Substrate
 […]
-III.8.2 Actionable design targets
+Actionable Design Targets
 […]
-IV Conclusions
-In this work, we performed a large-scale computational screening of $747$ TADF emitters to extract quantitative, statistically robust structure-property relationships for guiding molecular design.
-By systematically analyzing the impact of molecular architecture, conformational geometry, and electronic structure, we have moved beyond anecdotal evidence to establish a data-driven framework that integrates both the thermodynamic and kinetic principles of TADF engineering.
+Sensitivity Analysis of Design Guidelines
+To assess the robustness of our HTVS-derived design guidelines, we performed a comprehensive sensitivity analysis examining how variations in threshold choices affect the identification of promising TADF candidates.
+This analysis addresses the inherent subjectivity in defining quantitative cutoffs and provides confidence intervals for our screening criteria.
+Threshold Sensitivity for $\Delta E_{\text{ST}}$  Criteria
+We systematically varied the $\Delta E_{\text{ST}}$  threshold from 0.2 to 0.4 eV to examine its impact on candidate selection.
+The results demonstrate that threshold choice significantly influences the number of molecules classified as promising: $\Delta E_{\text{ST}}$  $\leq 0.2$ eV identifies 163 molecules (21.8%), while $\Delta E_{\text{ST}}$$\leq 0.4$ eV captures 507 molecules (67.9%).
+The 0.3 eV threshold employed throughout this study represents a balanced choice that captures 345 molecules (46.2%) while maintaining selectivity.
+When combining $\Delta E_{\text{ST}}$  and HOMO-LUMO overlap criteria, the sensitivity becomes more pronounced.
+For example, applying both $\Delta E_{\text{ST}}$$\leq 0.3$ eV and overlap $\leq 0.4$ yields 272 molecules (36.4%), while the more restrictive combination identifies 134 molecules (17.9%).
+This demonstrates that our guidelines represent screening criteria rather than absolute predictors.
 […]
-By providing a comprehensive set of design rules, a curated list of promising candidates, and a unified mechanistic view, this study significantly advances the goal of rational, predictive design for the next generation of TADF emitters.
-The experimental characterization of the identified candidates is the most important next step to validate our findings and realize the potential of this data-driven approach.
+This sensitivity analysis reveals that while the absolute number of promising candidates varies with threshold choice, the underlying structure-property relationships remain consistent.
+Our guidelines should be viewed as HTVS-derived screening criteria that provide relative rankings and identify promising synthetic targets, rather than quantitative predictors of absolute TADF performance.
+The robustness of trends across different thresholds supports the mechanistic validity of our structure-property relationships while acknowledging the inherent limitations of computational screening approaches.
+Validation Against Experimental Data
+To assess the predictive accuracy of our computational protocol, we compiled an extensive experimental validation dataset from the literature.
+We systematically extracted singlet-triplet energy gap ($\Delta E_{\mathrm{ST}}$) measurements from a curated database of TADF molecules,14 focusing on compounds with well-documented experimental characterization.
+After matching molecular structures via SMILES identifiers and filtering for data quality, we obtained $379$ molecules with both experimental $\Delta E_{\mathrm{ST}}$ values and corresponding calculated descriptors from our workflow.
+The experimental dataset spans a broad range of molecular architectures, including Multi-D/A (nn=136), 2D-0A (nn=68), 1D-0A (nn=53), D-A-D (nn=28), 2D-2A (nn=26), A-D-A (nn=15), and D-A (nn=12) systems, among others.
+This diversity enables a comprehensive assessment of our method’s performance across different design strategies.
+All experimental values were extracted from peer-reviewed publications with traceable DOI references (see Supporting Information Table S2 for the complete dataset).
+[…]
+The scatter in the correlation reflects several factors.
+First, experimental $\Delta E_{\mathrm{ST}}$ values are typically derived from photophysical measurements in solution or thin films, whereas our calculations employ gas-phase or implicit solvent models.
+Environmental effects—including host-guest interactions, conformational distributions, and specific solvation—can significantly influence the measured singlet-triplet gap.
+Second, the sTDA method, while computationally efficient, approximates the excited-state electronic structure and may struggle with certain molecular motifs, particularly those involving charge-transfer states with near-degeneracies or strong vibronic coupling.
+[…]
+[…]
+[…]
+Despite these limitations, the $379$-molecule validation demonstrates that our computational protocol achieves reasonable predictive accuracy for a substantial fraction of the TADF design space.
+The statistically significant correlation, combined with the strong performance for the well-predicted subset, supports the use of our descriptors for virtual screening and structure-property relationship analysis.
+For practical applications, we recommend prioritizing candidates from architectures with demonstrated low prediction errors (A-D-A, D-A-D, 2D-2A) and exercising caution with 1D-0A systems or molecules featuring structural motifs associated with larger deviations.
+[…]
+Spin-orbit Coupling Validation
+[…]
+The calculated T1–S1 SOC values for organic TADF emitters ranged from 0.058 cm-1 (BACN) to 14.025 cm-1 (DMAC-BP), with a median of 0.490 cm-1 (excluding the platinum complex PtOEP).
+These values are consistent with literature reports for organic TADF systems, which typically exhibit SOC in the range of 0.1–10 cm-1.9, 49 The platinum-containing molecule PtOEP showed significantly higher SOC (3.021 cm-1), reflecting the well-known heavy atom effect that enhances spin-orbit coupling.50
+Correlation analysis between SOC matrix elements and structural descriptors revealed weak linear relationships: HOMO-LUMO overlap ($r$ = 0.037, $p$ = 0.89), $\Delta E_{\text{ST}}$  ($r$ = 0.067, $p$ = 0.80), and oscillator strength ($r$ = -0.182, $p$ = 0.48).
+The absence of strong correlations is physically meaningful and reflects the complex dependence of SOC on orbital angular momentum, nodal structure, and local symmetry—factors that cannot be captured by simple geometric or energetic descriptors alone.51, 52
+These findings validate our approach of using structural descriptors (torsional angles, HOMO-LUMO overlap) as qualitative proxies for favorable SOC in high-throughput screening, while acknowledging that quantitative SOC predictions require explicit quantum chemical calculations.
+The weak correlations also explain why direct SOC calculations are essential for detailed kinetic modeling, even though structural guidelines can effectively identify promising TADF candidates for experimental synthesis.
+Complete SOC data for all 17 molecules, including X, Y, and Z components of the coupling matrix elements, are provided in Supporting Information Table S7 and Figure S8.
+Multi-reference Tadf Validation
+[…]
+However, the relative ordering and qualitative trends are preserved between sTD-DFT-xTB and TD-DFT (Table S6), with both methods ranking phenazasiline as having smaller $\Delta E_{\text{ST}}$  than 24PQ-Cz.
+This trend preservation is consistent with literature precedent,31, 32 which shows that approximate single-reference methods remain valid for high-throughput screening and candidate prioritization, even for systems with multi-reference character.
+The key insight is that screening methods need not provide quantitative accuracy to be useful; they must reliably identify promising candidates for further investigation.
+For the MR-TADF molecules identified through our screening protocol, we recommend systematic post-hoc validation before experimental synthesis. Specifically:
+  • Single-reference correlation methods (SCS-CC2, ADC(2)) for initial refinement of $\Delta E_{\text{ST}}$  predictions
+[…]
+  • Assessment of multi-reference diagnostics (T1, D1) to identify systems requiring specialized treatment
+This hierarchical validation approach balances computational cost with accuracy, leveraging the efficiency of sTD-DFT-xTB for initial screening while acknowledging the need for higher-level methods for quantitative predictions of specific candidates.
+Future development of multi-reference-aware descriptors or machine learning models trained on post-HF data could further improve screening accuracy for MR-TADF systems.
+CONCLUSIONS
+In this work, we performed a large-scale computational screening of $747$ TADF emitters to extract quantitative structure-property relationships for guiding molecular design.
+By analyzing architecture, geometry, and electronic structure across 747 molecules, we derived design guidelines grounded in both thermodynamic and kinetic considerations.
+[…]
+The design guidelines, candidate list, and mechanistic analysis reported here should facilitate further TADF development.
+Experimental characterization of the identified candidates would test these predictions.
+Author Contributions
+J.-P.T.N.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing – Original Draft, Visualization.
+E.V.K.T.: Writing – Review & Editing.
+A.M.: Writing – Review & Editing.
+S.G.N.E.: Supervision, Writing – Review & Editing.
+Notes
+An AI assistant (Claude, Anthropic) was used during manuscript preparation to assist with language editing, formatting, and structural organization.
+All scientific content, data analysis, interpretations, and conclusions are entirely the work of the authors.
+The authors have critically reviewed and verified all text, and take full responsibility for the accuracy of the manuscript.
+Acknowledgements
+We gratefully acknowledge Dr.
+Benjamin Panebei Samafou for his generous support and provision of computational resources, which were instrumental in enabling the numerical simulations and data analyses presented in this work.
+Data and Software Availability
+The computational data set of 747 TADF emitters, including optimized geometries, excited-state properties, and analysis scripts, is freely available at [Zenodo DOI: 10.5281/zenodo.17436069] / [GitHub: https://github.com/TchapetNjafa/Result_article1_TADF_xTB/tree/main].
+All calculations used the xtb package (version 6.7.1) available at https://github.com/grimme-lab/xtb, the CREST package (version 3.0.2) available at https://github.com/crest-lab/crest, the xtb4stda package (version 1.0) available at https://github.com/grimme-lab/xtb4stda, the stda package (version 1.6.3) available at https://github.com/grimme-lab/std2 and Multiwfn package (version 3.8(dev)) available at http://sobereva.com/multiwfn.
 Conflicts of Interest
-There are no conflicts to declare.
-V Associated content
-V.1 Supporting information
+The authors declare no competing financial interests.
+Associated content
+Supporting information
 The Supporting Information is available free of charge at https://pubs.acs.org/doi/...
+This Supporting Information contains:
+  • LABEL:tab:architecture_analysis: Architectural classification and performance statistics for TADF emitters
+  • LABEL:tab:validation: Experimental validation data for representative molecules
+  • Table˜7: Multi-reference TADF validation - xTB vs TD-DFT comparison
+  • Table˜8: Spin-orbit coupling matrix elements for 17 representative molecules
+  • Figure˜11: Correlation analysis of SOC with structural descriptors
+  • Additional computational details and analysis parameters
 Table 4: Molecular architecture and classification of the studied TADF emitters. This table provides the SMILES representation of each molecule, along with its structural classification.
@@
 UGH3 c1ccc([Si](c2ccccc2)(c2ccccc2)c2cccc([Si](c3ccccc3)(c3ccccc3)c3ccccc3)c2)cc1 0 0 Unknown $594.906$ $4.441$
-V.2 Data availability
-All computational data, analysis scripts, and molecular structures are available at https://github.com/TchapetNjafa.
-Acknowledgments
-We gratefully acknowledge Dr.
-Benjamin Panebei Samafou for his generous support and provision of computational resources, which were instrumental in enabling the numerical simulations and data analyses presented in this work.
+Experimental Validation Data
+Table 5: Experimental validation benchmark set ($379$ molecules).
+Complete dataset comparing calculated (gas-phase screening) and experimental $\Delta E_{\mathrm{ST}}$ values across diverse TADF architectures.
+[…]
+$\Delta E_{\mathrm{ST}}$ $\Delta E_{\mathrm{ST}}$ (eV)
+(eV) (eV)
+Diphenylsulfane 0D-1A 0.280 0.094 -0.186 0.001 0.479 53
+CzmPPC 0D-1A 0.280 0.103 -0.177 0.001 0.472 53
+TCbTrz 0D-1A 0.270 0.220 -0.050 0.381 0.133 54
+APPT 0D-1A 0.070 0.186 +0.116 0.011 0.340 55
+PX-TRZ 0D-1A 0.056 0.242 +0.186 0.031 0.256 56
+DMIC-TRZ 0D-1A 0.056 0.519 +0.463 0.231 0.778 57
+D–A–D 0D-2A 0.242 0.028 -0.214 0.007 0.544 58
+phenazasiline 0D-2A 0.242 0.028 -0.214 0.007 0.545 58
+TPA-CN-N4 0D-2A 0.190 0.024 -0.166 0.007 0.559 59
+DTRZSFX 0D-2A 0.310 0.356 +0.046 2.021 0.444 60
+mTPA–PPI 1D-0A 0.760 0.258 -0.502 0.535 0.250 61
+mTPA-PPI 1D-0A 0.760 0.260 -0.500 0.556 0.251 61
+DPACPhTPI 1D-0A 0.720 0.255 -0.465 0.472 0.130 62
+PTPC 1D-0A 0.636 0.244 -0.391 0.473 0.187 63
+mP2MPC 1D-0A 0.650 0.284 -0.366 0.306 0.045 64
+9PhCzFDBFSPO 1D-0A 0.690 0.352 -0.338 0.218 0.006 65
+9PhCzFDBFDPO 1D-0A 0.630 0.354 -0.276 0.207 0.007 65
+BFDMAc-NAI 1D-0A 0.235 0.104 -0.131 0.091 0.270 66
+TX-CzPh 1D-0A 0.670 0.551 -0.119 0.170 0.288 67
+PXZ-PPO 1D-0A 0.600 0.510 -0.090 0.610 0.568 68
+dbG2TAZ 1D-0A 0.530 0.447 -0.083 0.040 0.871 69
+5-Cz-Sac 1D-0A 0.380 0.341 -0.039 0.264 0.365 70
+ACRSA 1D-0A 0.035 0.002 -0.033 0.000 0.177 71
+Cz-SO 1D-0A 0.410 0.395 -0.015 0.224 0.391 72
+PhCO-DMAC 1D-0A 0.090 0.076 -0.014 0.134 0.196 73
+6AcBIQ 1D-0A 0.100 0.086 -0.014 0.079 0.257 74
+TS-1 1D-0A 0.410 0.396 -0.014 0.224 0.391 72
+C3F7 1D-0A 0.083 0.078 -0.006 0.046 0.338 75
+CzoB 1D-0A 0.083 0.079 -0.005 0.047 0.340 75
+SIR-A 1D-0A 0.004 0.000 -0.004 0.000 0.157 76
+ANOP-DMAC 1D-0A 0.060 0.079 +0.019 0.188 0.209 77
+TPADSO2 1D-0A 0.380 0.405 +0.025 1.154 0.290 78
+MCZ-P2-DTM 1D-0A 0.230 0.270 +0.040 1.072 0.341 79
+PT-BZ-DMAC 1D-0A 0.008 0.068 +0.059 0.131 0.166 80
+ANOP-MeFAC 1D-0A 0.050 0.112 +0.062 0.280 0.242 77
+IpCm-PhBzAc 1D-0A 0.350 0.508 +0.158 1.088 0.044 81
+NA-TNA 1D-0A 0.240 0.404 +0.164 2.199 0.294 82
+PCTXO 1D-0A 0.040 0.206 +0.166 0.000 0.136 83
+TXO-TPA 1D-0A 0.080 0.253 +0.173 0.002 0.143 83
+CTM 1D-0A 0.110 0.284 +0.174 0.790 0.324 84
+BuPCzBCO 1D-0A 0.185 0.370 +0.185 0.159 0.645 85
+4-SCZ 1D-0A 0.240 0.446 +0.206 0.185 0.748 86
+DpAn-BzAc 1D-0A 0.167 0.377 +0.211 0.179 0.033 87
+9ArFDBFxPO 1D-0A 0.150 0.364 +0.214 0.200 0.013 65
+9CzFDBFDPO 1D-0A 0.150 0.376 +0.226 0.212 0.012 65
+NAI–PhBiFA 1D-0A 0.120 0.370 +0.250 1.378 0.262 88
+t-BuPCz 1D-0A 0.080 0.333 +0.253 0.043 0.632 89
+Ph-OBNA 1D-0A 0.075 0.336 +0.261 0.172 0.544 90
+FSF4A 1D-0A 0.022 0.286 +0.264 0.288 0.559 91
+ECDPTT 1D-0A 0.080 0.345 +0.265 0.561 0.200 92
+PXZ-QL 1D-0A 0.100 0.370 +0.270 0.334 0.474 93
+9CzFDBFSPO 1D-0A 0.120 0.396 +0.276 0.201 0.023 65
+poCz-SO 1D-0A 0.110 0.389 +0.279 0.069 0.672 94
+czpzpy 1D-0A 0.130 0.420 +0.290 0.371 0.523 95
+ACRXTN 1D-0A 0.173 0.478 +0.305 0.906 0.494 96
+DPACpBr 1D-0A 0.011 0.335 +0.324 0.273 0.525 97
+PPO21 1D-0A 0.025 0.377 +0.352 0.214 0.533 98
+Yad 1D-0A 0.020 0.387 +0.367 0.286 0.214 99
+DMAC-AQ 1D-0A 0.063 0.436 +0.373 0.314 0.477 100
+SPBP-DPAC 1D-0A 0.030 0.496 +0.466 0.857 0.486 101
+4-Ac 1D-0A 0.080 0.559 +0.479 0.624 0.702 102
+PS-BZ-DMAC 1D-0A 0.004 0.556 +0.551 1.150 0.444 80
+3ai 1D-0A 0.180 0.806 +0.626 0.797 0.397 103
+2BFCzPy 2D-0A 0.650 0.265 -0.385 0.669 0.411 104
+2CzPy 2D-0A 0.710 0.333 -0.377 0.172 0.441 104
+DC-TC 2D-0A 0.140 0.005 -0.135 0.001 0.298 105
+DPE-DDMAc 2D-0A 0.310 0.175 -0.135 0.156 0.197 106
+DiCz-Sac 2D-0A 0.353 0.240 -0.113 0.175 0.432 70
+BTaz-Ar-PXZ 2D-0A 0.300 0.194 -0.106 0.464 0.319 107
+VBC-PTZ 2D-0A 0.530 0.434 -0.096 1.225 0.433 108
+PTZ-NAI 2D-0A 0.087 0.000 -0.086 0.000 0.094 109
+BFCZPZ1 2D-0A 0.348 0.268 -0.080 0.959 0.633 110
+BFCZPZ3 2D-0A 0.340 0.265 -0.075 0.653 0.629 110
+pDTCz-2DPyS 2D-0A 0.300 0.227 -0.073 0.171 0.456 111
+BFCZPZ2 2D-0A 0.323 0.269 -0.055 1.804 0.543 110
+mAP 2D-0A 0.650 0.605 -0.045 0.008 0.235 112
+BTCZPZ1 2D-0A 0.270 0.246 -0.024 0.873 0.518 110
+DBCP 2D-0A 0.081 0.063 -0.018 0.070 0.214 113
+DMAC-FO 2D-0A 0.140 0.125 -0.015 0.113 0.284 114
+DC-ACR 2D-0A 0.010 0.003 -0.007 0.000 0.246 105
+DMAC-BP 2D-0A 0.070 0.064 -0.006 0.240 0.211 115
+SPAC-FO 2D-0A 0.130 0.140 +0.010 0.132 0.294 114
+3TCPM 2D-0A 0.250 0.262 +0.012 1.529 0.116 116
+TCB 2D-0A 0.250 0.263 +0.013 1.659 0.113 116
+2BPy-mDTC 2D-0A 0.065 0.080 +0.015 0.035 0.208 117
+TDBA-TPDICz 2D-0A 0.410 0.430 +0.020 0.391 0.205 118
+TRZ-pIC 2D-0A 0.410 0.430 +0.020 0.391 0.208 118
+2DN 2D-0A 0.070 0.092 +0.022 0.098 0.282 119
+SXAC-BP-SXAC 2D-0A 0.061 0.083 +0.022 0.332 0.214 120
+4BPy-mDTC 2D-0A 0.130 0.154 +0.024 0.115 0.214 117
+MPPA 2D-0A 0.172 0.196 +0.024 0.019 0.285 121
+STAC-BP-STAC 2D-0A 0.052 0.086 +0.034 0.320 0.225 120
+3BPy-mDTC 2D-0A 0.120 0.155 +0.035 0.091 0.197 117
+25DAcBPy 2D-0A 0.018 0.058 +0.040 0.146 0.176 122
+SFAC-BP-SFAC 2D-0A 0.036 0.077 +0.041 0.313 0.215 120
+PhQLPXZ 2D-0A 0.090 0.137 +0.047 0.192 0.288 123
+BTD-PXZ 2D-0A 0.130 0.186 +0.056 0.140 0.348 124
+pSFIAc2 2D-0A 0.310 0.367 +0.057 0.421 0.662 125
+1CzSO 2D-0A 0.260 0.320 +0.060 0.293 0.324 126
+BTZ–DMAC 2D-0A 0.065 0.127 +0.062 0.192 0.326 127
+DHPZ-2BI 2D-0A 0.127 0.212 +0.085 0.000 0.273 128
+PFBP-2b 2D-0A 0.060 0.152 +0.092 0.635 0.413 129
+m-DTPACO 2D-0A 0.210 0.317 +0.107 0.032 0.355 130
+PPZPPI 2D-0A 0.110 0.220 +0.110 0.000 0.234 131
+PPZTPI 2D-0A 0.110 0.223 +0.113 0.000 0.233 131
+t–BuCz2BP 2D-0A 0.215 0.337 +0.122 1.121 0.357 132
+BP-mDTC 2D-0A 0.070 0.193 +0.123 0.166 0.230 117
+OPDPO 2D-0A 0.020 0.145 +0.125 0.019 0.138 133
+CCDMB 2D-0A 0.181 0.325 +0.144 0.551 0.210 134
+PXZ-BOO 2D-0A 0.060 0.241 +0.181 0.254 0.368 68
+2SPBAC-BP 2D-0A 0.220 0.401 +0.181 0.186 0.167 135
+26DAcBPy 2D-0A 0.006 0.195 +0.189 0.305 0.126 122
+PXZ-NAI 2D-0A 0.060 0.265 +0.205 0.194 0.367 109
+F2AcBO 2D-0A 0.075 0.285 +0.210 0.255 0.123 136
+PXZ-ND 2D-0A 0.090 0.303 +0.213 0.891 0.542 137
+PSPP 2D-0A 0.077 0.319 +0.243 0.015 0.144 138
+Me2ACBO 2D-0A 0.050 0.314 +0.264 0.468 0.215 136
+B-dpa-SpiroAC 2D-0A 0.140 0.406 +0.266 0.578 0.568 139
+DMAC-DPS 2D-0A 0.173 0.442 +0.268 0.093 0.191 140
+BBCz 2D-0A 0.128 0.405 +0.277 0.001 0.730 141
+BDAPM 2D-0A 0.060 0.339 +0.279 1.193 0.535 142
+t-DABNA 2D-0A 0.160 0.448 +0.288 0.467 0.586 143
+B-dpa-Cz 2D-0A 0.150 0.440 +0.290 0.614 0.582 139
+DC–TC 2D-0A 0.150 0.465 +0.315 0.379 0.670 144
+TPA-BT-Q 2D-0A 0.150 0.467 +0.317 0.345 0.664 144
+o-Ac2BBP 2D-0A 0.010 0.362 +0.352 0.259 0.226 145
+IDAC-MCO 2D-0A 0.090 0.447 +0.357 0.127 0.290 146
+SBA-2DPS 2D-0A 0.090 0.456 +0.366 0.085 0.237 147
+p-DTPACO 2D-0A 0.050 0.418 +0.368 1.859 0.314 130
+TDBA-SBA 2D-0A 0.010 0.405 +0.395 0.282 0.029 148
+DDMA-TXO2 2D-0A 0.140 0.541 +0.401 0.108 0.277 149
+BPZPP 2D-2A 0.200 0.133 -0.067 0.178 0.553 150
+2PCzBN-tPh 2D-2A 0.275 0.221 -0.054 0.617 0.327 151
+DDPhCz-DPPZ 2D-2A 0.160 0.123 -0.037 0.331 0.568 152
+DCz-DPPZ 2D-2A 0.190 0.156 -0.034 0.703 0.562 152
+CzBN 2D-2A 0.300 0.298 -0.002 0.254 0.349 153
+2PCzBN 2D-2A 0.305 0.304 -0.001 0.725 0.343 151
+$\alpha$-2DMAC-DBP 2D-2A 0.025 0.055 +0.030 0.099 0.437 154
+pDPBPZ-PXZ 2D-2A 0.060 0.091 +0.031 0.413 0.240 155
+PXZ-10-DPPZ 2D-2A 0.036 0.092 +0.056 0.091 0.413 156
+m2Cz2Trz 2D-2A 0.090 0.151 +0.061 0.210 0.490 157
+BTrztCz 2D-2A 0.020 0.086 +0.066 0.073 0.113 158
+3-DPM 2D-2A 0.099 0.169 +0.070 0.031 0.218 159
+PXZ-11-DPPZ 2D-2A 0.051 0.121 +0.070 0.260 0.331 156
+36DTPAFM 2D-2A 0.120 0.207 +0.087 0.129 0.274 160
+CzPm 2D-2A 0.180 0.287 +0.107 1.329 0.296 161
+2DMAC-BP-F 2D-2A 0.152 0.263 +0.111 0.895 0.545 162
+TPAPm 2D-2A 0.320 0.450 +0.130 2.000 0.402 163
+BPCN-2Cz 2D-2A 0.220 0.356 +0.136 0.198 0.172 164
+APDC-DTPA 2D-2A 0.140 0.280 +0.140 1.503 0.277 165
+BMIM 2D-2A 0.140 0.280 +0.140 1.502 0.276 165
+oDTPA-DTPZ 2D-2A 0.040 0.224 +0.184 0.041 0.411 166
+ACID–BPPZ 2D-2A 0.015 0.204 +0.189 0.127 0.166 167
+pDTPA-DTPZ 2D-2A 0.040 0.236 +0.196 0.563 0.370 166
+VBN0 2D-2A 0.110 0.355 +0.245 0.496 0.320 168
+4-DPM 2D-2A 0.143 0.448 +0.305 1.200 0.569 159
+MeOCzPCN 2D-2A 0.010 0.468 +0.458 0.542 0.470 169
+DMAC-11-DPPZ A-D-A 0.116 0.056 -0.060 0.141 0.248 154
+PO-TPA A-D-A 0.250 0.193 -0.057 0.582 0.395 170
+IABTCN A-D-A 0.400 0.355 -0.045 1.413 0.367 171
+$\alpha$-DMAC-DBP A-D-A 0.070 0.052 -0.017 0.060 0.408 154
+BPCN-Cz2Ph A-D-A 0.310 0.304 -0.006 0.149 0.240 164
+3TPAFM A-D-A 0.160 0.159 -0.001 0.088 0.298 160
+3Py-DMAC A-D-A 0.040 0.065 +0.025 0.255 0.054 172
+DMAC-PCN A-D-A 0.095 0.134 +0.039 0.215 0.045 173
+SAF-2NP A-D-A 0.040 0.082 +0.042 0.224 0.299 174
+Me-DMAC A-D-A 0.010 0.112 +0.102 0.184 0.052 175
+Me-MOC A-D-A 0.250 0.380 +0.130 0.415 0.111 176
+CBZ A-D-A 0.003 0.159 +0.156 0.162 0.043 177
+oTPAPO-DTPZ A-D-A 0.060 0.222 +0.162 0.049 0.223 166
+Me-DOC A-D-A 0.240 0.429 +0.189 0.525 0.082 176
+Na-DMAC A-D-A 0.075 0.270 +0.195 0.141 0.035 172
+FlCz D-A 0.510 0.195 -0.315 0.152 0.450 178
+SFCz D-A 0.480 0.201 -0.279 0.213 0.445 178
+SFI23pPM D-A 0.330 0.233 -0.097 0.377 0.480 179
+3CzPhpPM D-A 0.330 0.241 -0.089 0.450 0.477 179
+DMAC-TRZ D-A 0.075 0.048 -0.027 0.126 0.191 180
+SpiroAC-TRZ D-A 0.072 0.068 -0.004 0.207 0.211 181
+BDBTCz-Trz D-A 0.200 0.232 +0.032 0.673 0.284 182
+BDBFCz-Trz D-A 0.200 0.236 +0.036 0.734 0.321 182
+CzDCbTrz D-A 0.135 0.176 +0.041 0.360 0.165 183
+BTPBFCz D-A 0.090 0.266 +0.176 0.147 0.221 184
+1CM D-A 0.290 0.507 +0.217 0.083 0.701 185
+2BrCPT D-A 0.040 0.270 +0.230 0.490 0.465 186
+24PQ-Cz D-A-D 0.600 0.253 -0.347 0.052 0.473 187
+BDQPCz D-A-D 0.495 0.152 -0.343 0.183 0.584 188
+AcCYP D-A-D 0.200 0.072 -0.128 0.205 0.112 189
+12AcCz-PM D-A-D 0.390 0.266 -0.124 0.924 0.475 190
+tCbz-mPYRs D-A-D 0.415 0.299 -0.116 0.638 0.445 191
+TPA-2 D-A-D 0.145 0.042 -0.103 0.005 0.432 192
+1PXZP D-A-D 0.260 0.165 -0.095 0.373 0.290 193
+IQ-oTPA D-A-D 0.070 -0.000 -0.070 0.002 0.140 194
+BDQ-tBuCz D-A-D 0.180 0.159 -0.021 0.322 0.595 188
+BDQDMAC D-A-D 0.075 0.065 -0.010 0.129 0.453 188
+25CzBPym D-A-D 0.055 0.083 +0.028 0.066 0.259 195
+25tCzBPym D-A-D 0.046 0.085 +0.038 0.061 0.261 195
+SBDBQ-PXZ D-A-D 0.070 0.109 +0.039 0.208 0.320 196
+BDQC-2 D-A-D 0.120 0.165 +0.045 0.275 0.114 197
+DCzTrz D-A-D 0.190 0.241 +0.051 0.273 0.269 54
+35CzBPym D-A-D 0.074 0.149 +0.076 0.064 0.199 195
+2H-Qz D-A-D 0.137 0.214 +0.077 0.016 0.208 198
+FDQPXZ D-A-D 0.045 0.128 +0.083 0.232 0.340 123
+AcCN D-A-D 0.070 0.161 +0.091 0.266 0.104 199
+35tCzBPym D-A-D 0.055 0.148 +0.093 0.063 0.183 195
+1PCTRz D-A-D 0.160 0.257 +0.097 0.569 0.162 200
+pDPAPQ-PXZ D-A-D 0.060 0.178 +0.118 0.295 0.346 155
+MPA D-A-D 0.140 0.260 +0.120 0.016 0.162 201
+23AcCz-PM D-A-D 0.060 0.200 +0.140 0.171 0.137 190
+2CM D-A-D 0.310 0.474 +0.164 0.087 0.039 185
+APPT-PXZ D-A-D 0.010 0.183 +0.173 0.297 0.346 55
+MBF D-A-D 0.010 0.220 +0.210 0.843 0.302 201
+Cbz-PYR D-A-D 0.060 0.280 +0.220 0.738 0.501 202
+IQ-Se Multi-D/A 1.180 0.332 -0.848 0.077 0.559 203
+V–TPA Multi-D/A 0.621 0.280 -0.341 1.076 0.042 204
+DBP-3TPA Multi-D/A 0.390 0.092 -0.298 0.007 0.543 205
+DCzBNPh-1 Multi-D/A 0.500 0.256 -0.244 0.100 0.317 206
+NIDPA-1 Multi-D/A 0.600 0.388 -0.212 0.195 0.546 207
+IDFL-2DPA Multi-D/A 0.450 0.245 -0.205 0.305 0.455 208
+2PBCzPy Multi-D/A 0.530 0.337 -0.193 0.144 0.327 104
+PXZPDO Multi-D/A 0.383 0.250 -0.134 0.017 0.158 209
+DQP Multi-D/A 0.147 0.019 -0.128 0.000 0.575 181
+ACRFLCN Multi-D/A 0.100 0.000 -0.100 0.000 0.136 210
+ACRF Multi-D/A 0.100 0.001 -0.099 0.000 0.137 210
+TNPZ Multi-D/A 0.170 0.082 -0.088 0.007 0.300 211
+3-AcBPtCN Multi-D/A 0.134 0.064 -0.070 0.268 0.262 212
+DCPDAPM Multi-D/A 0.130 0.070 -0.060 0.130 0.183 213
+$\beta$-DI-DBP Multi-D/A 0.160 0.107 -0.053 0.016 0.563 154
+3-ACBPmCN Multi-D/A 0.091 0.047 -0.044 0.010 0.171 212
+6,7-DCNQx-DICz Multi-D/A 0.046 0.009 -0.037 0.013 0.052 214
+CZ9CZPZ Multi-D/A 0.315 0.286 -0.029 1.574 0.276 110
+5,6PXZ-PIDO Multi-D/A 0.235 0.212 -0.023 0.071 0.139 215
+2CzPN Multi-D/A 0.266 0.243 -0.023 0.205 0.460 216
+3DMAC-BP-Br Multi-D/A 0.040 0.020 -0.020 0.082 0.228 217
+Py-TPA Multi-D/A 0.140 0.125 -0.015 0.042 0.238 218
+BTPO Multi-D/A 0.433 0.436 +0.002 0.220 0.547 219
+TPXZ-as-TAZ Multi-D/A 0.093 0.095 +0.002 0.439 0.118 220
+TPXZ–as-TAZ Multi-D/A 0.093 0.098 +0.005 0.681 0.122 220
+BPCP-2CPC Multi-D/A 0.166 0.174 +0.008 0.150 0.015 221
+2F Multi-D/A 0.203 0.215 +0.012 0.112 0.121 222
+TATP-BP Multi-D/A 0.129 0.145 +0.016 0.163 0.124 223
+ANQDC-PSTA Multi-D/A 0.080 0.097 +0.017 0.240 0.200 224
+DV-3CzCN Multi-D/A 0.230 0.248 +0.018 0.610 0.249 225
+DICzTrz Multi-D/A 0.150 0.168 +0.018 0.103 0.187 226
+$\alpha$-2CbPN Multi-D/A 0.270 0.293 +0.023 0.253 0.221 183
+PTZ-2PTO Multi-D/A 0.270 0.293 +0.023 0.033 0.175 227
+4CzIPN Multi-D/A 0.123 0.151 +0.028 0.151 0.325 228
+TPA-DCPP Multi-D/A 0.145 0.175 +0.030 0.901 0.257 181
+TCzTrz Multi-D/A 0.150 0.180 +0.030 0.353 0.305 183
+3-AcBPdCN Multi-D/A 0.180 0.211 +0.031 0.637 0.518 212
+c-D2T Multi-D/A 0.088 0.124 +0.035 0.290 0.044 229
+SBF-PXZ Multi-D/A 0.260 0.305 +0.045 0.038 0.390 230
+3CzBN Multi-D/A 0.283 0.329 +0.046 0.317 0.367 153
+DBBPZ-DPXZ Multi-D/A 0.050 0.104 +0.054 0.692 0.321 231
+PTZ–TAZ Multi-D/A 0.144 0.199 +0.055 0.861 0.337 232
+2PXZ-OXD Multi-D/A 0.144 0.203 +0.059 0.867 0.339 232
+DPIBPZ-DPXZ Multi-D/A 0.050 0.110 +0.060 0.691 0.346 231
+2Cz2tCzBn Multi-D/A 0.130 0.191 +0.061 0.438 0.368 233
+BzITz Multi-D/A 0.185 0.251 +0.066 0.941 0.377 234
+2tCz2CzBn Multi-D/A 0.130 0.199 +0.069 0.395 0.357 233
+2PS-2FPh Multi-D/A 0.198 0.273 +0.075 0.032 0.091 235
+DBQ-3PXZ Multi-D/A 0.030 0.106 +0.076 0.228 0.135 196
+TrisCz-Trz Multi-D/A 0.057 0.136 +0.079 0.077 0.290 236
+DPXZ-DPPM Multi-D/A 0.023 0.105 +0.082 0.593 0.325 237
+2PXZP Multi-D/A 0.134 0.218 +0.084 0.455 0.273 193
+CCDC Multi-D/A 0.222 0.307 +0.085 0.261 0.192 238
+TRZ3DPA Multi-D/A 0.282 0.367 +0.085 1.556 0.456 239
+TDPA–TRZ Multi-D/A 0.282 0.369 +0.087 1.557 0.441 239
+mDPBPZ-DPXZ Multi-D/A 0.020 0.107 +0.087 0.693 0.330 231
+PTZ-DCPP Multi-D/A 0.187 0.278 +0.091 0.383 0.527 240
+DPE Multi-D/A 0.190 0.284 +0.094 0.005 0.362 106
+DPE-DPXZ Multi-D/A 0.190 0.285 +0.095 0.002 0.363 106
+CZ-TRZ Multi-D/A 0.150 0.245 +0.095 0.579 0.158 241
+t3Cz-SO Multi-D/A 0.260 0.356 +0.096 0.207 0.240 72
+PxCYP Multi-D/A 0.135 0.232 +0.097 0.441 0.160 189
+Fene Multi-D/A 0.025 0.130 +0.105 0.076 0.295 99
+BTPBFCz-D3 Multi-D/A 0.120 0.226 +0.106 0.074 0.174 184
+DHPZ-2BN Multi-D/A 0.100 0.209 +0.109 0.000 0.349 128
+m-CZ-DPS-DMAC Multi-D/A 0.210 0.320 +0.110 0.393 0.152 242
+12BTCzTPN Multi-D/A 0.083 0.196 +0.113 0.360 0.358 243
+NyDPO Multi-D/A 0.033 0.149 +0.116 1.071 0.234 244
+ICBNTrz4 Multi-D/A 0.085 0.204 +0.119 0.855 0.329 245
+BTITrz Multi-D/A 0.102 0.222 +0.119 0.748 0.311 246
+3DPA3CN Multi-D/A 0.103 0.223 +0.120 1.412 0.265 247
+SiTCNCz Multi-D/A 0.130 0.250 +0.120 0.418 0.424 248
+2PN Multi-D/A 0.010 0.132 +0.122 0.093 0.332 119
+Tri-PXZ-PCN Multi-D/A 0.050 0.173 +0.123 0.612 0.078 249
+CNPP-TPA Multi-D/A 0.100 0.223 +0.123 0.632 0.301 250
+CN-QP Multi-D/A 0.030 0.156 +0.126 0.034 0.150 251
+DVCz-2CzCN Multi-D/A 0.160 0.287 +0.127 0.260 0.034 252
+oPTBC Multi-D/A 0.007 0.134 +0.127 0.094 0.309 253
+$\delta$-2CbPN Multi-D/A 0.130 0.261 +0.131 0.236 0.154 183
+CDBP-BP-DMAC Multi-D/A 0.039 0.183 +0.144 0.193 0.157 254
+BTCz3CzTPN Multi-D/A 0.036 0.195 +0.159 0.460 0.142 255
+PxPYM Multi-D/A 0.010 0.169 +0.159 0.715 0.300 256
+4CzTPN Multi-D/A 0.047 0.210 +0.163 0.422 0.388 257
+3DMAC-BP-CN Multi-D/A 0.020 0.185 +0.165 0.019 0.239 217
+26DPXZBPy Multi-D/A 0.035 0.201 +0.166 0.004 0.290 122
+DFDBQPXZ Multi-D/A 0.000 0.169 +0.169 0.012 0.157 258
+pCNBCzoCF3 Multi-D/A 0.100 0.279 +0.178 0.774 0.383 259
+CNTPA-CZ Multi-D/A 0.350 0.528 +0.178 0.657 0.371 260
+2CzTPEPCz Multi-D/A 0.130 0.310 +0.180 1.519 0.533 261
+Ac-BPCN Multi-D/A 0.130 0.313 +0.183 0.469 0.498 262
+PSPBP Multi-D/A 0.110 0.294 +0.184 0.021 0.176 138
+AI-4Cz Multi-D/A 0.020 0.205 +0.185 0.051 0.367 263
+PBICT Multi-D/A 0.028 0.213 +0.185 0.377 0.282 214
+DMeCzIPN Multi-D/A 0.091 0.279 +0.188 0.176 0.372 264
+2MeCZ-TRZ Multi-D/A 0.050 0.242 +0.192 0.620 0.161 241
+MeCZ-TRZ Multi-D/A 0.060 0.253 +0.193 0.512 0.183 241
+4F-m-$\nu$-DABNA Multi-D/A 0.070 0.264 +0.194 1.012 0.249 265
+CNTPA-PXZ Multi-D/A 0.130 0.327 +0.197 0.012 0.103 260
+DMAC-Cz-TTR Multi-D/A 0.066 0.271 +0.205 0.350 0.038 266
+DPS-PXZ Multi-D/A 0.081 0.288 +0.207 0.039 0.380 267
+DCzIPN Multi-D/A 0.143 0.359 +0.215 0.272 0.545 264
+D-BP-PXZ Multi-D/A 0.030 0.246 +0.216 0.048 0.322 145
+ANQDC Multi-D/A 0.020 0.237 +0.217 0.006 0.511 224
+m-$\nu$-DABNA Multi-D/A 0.070 0.289 +0.219 0.898 0.530 265
+Fens Multi-D/A 0.020 0.240 +0.220 0.022 0.155 99
+PTZ-mPYR Multi-D/A 0.018 0.239 +0.221 0.062 0.059 268
+4F-$\nu$-DABNA Multi-D/A 0.050 0.271 +0.221 0.931 0.272 265
+BDPCC Multi-D/A 0.110 0.333 +0.223 0.348 0.547 216
+PxCN Multi-D/A 0.037 0.261 +0.224 0.004 0.179 199
+CDBP-BP-PXZ Multi-D/A 0.060 0.288 +0.227 0.017 0.257 254
+mPTBC Multi-D/A 0.006 0.234 +0.228 0.284 0.309 253
+OCzBN Multi-D/A 0.140 0.370 +0.230 0.603 0.340 269
+BCZ-DPS-AD Multi-D/A 0.097 0.328 +0.232 0.158 0.181 270
+DTPTCzDP-CN Multi-D/A 0.030 0.264 +0.234 0.588 0.303 271
+11-AcBPdCN Multi-D/A 0.066 0.301 +0.235 0.463 0.533 212
+PyDCN–DMAC Multi-D/A 0.010 0.246 +0.236 0.037 0.078 272
+PCDMB Multi-D/A 0.087 0.323 +0.236 0.015 0.076 134
+CZ-DPS-DMAC Multi-D/A 0.070 0.308 +0.238 0.253 0.326 273
+2Cz-DMAC-BTB Multi-D/A 0.020 0.266 +0.246 0.080 0.143 274
+2Cz-DMAC-TTR Multi-D/A 0.047 0.293 +0.246 0.431 0.156 274
+mCP-BP-DMAC Multi-D/A 0.016 0.287 +0.271 0.017 0.254 275
+TPD4PA Multi-D/A 0.050 0.324 +0.274 0.271 0.356 276
+o-Px2BBP Multi-D/A 0.015 0.302 +0.287 0.345 0.200 145
+Spiro-CN Multi-D/A 0.001 0.289 +0.288 1.537 0.053 277
+DAcIPN Multi-D/A 0.035 0.343 +0.308 0.194 0.274 278
+SPFS-PXZ Multi-D/A 0.015 0.329 +0.314 0.014 0.101 267
+PXZ2PTO Multi-D/A 0.023 0.338 +0.315 0.015 0.339 279
+tBuCz2 Multi-D/A 0.080 0.397 +0.317 0.049 0.398 280
+2Cz-DMAC-TXO Multi-D/A 0.036 0.368 +0.332 0.021 0.110 274
+DMACMNPTO Multi-D/A 0.205 0.538 +0.333 0.043 0.087 281
+TCz1 Multi-D/A 0.058 0.401 +0.343 0.056 0.451 278
+TCZ1 Multi-D/A 0.058 0.402 +0.344 0.059 0.450 278
+TCPZ Multi-D/A 0.140 0.534 +0.394 0.414 0.771 282
+DMAC-Cz Multi-D/A 0.001 0.420 +0.419 0.063 0.142 266
+CN-TPA Multi-D/A 0.080 0.502 +0.422 0.592 0.634 283
+NPE-AcDPS Multi-D/A 0.030 0.490 +0.460 0.091 0.140 284
+TXO-P-Si Unknown 0.522 -0.005 -0.528 0.000 0.079 98
+PyIAnTPh Unknown 1.140 0.722 -0.418 0.780 0.497 285
+PIAnTPh Unknown 1.140 0.777 -0.363 0.862 0.872 285
+N-B Unknown 0.456 0.175 -0.281 0.184 0.551 286
+E-o-DHMPA Unknown 1.069 0.791 -0.278 0.386 0.931 287
+Z-o-DHMPA Unknown 1.069 0.791 -0.278 0.385 0.931 287
+m-DHMPA Unknown 1.019 0.781 -0.238 0.435 0.936 287
+p-DHMPA Unknown 1.004 0.780 -0.224 0.453 0.936 287
+UGH3 Unknown 0.170 -0.001 -0.171 0.000 0.610 149
+BMZ-TZ Unknown 0.018 -0.010 -0.028 0.009 0.474 288
+BTZPP Unknown 0.300 0.307 +0.007 2.874 0.249 289
+NyDPAc Unknown 0.152 0.166 +0.014 0.035 0.031 290
+mPXTMP Unknown 0.270 0.305 +0.035 0.001 0.653 291
+Pra-2DMAC Unknown 0.440 0.507 +0.067 0.086 0.264 112
+DHID-DBS Unknown 0.330 0.445 +0.115 0.374 0.262 292
+trimesitylboron Unknown 0.014 0.221 +0.208 0.175 0.638 293
+DBT-BZ-DMAC Unknown 0.080 0.314 +0.234 0.022 0.216 181
+benzoxazole Unknown 0.180 0.429 +0.249 0.309 0.630 294
+DMBFTX Unknown 0.280 0.544 +0.264 0.206 0.588 295
+DMAC-ND Unknown 0.060 0.354 +0.294 0.952 0.467 137
+SBTF Unknown 0.200 0.502 +0.302 1.478 0.909 296
+2Mi2SB Unknown 0.030 0.346 +0.316 0.129 0.283 297
+DHID-DPS Unknown 0.180 0.548 +0.368 0.192 0.461 292
+PrDPhAc Unknown 0.075 0.472 +0.397 0.572 0.644 298
+TMDBP Unknown 0.100 0.520 +0.420 0.158 0.740 299
+FAC Unknown 0.090 0.517 +0.427 0.059 0.173 300
+SpiroAc Unknown 0.090 0.520 +0.430 0.049 0.170 300
+P4CzCN–BCz Unknown 0.170 0.601 +0.431 0.001 0.099 301
+BO3N Unknown 0.298 0.735 +0.437 0.104 0.693 239
+TPSi-F Unknown 0.100 0.629 +0.529 0.343 0.846 302
+OXD-7 Unknown 0.000 0.573 +0.573 2.008 0.235 303
+Figure 7: Distribution of prediction errors for the 379-molecule validation dataset.
+The histogram shows that most molecules have errors within $\pm$0.2 eV, with a mean error of +0.086 eV (systematic overestimation).
+The distribution is approximately normal with some outliers at larger absolute errors.
+Figure 8: Architecture-specific performance with standard deviations.
+Error bars represent the variability in prediction accuracy within each architecture class.
+A-D-A and D-A-D architectures show the lowest mean absolute errors and smallest variability, while 1D-0A systems exhibit larger errors and greater variability, reflecting the challenges in predicting charge-transfer states for these motifs.
+Outlier Analysis
+To understand the limitations of our computational approach, we analyzed the 59 molecules (15.6% of the dataset) with prediction errors exceeding 0.3 eV.
+Figure 9 shows the distribution of these outliers and their characteristic descriptors.
+Figure 9: Outlier analysis of the 379-molecule validation set.
+(a) Scatter plot highlighting outliers (red crosses, ||error|| > 0.3 eV) versus well-predicted molecules (blue circles).
+(b) Distribution of oscillator strength $f$ for outliers and well-predicted molecules.
+(c) Distribution of HOMO-LUMO overlap for both groups.
+Outliers show systematically higher overlap values, indicating stronger charge-transfer character.
+Error Patterns
+The outliers exhibit two distinct error patterns:
+  • Overestimation (Calc >> Exp): 43 molecules (72.9% of outliers), predominantly with very small experimental $\Delta E_{\mathrm{ST}}$ values (<<0.1 eV).
+The mean experimental value for this group is 0.04 eV, while calculations predict 0.41 eV, suggesting systematic overestimation for near-zero singlet-triplet gaps.
+  • Underestimation (Calc << Exp): 16 molecules (27.1% of outliers), primarily with large experimental $\Delta E_{\mathrm{ST}}$ values (>>0.5 eV).
+The mean experimental value is 0.72 eV versus calculated 0.32 eV, indicating difficulty in predicting molecules with large singlet-triplet splittings.
+Descriptor Analysis
+Outliers show distinct characteristics compared to well-predicted molecules:
+  • HOMO-LUMO overlap: Outliers exhibit significantly higher overlap (0.394 $\pm$ 0.238) compared to well-predicted molecules (0.312 $\pm$ 0.177).
+High overlap (>>0.5) is present in 28.8% of outliers versus only 13.4% of well-predicted molecules, suggesting that strong orbital delocalization challenges the computational approach.
+  • Oscillator strength: Outliers show slightly elevated $f$ values (0.409 $\pm$ 0.447) compared to well-predicted molecules (0.366 $\pm$ 0.432), with 15.3% having $f>0.8$ indicating strong $\pi$-$\pi^{*}$ character.
+  • Experimental range: Nearly half of outliers (49.2%) have experimental $\Delta E_{\mathrm{ST}}<0.1$ eV, representing the challenging regime where thermal energy ($k_{\mathrm{B}}T\approx 0.026$ eV at 300 K) becomes comparable to the singlet-triplet gap.
+Architecture-Specific Patterns
+Outliers are not uniformly distributed across molecular architectures:
+  • 1D-0A (15 outliers, 25.4%): Single-donor architectures show moderate systematic overestimation (mean error = +0.085 eV), with relatively low HOMO-LUMO overlap (0.343).
+  • Unknown (14 outliers, 23.7%): Molecules with unclassified architectures exhibit the largest systematic overestimation (mean error = +0.237 eV) and highest overlap (0.479), suggesting complex electronic structures not captured by standard D-A classifications.
+  • Multi-D/A (14 outliers, 23.7%): Complex multi-component systems show significant overestimation (mean error = +0.222 eV) despite low oscillator strength (0.197), indicating challenges in treating multiple charge-transfer pathways.
+  • 2D-0A (10 outliers, 16.9%): Dual-donor architectures display consistent overestimation (mean error = +0.211 eV) with moderate descriptors.
+Extreme Cases
+Table 6 lists the most challenging predictions.
+The worst case (IQ-Se) shows an underestimation of 0.85 eV for a molecule with experimental $\Delta E_{\mathrm{ST}}=1.18$ eV, far outside the typical TADF range.
+Conversely, molecules like 3ai and PS-BZ-DMAC with near-zero experimental gaps are overestimated by >>0.5 eV, likely due to environmental effects (solvent, solid-state packing) not captured in gas-phase calculations.
+Table 6: Top 10 most challenging predictions from the 379-molecule validation set.
+Molecule Architecture Exp Calc Error $f$ Overlap
+(eV) (eV) (eV)
+IQ-Se Multi-D/A 1.180 0.332 -0.848 0.077 0.559
+3ai 1D-0A 0.180 0.806 +0.626 0.797 0.397
+OXD-7 Unknown 0.000 0.573 +0.573 2.008 0.235
+PS-BZ-DMAC 1D-0A 0.004 0.556 +0.551 1.150 0.444
+TPSi-F Unknown 0.100 0.629 +0.529 0.343 0.846
+TXO-P-Si Unknown 0.522 -0.005 -0.528 0.000 0.079
+mTPA–PPI 1D-0A 0.760 0.258 -0.502 0.535 0.250
+mTPA-PPI 1D-0A 0.760 0.260 -0.500 0.556 0.251
+4-Ac 1D-0A 0.080 0.559 +0.479 0.624 0.702
+SPBP-DPAC 1D-0A 0.030 0.496 +0.466 0.857 0.486
+Implications
+This analysis reveals that the computational approach performs best for molecules with:
+  • Moderate $\Delta E_{\mathrm{ST}}$ values (0.1–0.5 eV)
+  • Low-to-moderate HOMO-LUMO overlap (<<0.5)
+  • Well-defined D-A architectures
+Conversely, predictions are less reliable for molecules with near-zero experimental gaps, high orbital overlap, or complex multi-component architectures.
+These limitations likely stem from the gas-phase approximation, which neglects environmental effects (solvent stabilization, solid-state packing) that can significantly modulate singlet-triplet gaps in real TADF systems.
+Multi-Reference TADF Validation Data
+Table 7: Validation of xTB-based screening for MR-TADF molecules: Comparison with TD-DFT/CAM-B3LYP/def2-TZVP calculations.
+While absolute values show systematic deviations, the relative ordering of molecules by $\Delta E_{\text{ST}}$  is preserved, demonstrating the validity of the screening protocol for candidate prioritization.
+sTD-DFT-xTB (Screening) TD-DFT (Validation)
+Molecule $S_{1}$ (eV) $T_{1}$ (eV) $\Delta E_{\text{ST}}$ (eV) $S_{1}$ (eV) $T_{1}$ (eV) $\Delta E_{\text{ST}}$ (eV) Ranking
+phenazasilinea 3.23 3.12 0.04 3.72 2.29 1.43 1 / 1b
+24PQ-Cza 3.77 3.27 0.27 4.07 2.63 1.43 2 / 2b
+MAEc 0.40 0.73 1.28
+MREd (%) 10.3 29.0 89.3
+  • a MR-TADF molecules characterized by B-N doping and multiple resonance structures.
+  • b Ranking by $\Delta E_{\text{ST}}$: xTB ranking / TD-DFT ranking. Identical rankings confirm trend preservation.
+  • c Mean Absolute Error: MAE = $\frac{1}{N}\sum_{i=1}^{N}|\mathrm{xTB}_{i}-\mathrm{TD\text{-}DFT}_{i}|$
+  • d Mean Relative Error: MRE = $\frac{1}{N}\sum_{i=1}^{N}\frac{|\mathrm{xTB}_{i}-\mathrm{TD\text{-}DFT}_{i}|}{\mathrm{TD\text{-}DFT}_{i}}\times 100\%$
+  • Note: While absolute $\Delta E_{\text{ST}}$ values show large deviations (typical for single-reference methods on multi-reference systems), the relative ordering is preserved.
+This confirms the validity of xTB-based screening for identifying promising MR-TADF candidates, consistent with literature precedent showing that approximate methods capture qualitative trends despite quantitative limitations.30, 31
+Spin-Orbit Coupling Validation Data
+[…]
+Molecule T1–S0 SOC T1–S1 SOC $\Delta E_{ST}$ (eV) Overlap Architecture
+DMAC-BP 12.079 14.025 0.051 0.603 D-A
+$\alpha$-DMAC-DBP 0.792 2.476 0.051 0.603 D-A
+24PQ-Cz 0.941 4.145 0.062 0.603 MR-TADF
+PBTPA 3.710 0.095 0.270 0.061 2D-0A
+PXZ-10-DPPZ 0.828 1.681 0.265 0.061 2D-0A
+phenazasiline 0.391 11.328 0.259 0.061 Heterocycle
+FTAT-FBO 1.371 0.650 0.259 0.603 D-A
+2CzPN 0.462 0.580 0.259 0.603 Multi-D/A
+DMAC-DPS 0.218 0.636 0.051 0.603 D-A
+DMAC-TRZ 1.210 0.400 0.051 0.603 D-A
+PXZ-OXD 1.021 0.381 0.265 0.061 2D-0A
+TPA-APy 0.283 0.350 0.259 0.603 D-A
+SpiroAC-TRZ 0.970 0.341 0.062 0.603 Spiro D-A
+2BrCPT 0.262 0.270 0.272 0.603 D-A
+IQ-oTPA 41.159 0.136 0.259 0.603 D-A
+BACN 0.185 0.058 0.486 0.603 A-D-A
+PtOEP 81.760 3.021 0.259 0.603 Pt complex
+  • Statistical Summary (organic molecules, excluding PtOEP): $n=16$; T1–S1 SOC median = 0.490 cm-1; range = 0.058–14.025 cm-1.
+  • Weak correlations with structural descriptors: HOMO-LUMO overlap ($r$ = 0.037, $p$ = 0.89), $\Delta E_{\text{ST}}$  ($r$ = 0.067, $p$ = 0.80).
+  • PtOEP shows enhanced SOC due to heavy atom effect (Pt).
+Figure 10: Correlation analysis of spin-orbit coupling with structural descriptors.
+(a) SOC vs.
+HOMO-LUMO overlap, (b) SOC vs.
+singlet-triplet gap, (c) SOC vs.
+oscillator strength, (d) SOC distribution for organic molecules.
+Weak correlations ($r<0.2$) indicate that SOC depends on complex orbital angular momentum factors beyond simple geometric descriptors.
+PtOEP (red star) shows enhanced SOC due to heavy atom effect.
+Adiabatic $\Delta E_{\text{ST}}$  Validation
+Computational Methodology
+[…]
+Choice of Functional
+[…]
+[…]
+Dispersion correction: The D4 dispersion correction304 represents the latest generation of Grimme’s dispersion schemes, providing improved treatment of non-covalent interactions important in $\pi$-stacked and sterically hindered TADF architectures.
+[…]
+Comparison with xTB Adiabatic Estimates
+Table 9 compares xTB adiabatic $\Delta E_{\mathrm{ST}}$ estimates (from both sTDA and sTD-DFT methods) with TD-DFT adiabatic values obtained from excited-state geometry optimizations.
+Figure 11 shows the correlation between xTB and TD-DFT adiabatic energies.
+Table 9: Comparison of xTB adiabatic $\Delta E_{\mathrm{ST}}$ estimates with TD-DFT adiabatic values from excited-state geometry optimizations. All values in eV.
+Molecule xTB sTDA xTB sTD-DFT TD-DFT Adiabatic
+DMAC-TRZ 0.049 0.051 1.018
+SpiroAC-TRZ 0.060 0.062 1.000
+2CzPN 0.268 0.259 1.046
+PBTPA 0.263 0.270 1.233
+2BrCPT 0.302 0.272 1.041
+BACN 0.523 0.486 1.171
+Statistics:
+sTDA: $R^{2}=0.395$, Spearman $\rho=0.543$ ($p=0.266$)
+sTD-DFT: $R^{2}=0.454$, Spearman $\rho=0.657$ ($p=0.156$)
+Figure 11: Correlation between xTB adiabatic $\Delta E_{\mathrm{ST}}$ estimates and TD-DFT adiabatic values from excited-state geometry optimizations.
+(a) sTDA method shows moderate correlation ($R^{2}=0.40$, $\rho=0.54$).
+(b) sTD-DFT method shows improved correlation ($R^{2}=0.45$, $\rho=0.66$) due to better treatment of electron correlation.
+Both methods preserve molecular rankings, validating the use of xTB for high-throughput screening.
+Red dashed lines show linear fits.
+Systematic Comparison with Experiment
+To assess the accuracy of our TD-DFT approach, we compared adiabatic $\Delta E_{\mathrm{ST}}$ values with available experimental data (Table 10).
+For the three molecules with experimental measurements, we observe systematic overestimation by approximately 0.9–1.0 eV.
+This discrepancy is consistent with known limitations of TD-DFT for strongly CT-dominated excited states in TADF systems,49, 8 where the single-reference nature of TD-DFT struggles to fully capture the multi-configurational character of near-degenerate singlet-triplet states.
+Table 10: Comparison of TD-DFT adiabatic $\Delta E_{\mathrm{ST}}$ with experimental values. All values in eV.
+Molecule Experimental TD-DFT Difference
+DMAC-TRZ 0.111 1.018 +0.91
+SpiroAC-TRZ 0.11305 1.000 +0.89
+2CzPN 0.08306 1.046 +0.97
+Mean overestimation: 0.92 $\pm$ 0.04 eV
+[…]
+Conclusions from Adiabatic Validation
+The adiabatic validation demonstrates that:
+  1. 1.
+    xTB adiabatic estimates show moderate correlation with TD-DFT adiabatic values ($R^{2}=0.40$–0.45, $\rho=0.54$–0.66).
+  2. 2.
+    sTD-DFT performs slightly better than sTDA, as expected from improved electron correlation treatment.
+  3. 3.
+    Molecular rankings are preserved despite systematic underestimation by xTB.
+  4. 4.
+    TD-DFT systematically overestimates experimental $\Delta E_{\text{ST}}$  by $\sim$1 eV, a known limitation for CT states.
+  5. 5.
+    The vertical approximation employed in our screening is validated for identifying promising TADF candidates through relative comparisons.